### v4

# On $\mathcal{F}$-semi-transitive families of operators

Abstract.
Motivated by the concept of disjoint topological transitivity for supercyclicity given in [30], in this paper, we introduce the concept of disjoint Furstenberg-semi-transitivity. Then we provide sufficient conditions for cosine operator functions generated by adjoints of weighted composition operators to be Furstenberg-semi-transitive on the space of Radon measures on a locally compact, non-compact Hausdorff space. Also, we characterize disjoint Furstenberg-semi-transitive generalized bilateral weighted shifts on the standard Hilbert module over the C*-algebra of compact operators on a separable Hilbert space. As a special case, we consider the case when the corresponding Furstenberg family under consideration is the family of all infinite subsets of the set of natural numbers. Finally, we illustrate our results with concrete examples.
Keywords:
Mathematics Subject Classification (2010) Primary MSC 47A16, Secondary MSC 54H20.
1. Introduction
The research on the dynamics of weighted composition operators has started from pioneering papers by Salas on the dynamics on weighted shifts, see [33, 35]. Since then, numerous papers on the dynamics of weighted composition operators on various function spaces have been published. In particular, the dynamics of cosine operator functions generated by weighted composition operators has been studied in for instance [10, 11, 12, 13, 14, 23, 37]. Moreover, the dynamics of cosine operator functions generated by some other operators has been investigated in for instance [22, 27, 28]. Further, Furstenberg topological transitivity of weighted composition operators, as well as of some other classes of operators, has been studied in several papers, see [2, 7, 18, 19, 29]. Finally, quite recently, topological transitivity of the adjoints of weighted composition operators has been investigated and studied in [22]. These adjoints considered in [22] act on the space of Radon measures on a locally compact, non-compact Hausdorff space. All these facts motivated us to study in this paper the Furstenberg-dynamics of cosine operator functions generated by the adjoints of weighted composition operators. In Section 3 in this paper, we introduce first the concept of $\mathcal{F}$-semi-transitivity, as a generalization of the concept of topological transitivity for supercyclicity given in [30] and as an extension of the notion of $\mathcal{F}$--transitivity Then, in Theorem 3.1, we obtain sufficient conditions for cosine operator functions generated by the adjoints of weighted composition operators to be $\mathcal{F}$-semi-transitive on the space of Radon measures (on a locally compact, non-compact Hausdorff space) equipped with the total variation norm. A special case of this theory is the case when the corresponding Furstenberg family under consideration is the family of all infinite subsets of $\mathbb{N},$ as given in Corollary 3.2. We provide also concrete examples where the sufficient conditions of Theorem 3.1 and Corollary 3.2 in this paper are satisfied, but where the sufficient conditions of [22, Proposition 3.3] are not satisfied, illustrating thus the difference between (ordinary) topological transitivity and $\mathcal{F}$-semi-transitivity of cosine operator functions.
As mentioned above, our concept of $\mathcal{F}$-semi-transitivity is motivated by the concept of topological transitivity for supercyclicity given in [30]. We recall in this context that supercyclicity is a fundamental concept in the dynamics of operators and numerous papers have been written on the topics related to supercyclicity of operators. For example, Hilden and Wallen in [20] proved that any unilateral backward weighted shift is supercyclic. Afterwards, Salas in [35] characterized supercyclic bilateral weighted shift operators on $l^{p}(\mathbb{Z})$ in terms of a supercyclicity criterion. Supercyclicity of several kinds of operators has also been studied in for instance [16, 30, 32, 38]. Moreover, there is a close connection between supercyclicity and semi-Fredholm theory, see [1, 9, 17]. On the other hand, after pioneering papers by Salas on the dynamics on weighted shifts, several papers on the dynamics of various kinds of generalized weighted shifts have been published, see [24, 25, 39]. In particular, the dynamics of generalized bilateral weighted shifts on the standard Hilbert module over the C*-algebra of compact operators on a separable Hilbert space has been studied in [25, 34], however, the concepts such as supercyclicity and disjoint supercyclicity have not been considered in [25, 34]. This motivated us to study disjoint supercyclicity and related concepts for the generalized weighted shift operator considered in [25, 34]. In Section 4 in this paper, we present first the concept of disjoint $\mathcal{F}$-semi-transitivity as a generalization of disjoint topological transitivity for supercyclicity introduced in [30], and then, as a continuation and extension of the results in [25, 34], in Theorem 4.2 we characterize disjoint $\mathcal{F}$-semi-transitive generalized weighted bilateral shift operators on the standard Hilbert module over the C*-algebra of compact operators on a separable Hilbert space in the case when the corresponding Furstenberg family is finitely invariant. Once again it turns out that a special case of this theory is the case when the corresponding Furstenberg family under consideration is the family of all infinite subsets of $\mathbb{N}.$ This special case is treated in Corollary 4.3 and Proposition 4.4 in this paper. At the end of the paper, we give concrete examples of $\mathcal{F}$-semi-transitive generalized weighted bilateral shift operators (Example 4.5).
2. Preliminaries
Throughout this paper, we let $\Omega$ be a locally compact non-compact Hausdorff space and $\alpha$ be an a homeomorphism of $\Omega.$ As usual, $C_{0}(\Omega)$ denotes the space of all continuous functions on $\Omega$ vanishing at infinity, $C_{b}(\Omega)$ denotes the space of bounded, continuous functions on $\Omega$, whereas $C_{c}(\Omega)$ stands for the set of all continuous, compactly supported functions on $\Omega$. Both $C_{0}(\Omega)$ and $C_{b}(\Omega)$ are equipped with the supremum norm. Moreover, we let $w$ be a positive bounded function on $\Omega$ such that also $w^{-1}:=\frac{1}{w}\in C_{b}(\Omega)$ and we put then $T_{\alpha,w}$ to be the weighted composition operator on $C_{0}(\Omega)$ with respect to $\alpha$ and $w,$ that is $T_{\alpha,w}(f)=w\cdot(f\circ\alpha)$ for all $f\in C_{0}(\Omega).$ Easily, one can see that by the above assumptions $T_{\alpha,w}$ is well-defined and $\|T_{\alpha,w}\|\leq\|w\|_{\sup}$. Since $\frac{1}{w}$ is also bounded, then $T_{\alpha,w}$ is invertible and we have
$T_{\alpha,w}^{-1}f=\frac{f\circ\alpha^{-1}}{w\circ\alpha^{-1}},\quad(f\in C_{0}(\Omega)).$
Simply we denote $S_{\alpha,w}:=T_{\alpha,w}^{-1}$.
By some calculation one can see that for each $n\in\mathbb{N}$ and $f\in C_{0}(\Omega)$,
$T^{n}_{\alpha,w}f=\left(\prod_{j=0}^{n-1}(w\circ\alpha^{j})\right)\cdot(f\circ\alpha^{n})\text{ (1)}$
and
$S^{n}_{\alpha,w}f=\left(\prod_{j=1}^{n}(w\circ\alpha^{-j})\right)^{-1}\cdot(f\circ\alpha^{-n}).\text{ (2)}$
The adjoint $T_{\alpha,w}^{\ast}$ is a bounded operator on $M(\Omega)$ where $M(\Omega)$ stands for the Banach space of all Radon measures on $\Omega$ equipped with the total variation norm. It is straightforward to check that
$T_{\alpha,w}^{\ast}(\mu)(E)=\int_{E}w\circ\alpha^{-1}\,d\mu\ \circ\alpha^{-1}$
for every $\mu\in M(\Omega)$ and every measurable subset $E$ od $\Omega.$ Here $\mu\ \circ\alpha^{-1}(E)=\mu(\alpha^{-1}(E))$ for every $\mu\in M(\Omega)$ and every measurable subset $E$ od $\Omega.$ By $(1)$ and $(2)$ it follows then that for every $n\in\mathbb{N},$ $\mu\in M(\Omega)$ and a Borel measurable subset $E\subseteq\Omega$ we have that
${T_{\alpha,w}^{\ast n}}(\mu)(E)=\int_{E}\prod_{j=0}^{n-1}w\circ\alpha^{j-n}\,d\mu\ \circ\alpha^{-n}$
and
${S_{\alpha,w}^{\ast n}}(\mu)(E)=\int_{E}\prod_{j=1}^{n}(w\circ\alpha^{n-j})^{-1}\,d\mu\ \circ\alpha^{n}.$
For every Radon measure $\mu$ on $\Omega,$ we let as usual $|\mu|$ denote the total variation of $\mu.$
In this paper, $\mathcal{F}$ represents a Furstenberg family, which is a specific collection of subsets satisfying the following conditions:
  1. (1)
    The empty set $\emptyset$ is not an element of $\mathcal{F}$.
  2. (2)
    If $A$ belongs to $\mathcal{F}$ and $B$ contains $A$, then $B$ also belongs to $\mathcal{F}$.
3. $\mathcal{F}$-semi-transitive cosine operator functions
Assume that $S$ is a set and $\mathcal{F}$ is a proper Furstenberg family of subsets of $S.$
Let $(\alpha_{z})_{z\in\in\mathcal{S}}$ be a family of homeomorphisms of $\Omega$ and $(w_{z})_{z\in\in\mathcal{S}}$ be a family of positive bounded continuous functions on $\Omega$ such that $\frac{1}{w_{z}}$ is also bounded for all $z\in\mathcal{S}$. For each $z\in\mathcal{S}$, we let $T_{z}^{\ast}:=T_{\alpha_{z},w_{z}}^{\ast}$ and let $C_{z}^{\ast}:=\frac{1}{2}(T_{z}^{\ast}+S_{z}^{\ast})$. We will say that $\{C_{z}^{\ast}\}_{z\in\mathcal{S}}$ is $\mathcal{F}$-semi-transitive if for every non-empty open subsets $U_{1}$ and $U_{2}$ of $M(\Omega)$, the set
$\{z\in\mathcal{S}\text{ : }\exists\lambda_{z}\in\mathbb{R}^{+}\text{ such that }\lambda_{z}C_{z}^{\ast}\ (U_{1})\cap U_{2}\neq\emptyset\}$
is an element of $\mathcal{F}$.
The next theorem gives some sufficient conditions for the family $\{C_{z}^{\ast}\}_{z\in\mathcal{S}}$ to be $\mathcal{F}$-semi-transitive.
Theorem 3.1.
We have $(ii)\Rightarrow(i)$.
  1. $(i)$
    $\{C_{z}^{\ast}\}_{z\in\mathcal{S}}$ is $\mathcal{F}$-semi-transitive;
  2. $(ii)$
    for each compact subset $K$ of $\Omega$, each $\varepsilon>0$, and any two non-zero measures $\mu$ $\nu\in M(\Omega)$ with $|\mu|(K^{c})=|\nu|(K^{c})=0$, there exists some $F\in\mathcal{F}$ and collections of Borel subsets $\{A_{z}\}_{z\in F}$, $\{D_{z}\}_{z\in F}$$\{E_{z}\}_{z\in F}$ of $K$ such that
    $|\mu|(A_{z})<\varepsilon\text{, }|\nu|(A_{z})<\varepsilon\text{, }D_{z}\cap E_{z}=\emptyset\text{, }A^{c}_{z}\cap K=D_{z}\cup E_{z},$
    for all $z\in F$, and
    $\left(\sup_{t\in A^{c}_{z}\cap K}w_{z}(t)\right)\cdot\left(\sup_{t\in D_{z}}w_{z}(t)\right)<\varepsilon,$
    $\left(\sup_{t\in A^{c}_{z}\cap K}w_{z}(t)\right)\cdot\left(\sup_{t\in E_{z}}\left(w_{z}\circ\alpha^{-1}\right)^{-1}(t)\right)<\varepsilon,$
    $\left(\sup_{t\in A^{c}_{z}\cap K}\left(w_{z}\circ\alpha^{-1}\right)^{-1}(t)\right)\cdot\left(\sup_{t\in D_{z}}w_{z}(t)\right)<\varepsilon,$
    $\left(\sup_{t\in A^{c}_{z}\cap K}\left(w_{z}\circ\alpha^{-1}\right)^{-1}(t)\right)\cdot\left(\sup_{t\in E_{z}}\left(w_{z}\circ\alpha^{-1}\right)^{-1}(t)\right)<\varepsilon,$
    $\sup_{t\in D_{z}}\left(\prod_{j=0}^{1}\left(w_{z}\circ\alpha_{z}^{j}\right)(t)\right)<\varepsilon,$
    $\sup_{t\in E_{z}}\left(\prod_{j=1}^{2}\left(w_{z}\circ\alpha_{z}^{-j}\right)^{-1}(t)\right)<\varepsilon\ \ \text{ for all }z\in F.$
Proof.
Let $U_{1}$ and $U_{2}$ be two open non-empty subsets of $M(\Omega)$. Then $U_{1}\setminus\{0\}$ and $U_{2}\setminus\{0\}$ are also open and non-empty subsets of $M(\Omega)$. As in the proof of [22, Proposition 3.1] part $(ii)\Rightarrow(i)$, we can find two measures $\mu\in U_{1}\setminus\{0\}$, $\nu\in U_{2}\setminus\{0\}$, and a compact subset $K$ of $\Omega$ such that $|\mu|(K^{c})=|\nu|(K^{c})=0$. Also, we can find some $\delta>0$ such that $B(\mu,\delta)\subset U_{1}\setminus\{0\}$ and $B(\nu,\delta)\subset U_{2}\setminus\{0\}$. Choose $F\in\mathcal{F}$ and collections of Borel subsets $\{A_{z}\}_{z\in F}$, $\{D_{z}\}_{z\in F}$$\{E_{z}\}_{z\in F}$ satisfying the assumptions of $(ii)$ with respect to $\mu$, $\nu$, and
$\varepsilon:=\min\left\{\frac{\delta}{4}\text{, }\frac{\delta^{2}}{64\|\mu\|\|\nu\|}\text{, }\frac{\delta}{8\|\nu\|}\right\}.$
For each $z\in F$, let $\tilde{\mu_{z}}$, $\tilde{\nu_{z}}$, $\tilde{\tilde{\nu_{z}}}$ be the elements of $M(\Omega)$ given by $\tilde{\mu_{z}}(B)=\mu(B\cap A_{z}^{c}\cap K)$, $\tilde{\nu_{z}}(B)=\nu(B\cap D_{z})$, and $\tilde{\tilde{\nu_{z}}}(B)=\nu(B\cap E_{z})$ for every $z\in F$ and any Borel subset $B$ of $\Omega$. Then
$|\mu-\tilde{\mu_{z}}|(\Omega)=|\mu|(A_{z})<\varepsilon<\delta\ \ \text{ and }\ \ |\mu-\tilde{\nu_{z}}-\tilde{\tilde{\nu_{z}}}|(\Omega)=|\nu|(A_{z})<\varepsilon<\delta,$
[…]
$\|T^{\ast}_{z}(\tilde{\mu_{z}})\|+\|S^{\ast}_{z}(\tilde{\mu_{z}})\|>0\ \ \text{ and }\|T^{\ast}_{z}(\tilde{\nu_{z}})\|+\|S^{\ast}_{z}(\tilde{\tilde{\nu_{z}}})\|>0.$
For each $z\in F$, put
[…]
Then
$\|\phi_{z}-\mu\|\leq\|\mu-\tilde{\mu_{z}}\|+2\sqrt{\|T^{\ast}_{z}(\tilde{\mu_{z}})\|+\|S^{\ast}_{z}(\tilde{\mu_{z}})\|}\sqrt{\|T^{\ast}_{z}(\tilde{\nu_{z}})\|+\|S^{\ast}_{z}(\tilde{\tilde{\nu_{z}}})\|}$
and
$\left\|\frac{\sqrt{\|T^{\ast}_{z}(\tilde{\nu_{z}})\|+\|S^{\ast}_{z}(\tilde{\tilde{\nu_{z}}})\|}}{\sqrt{\|T^{\ast}_{z}(\tilde{\mu_{z}})\|+\|S^{\ast}_{z}(\tilde{\mu_{z}})\|}}C^{\ast}_{z}(\phi_{z})-\nu\right\|$ $\leq\frac{1}{2}\sqrt{\|T^{\ast}_{z}(\tilde{\nu_{z}})\|+\|S^{\ast}_{z}(\tilde{\tilde{\nu_{z}}})\|}\ \sqrt{\|T^{\ast}_{z}(\tilde{\mu_{z}})\|+\|S^{\ast}_{z}(\tilde{\mu_{z}})\|}$
$+\|(T^{\ast}_{z})^{2}(\tilde{\nu_{z}})\|+\|(S^{\ast}_{z})^{2}(\tilde{\tilde{\nu_{z}}})\|+\|\tilde{\nu_{z}}+\tilde{\tilde{\nu_{z}}}-\nu\|,$
for all $z\in F$. Now, by exactly the same arguments as in the proof of [22, Proposition 3.1] part $(ii)\Rightarrow(i)$, one can show that for each $z\in F$ we have
$\left\|T^{\ast}_{z}(\tilde{\mu_{z}})\right\|\leq\sup_{t\in A^{c}_{z}\cap K}\left\{w_{z}(t)\|\mu\|\right\}\text{,}$
$\left\|S^{\ast}_{z}(\tilde{\mu_{z}})\right\|\leq\sup_{t\in A^{c}_{z}\cap K}\left\{(w_{z}\circ\alpha_{z}^{-1})^{-1}(t)\|\mu\|\right\}\text{,}$
$\left\|T^{\ast}_{z}(\tilde{\nu_{z}})\right\|\leq\sup_{t\in D_{z}}\left\{w_{z}(t)\|\nu\|\right\}\text{,}$
$\left\|S^{\ast}_{z}(\tilde{\tilde{\nu_{z}}})\right\|\leq\sup_{t\in E_{z}}\left\{(w_{z}\circ\alpha_{z}^{-1})^{-1}(t)\|\nu\|\right\}\text{,}$
$\left\|(T^{\ast}_{z})^{2}(\tilde{\nu_{z}})\right\|\leq\sup_{t\in D_{z}}\left\{\prod_{j=0}^{1}(w_{z}\circ\alpha_{z}^{j})(t)\|\nu\|\right\}\text{,}$
and
$\left\|(S^{\ast}_{z})^{2}(\tilde{\tilde{\nu_{z}}})\right\|\leq\sup_{t\in E_{z}}\left\{\prod_{j=1}^{2}(w_{z}\circ\alpha_{z}^{-j})^{-1}(t)\|\nu\|\right\}\text{.}$
Utilizing this fact together with assumptions $(ii)$ w. r. to $\varepsilon=\min\left\{\frac{\delta}{4}\text{, }\frac{\delta^{2}}{64\|\mu\|\|\nu\|}\text{, }\frac{\delta}{8\|\nu\|}\right\}$ and the fact that $B(\mu,\delta)\subset U_{1}$ and $B(\nu,\delta)\subset U_{2},$ we deduce that
$\frac{\sqrt{\|T^{\ast}_{z}(\tilde{\nu_{z}})\|+\|S^{\ast}_{z}(\tilde{\tilde{\nu_{z}}})\|}}{\sqrt{\|T^{\ast}_{z}(\tilde{\mu_{z}})\|+\|S^{\ast}_{z}(\tilde{\mu_{z}})\|}}C^{\ast}_{z}(\phi_{z})\in U_{1}$
for all $z\in F$. Hence
$F\subseteq\{z\in\mathcal{S}\text{ : }\exists\lambda_{z}\in\mathbb{R}^{+}\text{ such that }\lambda_{z}C^{\ast}(U_{1})\cap U_{2}\neq\emptyset\}.$
Since $\mathcal{F}$ is Furstenberg , we obtain that
$\{z\in\mathcal{S}\text{ : }\exists\lambda_{z}\in\mathbb{R}^{+}\text{ such that }\lambda_{z}C^{\ast}(U_{1})\cap U_{2}\neq\emptyset\}\in\mathcal{F}.$
Hence, $\{C_{z}^{\ast}\}_{z\in\mathcal{S}}$ is $\mathcal{F}$-semi-transitive since $U_{1}$ and $U_{2}$ were arbitrary. ∎
Corollary 3.2.
Let $w$ be a continuous positive function on $\Omega$ such that $w$ and $\frac{1}{w}$ are bounded and let $\alpha$ be a homeomorphism of $\Omega$. Suppose that for each compact subset $K$ of $\Omega$ it holds that
$\lim_{n\longrightarrow\infty}\left(\sup_{t\in K}\prod_{j=0}^{n-1}\left(w\circ\alpha^{j}\right)(t)\right)\left(\sup_{t\in K}\prod_{j=1}^{n}\left(w\circ\alpha^{-j}\right)^{-1}(t)\right)=0$
and either
$\lim_{n\longrightarrow\infty}\left(\sup_{t\in K}\prod_{j=0}^{2n-1}\left(w\circ\alpha^{j}\right)(t)\right)=0$
or
$\lim_{n\longrightarrow\infty}\left(\sup_{t\in K}\prod_{j=1}^{2n}\left(w\circ\alpha^{-j}\right)^{-1}(t)\right)=0.$
Then, for each non-empty open subsets $U_{1}$ and $U_{2}$ of $M(\Omega)$, there exists some $N\in\mathbb{N}$ and some sequence $\{\lambda_{n}\}_{n=N}^{\infty}$ of positive numbers such that $\lambda_{n}C_{n}^{\ast}(U_{1})\cap U_{2}\neq\emptyset\text{ for all }n\geq N,$ where $C_{n}^{\ast}=\frac{1}{2}\left(T^{\ast n}_{\alpha,w}+S^{\ast n}_{\alpha,w}\right)$.
Proof.
Let $S=\mathbb{N}$ and $\mathcal{F}$ be the family of all infinite subset of $\mathbb{N}$. For each $n\in\mathbb{N}$, we let $\alpha_{n}:=\alpha^{n}$ and $w_{n}:=\prod_{j=0}^{n-1}\left(w\circ\alpha^{j}\right)$. Then we consider the case when $A_{n}=E_{n}=\emptyset$ and $D_{n}=K$ (such that $\tilde{\tilde{\nu_{n}}}=0$) for all $n\in\mathbb{N}$, and the case when $A_{n}=D_{n}=\emptyset$ and $E_{n}=K$ for all $n\in\mathbb{N}$, (such that $\tilde{\nu_{n}}=0$ for all $n\in\mathbb{N}$). By adapting the proof of Theorem 3.1 for these special cases, we can deduce the statement of corollary. ∎
Now we will provide a concrete example where the conditions of Corollary 3.2 are satisfied.
Example 3.3.
Let $X=\mathbb{R}$ and $M,\delta$ be positive constants such that $M\geq 2+2\delta$ and $\delta\geq 1.$ Put $\alpha$ to be the function on $\mathbb{R}$ given by $\alpha(t)=t+1$ for all $t\in\mathbb{R}.$ Set
$w(t)=\begin{cases}M\text{ for }t\leq-1,\\ M+\frac{t+1}{2}(1+\delta-M)\text{ for }t\in[-1,1],\\ 1+\delta\text{ for }t\geq 1.\\ \end{cases}.$
By some calculations one can check that for each compact subset $K$ of $\Omega$ it holds that
$\lim_{n\longrightarrow\infty}\left(\sup_{t\in K}\prod_{j=0}^{n-1}\left(w\circ\alpha^{j}\right)(t)\right)\left(\sup_{t\in K}\prod_{j=1}^{n}\left(w\circ\alpha^{-j}\right)^{-1}(t)\right)=0$
and
$\lim_{n\longrightarrow\infty}\left(\sup_{t\in K}\prod_{j=1}^{2n}\left(w\circ\alpha^{-j}\right)^{-1}(t)\right)=0,$
so the conditions of Corollary 3.2 are satisfied. However, it is not hard to check by some calculations that for every strictly increasing sequence $\{n_{k}\}\subseteq\mathbb{N}$ and every $t\in\mathbb{R}$ it holds that
$\lim_{k\rightarrow\infty}\text{ }\prod_{j=0}^{n_{k}-1}(w\circ\alpha^{j})(t)=\infty,$
hence the conditions of [22, Proposition 3.3] are not satisfied. This illustrates the difference between the sufficient conditions of Corollary 3.2 and the sufficient conditions of [22, Proposition 3.3].
4. Disjoint $\mathcal{F}$-semi-transitive generalized weighted shifts
We start with the following definition.
Definition 4.1.
Let $S$ be a set and let $\mathcal{F}$ be a proper Furstenberg family of subsets of $S$ and $\{T_{t,1}\}_{t\in S},\dots,\{T_{t,N}\}_{t\in S}$ be families of bounded linear operators on a Banach space $X$. We say that $\{T_{t,1}\}_{t\in S},\dots,\{T_{t,N}\}_{t\in S}$ are disjoint $\mathcal{F}$-semi-transitive, or shortly $d\mathcal{F}$-semi-transitive, if for every collection of non-empty open subsets $\mathcal{O},V_{1},\dots,V_{N}$ of $X$, the set
$\{t\in S\ |\text{ there exists }\lambda_{t}\in\mathbb{R}^{+}\text{ such that }\mathcal{O}\cap\lambda_{t}T_{t,1}^{-1}(V_{1})\cap\dots\cap\lambda_{t}T_{t,N}^{-1}(V_{N})\neq\varnothing\}$
belongs to $\mathcal{F}$.
In this section, for a separable Hilbert space $H$ we let $B(H)$ be the space of all bounded linear operators on $H$, and $\mathcal{C}$ be the $C^{*}$-algebra of compact operators on $H$. For an orthonormal basis $\{e_{j}\}_{j\in\mathbb{Z}}$ for $H$, we let for each $m\in\mathbb{N}$, $P_{m}$ be the orthogonal projection onto $\text{Span}\{e_{-m},\dots,e_{m}\}$. We will denote by $\ell_{2}(\mathcal{C})$ the standard Hilbert module over $\mathcal{C}$, see [31, Example 1.3.5]. Notice that $\ell_{2}(\mathcal{C})$ is a Banach algebra. Indeed, we can define multiplication on $\ell_{2}(\mathcal{C})$ as pointwise multiplication, i.e., if $\{x_{j}\}_{j\in\mathbb{Z}},\{y_{j}\}_{j\in\mathbb{Z}}\in\ell_{2}(\mathcal{C})$, then
$\{x_{j}\}_{j\in\mathbb{Z}}\cdot\{y_{j}\}_{j\in\mathbb{Z}}=\{x_{j}y_{j}\}_{j\in\mathbb{Z}}.$
To see that $\{x_{j}y_{j}\}_{j\in\mathbb{Z}}$ belongs to $\ell_{2}(\mathcal{C})$, it suffices to observe that
$\sum_{j}y_{j}^{*}x_{j}^{*}x_{j}y_{j}\leq\sum_{j}y_{j}^{*}\|x_{j}\|^{2}y_{j}\leq\|x\|^{2}\sum_{j}y_{j}^{*}y_{j}\leq\parallel\{x_{j}\}_{j\in\mathbb{Z}}\parallel^{2}\sum y_{j}^{*}y_{j}.$
For each $m,J\in\mathbb{N}$ we let $\tilde{p}_{J,m}\in\ell_{2}(\mathcal{C})$ be given by
[…]
As shown in the proof of [25, Proposition 3.1], the set $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ forms a left approximate unit for $\ell_{2}(\mathcal{C})$.
For $N\in\mathbb{N},$ let $r_{1},\dots,r_{N}\in\mathbb{N}$ with $r_{1}<r_{2}<\dots<r_{N}$. Further, for each $n\in\mathbb{N}$ and $s\in\{1,\dots,N\},$ let $\{W_{n,j}^{(s)}\}_{j\in\mathbb{Z}}$ be a family of operators in $B(H)$ which is uniformly bounded in norm and such that each $W_{n,j}^{(s)}$ has a bounded inverse. Also, assume that $\mathcal{U}^{(1)},\dots,\mathcal{U}^{(N)}$ are unitary operators on $H$. Then, for each $n\in\mathbb{N}$ and $s\in\{1,\dots,N\},$ we let $T_{n,s}$ be the operator on $\ell_{2}(\mathcal{\mathcal{C}})$ defined as
$\big(T_{n,s}(\{x_{i}\}_{i\in\mathbb{Z}})\big)=\{W_{n,j}^{(s)}\,x_{j-r_{s}n}\mathcal{U}^{(s)r_{s}n}\}_{j\in\mathbb{Z}}.$
By [25], $T_{n,s}$ is a bounded linear operator on $\ell_{2}(\mathcal{\mathcal{C}})$ for each $n\in\mathbb{N}$ and $s\in\{1,\dots,N\}$.
In what follows, we will assume that $\mathcal{F}$ is a Furstenberg family of subsets of $\mathbb{N}$ which is finitely invariant, that is, for each $n\in\mathbb{N}$ and $F\in\mathcal{F}$ it holds that $F\setminus\{1,\dots,n\}\in\mathcal{F}$ also.
Now we are ready to give the main result of this section.
Theorem 4.2.
Under the above notation and assumptions, the following statements are equivalent.
(1) The families of operators $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
(2) For every $J,m\in\mathbb{N}$ and $\varepsilon>0$, there exists some $F\in\mathcal{F}$ such that for each $j\in\mathbb{Z}$ with $J\leq j\leq J$ we can find families $\{D_{n,j}\}_{n\in F},\quad\{G_{n,j}^{(1)}\}_{n\in F},\dots,\{G_{n,j}^{(N)}\}_{n\in F}$ of compact operators on $H$ satisfying
$\|D_{n,j}-P_{m}\|<\varepsilon,\quad\|G_{n,j}^{(s)}-P_{m}\|<\varepsilon,$
and
$\|W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|\,\|W_{n,j}^{(s)\,-1}G_{n,j}^{(s)}\|<\varepsilon^{2}\quad\text{for all }n\in F\text{ and }\ell,s\in\{1,\dots,N\}.$
and satisfying in addition that for each distinct $s,\ell\in\{1,\dots,N\}$ and all $n\in F$
$\|W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)-1}\,G_{n,j}^{(s)}\|<\varepsilon.$
Proof.
$(1)\Rightarrow(2)$ Let $J,m\in\mathbb{N}$ be given. Since the families $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ are disjoint $\mathcal{F}$-semi-transitive and $\mathcal{F}$ is finitely invariant. given $\varepsilon>0$, we can find some $F\subseteq\mathcal{F}$ with $\{1,\dots,2J\}\cap F=\varnothing$ such that for each $n\in F$, there exists some $x_{n}\in\ell_{2}(\mathcal{C})$ and some $\lambda_{n}\in\mathbb{R}^{+}$ with
$\|(x_{n})-\tilde{p}_{J,m}\|<\varepsilon,$
$\|\lambda_{n}(T_{n,\ell}x_{n})-\tilde{p}_{J,m}\|<\varepsilon\text{ for all }\ell\in\{1,\dots,N\}.$
This gives further that
$\|(x_{n})_{j}-P_{m}\|<\varepsilon,$
and
$\|\lambda_{n}W_{n,j+r_{\ell}n}^{(\ell)}(x_{n})_{j}\,\mathcal{U}^{({\ell})r_{\ell}n}\|<\varepsilon.$
for all $j\in\mathbb{Z}$ with $J\leq j\leq J$ and $\ell\in\{1,\dots,N\}$, because
$(T_{n,\ell}(x_{n}))_{j+r_{\ell}n}=W_{n,j+r_{\ell}n}^{(\ell)}\,(x_{n})_{j}\,\mathcal{U}^{({\ell})r_{\ell}n},$
and $j+r_{\ell}n>J$ for all $j\in\mathbb{Z}$ with $J\leq j\leq J$ and $n\in F$.
Hence
$|\lambda_{n}|\,\|W_{n,j+r_{\ell}n}^{(\ell)}(x_{n})_{j}\|=$
$\|\lambda_{n}W_{n,j+r_{\ell}n}^{(\ell)}(x_{n})_{j}\,\mathcal{U}^{(\ell)r_{\ell}n}\|<\varepsilon$
for all $n\in F$, $j\in\mathbb{Z}$ with $J\leq j\leq J$ and $\ell\in\{1,\dots,N\}$. Further, since
$\|\lambda_{n}(T_{n,s}\,x_{n})_{j}-P_{m}\|<\varepsilon\quad\text{for all }j\in\mathbb{Z}\text{ with }-J\leq j\leq J\text{ and }s\in\{1,\dots,N\},$
it follows that
$\|\lambda_{n}W_{n,j}^{(s)}(x_{n})_{j-r_{s}n}\,\mathcal{U}^{(s)r_{s}n}-P_{m}\|<\varepsilon,$
however,
$\frac{1}{|\lambda_{n}|}\,\|W_{n,j}^{(s)-1}\lambda_{n}W_{n,j}^{(s)}(x_{n})_{j-r_{s}n}\,\mathcal{U}^{(s)r_{s}n}\|=\|(x_{n})_{j-r_{s}n}\,\mathcal{U}^{(s)r_{s}n}\|$
$\|(x_{n})_{j-r_{s}n}\|\leq\|x_{n}-\tilde{p}_{J,m}\|<\varepsilon,$
once again because $j-r_{s}n<J-2J\leq-J.$ For each $n\in F,$ $\ell,s\in\{1,\dots,N\}$ and $j\in\mathbb{Z}$ with $J\leq j\leq J$, define
$D_{n,j}:=(x_{n})_{j}\text{ and }\quad G_{n,j}^{(s)}=\lambda_{n}W_{n,j}^{(s)}(x_{n})_{j-r_{s}n}\,\mathcal{U}^{(s)r_{s}n}.$
Then
$\|D_{n,j}-P_{m}\|<\varepsilon,\quad\|G_{n,j}^{(s)}-P_{m}\|<\varepsilon,$
and
$|\lambda_{n}|\,\|W_{n,j+r_{\ell}n}^{(\ell)\,}D_{n,j}\|\,\tfrac{1}{|\lambda_{n}|}\,\|W_{n,j}^{(s)\,-1}G_{n,j}^{(s)}\|<\varepsilon^{2}.$
Finally, if $\ell,s\in\{1,\dots,N\}$ with $\ell\neq s$, then $|r_{\ell}-r_{s}|n>2J$ for all $n\in F$. Hence, either $j-r_{s}n+r_{\ell}n<-J$ or $j-r_{s}n+r_{\ell}n>J$ for any $j\in\mathbb{Z}$ with $J\leq j\leq J$. Therefore, $(\tilde{p}_{J,m})_{j-r_{s}n+r_{\ell}n}=0,$ hence
$\|W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)\,-1}\,G_{n,j}^{(s)}\|=\big\|\lambda_{n}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}(x_{n})_{j-r_{s}n}\mathcal{U}^{(s)r_{s}n}\big\|$
$\big\|\lambda_{n}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}(x_{n})_{j-r_{s}n}\big\|=\big\|\lambda_{n}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}(x_{n})_{j-r_{s}n}\mathcal{U}^{(\ell)r_{\ell}n}\big\|$
$\big\|\lambda_{n}(T_{n,\ell}(x_{n}))_{j-r_{s}n+r_{\ell}n}\big\|=\big\|(\lambda_{n}T_{n,\ell}(x_{n})-\tilde{p}_{J,m})_{j-r_{s}n+r_{\ell}n}\big\|<\varepsilon.$
for all $n\in F$, $j\in\mathbb{Z}$ with $J\leq j\leq J$, and $s,\ell\in\{1,\dots,N\}$ with $s\neq\ell$.
$(2)\Rightarrow(1)$ Let $\mathcal{O},V_{1},\dots,V_{N}$ be non-empty open subsets of $\ell_{2}(\mathcal{C})$. Then there exist some $J,m\in\mathbb{N}$ and some $x\in\mathcal{O}$, $y^{(1)}\in V_{1},\dots,y^{(N)}\in V_{N}$ such that $\tilde{p}_{J,m}x=x$ and $\tilde{p}_{J,m}y^{(\ell)}=y^{(\ell)}$ for each $\ell\in\{1,\dots,N\}$ since the set $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ forms a left approximate unit for $\ell_{2}(\mathcal{C})$.
For each $a\in\ell_{2}(\mathcal{C}),$ let $B_{\delta}(a)$ denote the open ball with centre in $a$ and radius $\delta$ . We can find some $\delta>0$ such that $B_{\delta}(x)\subseteq\mathcal{O}$ and $B_{\delta}(y^{(\ell)})\subseteq V_{\ell}$ for each $\ell\in\{1,\dots,N\}.$ Put $C_{y}=\sum_{s=1}^{N}\|y^{(s)}\|$ and
$\varepsilon:=\frac{\delta}{2JN(\|x\|+2C_{y}+\delta)}.$
Choose $F\in\mathcal{F}$ and families
$\{D_{n,j}\}_{n\in F},\quad\{G_{n,j}^{(1)}\}_{n\in F},\dots,\{G_{n,j}^{(N)}\}_{n\in F}$
for each $j\in\mathbb{Z}$ with $J\leq j\leq J$, satisfying the assumptions of (2) with respect to $J$, $m$, and $\varepsilon$.
For each $n\in F$, let $u_{n},v_{n}^{(1)},\dots,v_{n}^{(N)}\in\ell_{2}(\mathcal{C})$ be given by
$(u_{n})_{j}=\begin{cases}D_{n,j}\,x_{j},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else.}\end{cases}$
and
$(v_{n}^{(\ell)})_{j-r_{\ell}n}=\begin{cases}W_{n,j}^{(\ell)\,-n}\,G_{n,j}^{(\ell)}\,(y^{(\ell)})_{j}\,\mathcal{U}^{*(\ell)\,r_{\ell}n},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else.}\end{cases}$
Since $\|D_{n,j}-P_{m}\|<\varepsilon$ and $\|G_{n,j}^{(\ell)}-P_{m}\|<\varepsilon<1$ for all $n\in F$, $\ell\in\{1,\dots,N\}$ and $j\in\mathbb{Z}$ with $J\leq j\leq J$, it follows that $D_{n,j}\neq 0$ and $G_{n,j}^{(\ell)}\neq 0,$ hence $W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\neq 0$, and $W_{n,j}^{(\ell)-1}G_{n,j}^{(\ell)}\neq 0$ for all $n\in F$, $\ell\in\{1,\dots,N\}$ and $j\in\mathbb{Z}$ with $J\leq j\leq J$. Now we notice that
$\|T_{n,\ell}(u_{n})\|\leq\sum_{j=-J}^{J}\|W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|\,\|x\|$
because
$(T_{n,\ell}(u_{n}))_{j+r_{\ell}n}=\begin{cases}W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\,x_{j}\,\mathcal{U}^{(\ell)\,r_{\ell}n},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else}.\end{cases}$
Moreover, since for each distinct $s,\ell\in\{1,\dots,N\}$ and all $n\in F$,
$(T_{n,\ell}(v_{n}^{(s)}))_{j-r_{s}n+r_{\ell}n}=$
$\begin{cases}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)-1}\,G_{n,j}^{(s)}(y^{(s)})_{j}\,\mathcal{U}^{*(s)\,r_{s}n}\,\mathcal{U}^{(\ell)\,r_{\ell}n},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else.}\end{cases}$
it follows that
$\|T_{n,\ell}(v_{n}^{(s)})\|\leq\sum_{j=-J}^{J}\|W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)\,-1}\,G_{n,j}^{(s)}\|\,\|y^{(s)}\|.$
Finally, it is easy to see that by the construction of $v_{n}^{(\ell)}$, we have
$T_{n,\ell}(v_{n}^{(\ell)})_{j}=G_{n,j}^{(\ell)}(y^{(\ell)})_{j}\quad\text{if }-J\leq j\leq J,\quad\text{and }0\text{ else.}$
For each $n\in F$, set
$x_{n}:=u_{n}+\frac{\sqrt{\sum_{\ell=1}^{N}\sum_{j=-J}^{J}\|W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|}}{\sqrt{\sum_{\ell=1}^{N}\sum_{j=-J}^{J}\|W_{n,j}^{(\ell)\,-1}G_{n,j}^{(\ell)}\|}}\sum_{\ell=1}^{N}v_{n}^{(\ell)}.$
By the triangle inequality and the choice of $\varepsilon$, it is not hard to deduce by some calculations that
$x_{n}\in\mathcal{O}\quad\text{and}\quad\frac{\sqrt{\sum_{\ell=1}^{N}\sum_{j=-J}^{J}\|W_{n,j}^{(\ell)\,-1}G_{n,j}^{(\ell)}\|}}{\sqrt{\sum_{\ell=1}^{N}\sum_{j=-J}^{J}\|W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|}}T_{n,\ell}(x_{n})\in V_{\ell}$
for all $n\in F$ and $\ell\in\{1,\dots,N\}$.
Let now $\mathcal{F}$ be the family of all infinite subsets of $\mathbb{N}$ and
$\{W_{j}^{(1)}\}_{j\in\mathbb{Z}},\dots,\{W_{j}^{(N)}\}_{j\in\mathbb{Z}}\subset B(H)$
be such that each $W_{j}^{(\ell)}$ $(j\in\mathbb{Z},\,1\leq\ell\leq N)$ is invertible, and such that
$\|W_{j}^{(\ell)}\|\leq M\quad\text{for all }j\in\mathbb{Z},\,\ell\in\{1,\dots,N\},$
and some fixed $M>0$. For each $n\in\mathbb{N},i\in\mathbb{Z}$ and $\ell\in\{1,\dots,N\}$, we let $W_{n,i}^{(l)}\in B(H)$ be given by
$W_{n,i}^{(l)}=W_{i}^{(\ell)}W_{i-1}^{(\ell)}\cdots W_{i-r_{\ell}n+1}^{(\ell)}.$
Finally, for each $\ell\in\{1,\dots,N\}$ we let $\tilde{T}_{\ell}$ be the generalized bilateral weighted shift operator on $\ell_{2}(\mathcal{C})$ (introduced in [25]) with respect to the sequence $\{W_{j}^{(\ell)}\}_{j\in\mathbb{Z}}$ and unitary $\mathcal{U}^{(\ell)}$, i.e.
[…]
Then it is not hard to check that for each $\ell\in\{1,\dots,N\}$ it holds that
[…]
We will now present an application of Theorem 4.2 in this concrete case.
Corollary 4.3.
Under the above notation and assumptions, the following statements are equivalent:
(1) The sequences of operators $\{\tilde{T}_{1}^{r_{1}n}\}_{n\in\mathbb{N}},\dots,\{\tilde{T}_{N}^{r_{N}n}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
(2) For every $J,m\in\mathbb{N}$ there exist a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ and sequences of operators
$\{D_{j}^{(k)}\}_{k\in\mathbb{N}},\quad\{G_{j,1}^{(k)}\}_{k\in\mathbb{N}},\dots,\{G_{j,N}^{(k)}\}_{k\in\mathbb{N}}.$
in $\mathcal{C}$ for every $j\in\mathbb{Z}$ with $J\leq j\leq J$ such that
$\lim_{k\to\infty}\|D_{j}^{(k)}-P_{m}\|=\lim_{k\to\infty}\|G_{j,\ell}^{(k)}-P_{m}\|=0,$
and
$\lim_{k\to\infty}\|W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}G_{j,s}^{(k)}\|\|W_{j+r_{\ell}n_{k}}^{(\ell)}\cdots W_{j+1}^{(\ell)}D_{j}^{(k)}\|=0.$
for all $J\in\mathbb{Z}$ with $J\leq j\leq J$ and $\ell,s\in\{1,\dots,N\}$. Moreover, for each distinct $\ell,s\in\{1,\dots,N\}$ and all $j\in\mathbb{Z}$ with $J\leq j\leq J$, it holds that
$\lim_{k\to\infty}\|W_{j+r_{\ell}n_{k}-r_{s}n_{k}}^{(\ell)}\cdots W_{j-r_{s}n_{k}+1}^{(\ell)}W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}G_{j,s}^{(k)}\|=0.$
In order to construct concrete examples satisfying the conditions $(2)$ of Corollary 4.3, we need first the following technical result.
Proposition 4.4.
Under the above notation and assumptions, if for each $j\in\mathbb{Z}$ there exist dense subsets $H_{0}^{(j)},H_{1}^{(j)},\dots,H_{N}^{(j)}$ of $H$ and a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ such that for each $s,\ell\in\{1,\dots,N\}$ and every $x^{(j)}\in H_{0}^{(j)}$, $y_{s}^{(j)}\in H_{s}^{(j)}$ we have
$\lim_{k\to\infty}\big\|W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}\,y_{s}^{(j)}\big\|\big\|W_{j+r_{\ell}n_{k}}^{(\ell)}\cdots W_{j+1}^{(\ell)}\,x^{(j)}\big\|=0,$
and for each distinct $s,\ell\in\{1,\dots,N\}$ and every $y_{s}^{(j)}\in H_{s}^{(j)}$,we have
$\lim_{k\to\infty}\|W_{j+r_{\ell}n_{k}-r_{s}n_{k}}^{(\ell)}\cdots W_{j-r_{s}n_{k}+1}^{(\ell)}W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}y_{s}^{(j)}\|=0.,$
then he sequences of operators $\{\tilde{T}_{1}^{r_{1}n}\}_{n\in\mathbb{N}},\dots,\{\tilde{T}_{N}^{r_{N}n}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
Proof.
The main idea in this proof is motivated by the proofs of [25, Theorem 3.2] and [26, Proposition 2.7]. Given $m\in\mathbb{N}$ and $j\in\mathbb{Z}$ with $J\leq j\leq J$, for every $\eta\in\{-m,\dots,m\}$ we can find sequences
$\{f_{i,j}^{(\eta)}\}_{i\in\mathbb{N}},\subseteq H_{0}^{(j)},\{h_{i,l,j}^{(\eta)}\}_{i\in\mathbb{N}}\subseteq H_{l}^{(j)}\text{ for each }\ell\in\{1,\dots,N\}$
such that
$f_{i,j}^{(\eta)}\to e_{\eta},\text{ and }h_{i,\ell,j}^{(\eta)}\to e_{\eta}\quad\text{as }i\to\infty,$
for all $\eta\in\{-m,\dots,m\}$ and $\ell\in\{1,\dots,N\}$. Since by the assumption we have for all $\ell,s\in\{1,\dots,N\}$ and $\eta\in\{-m,\dots,m\}$ that
$\lim_{k\to\infty}\big\|W_{j-r_{s}n_{k}+1}^{(s)^{-1}}\cdots W_{j}^{(s)^{-1}}h_{i,s,j}^{(\eta)}\big\|\big\|W_{j+r_{\ell}n_{k}}^{(\ell)}\cdots W_{j+1}^{(\ell)}f_{i,j}^{(\eta)}\big\|=0,$
and for each distinct $s,\ell\in\{1,\dots,N\}$ and $\eta\in\{-m,\dots,m\}$ we have
$\lim_{k\to\infty}\big\|W_{j+r_{\ell}n_{k}-r_{s}n_{k}}^{(\ell)}\dots W_{j-r_{\ell}n_{k}+1}^{(\ell)}W_{j-r_{s}n_{k}+1}^{(s)^{-1}}\dots W_{j}^{(s)^{-1}}\,h_{i,s,j}^{(\eta)}\big\|=0,$
we can inductively easily construct a subsequence $\{n_{k_{i}}\}_{i\in\mathbb{N}}$ such that
$\big\|W_{j-r_{s}n_{k_{i}}}^{(s)^{-1}}\cdots W_{j}^{(s)^{-1}}\,h_{i,s,j}^{(\eta)}\big\|\big\|W_{j+r_{\ell}n_{k_{i}}}^{(\ell)}\cdots W_{j+1}^{(\ell)}\,f_{i,j}^{(\eta)}\big\|<\tfrac{1}{4m^{2}4^{i}},$
and
$\big\|W_{j+r_{\ell}n_{k_{i}}-r_{s}n_{k_{i}}}^{(\ell)}\cdots W_{j-r_{s}n_{k_{i}}+1}^{(\ell)}W_{j-r_{s}n_{k_{i}}+1}^{(s)^{-1}}\cdots W_{j}^{(s)^{-1}}\,h_{i,s,j}^{(\eta)}\big\|<\tfrac{1}{2_{m}2^{i}}.$
For each $i\in\mathbb{N}$ and $\ell\in\{1,\dots,N\}$ we define operators $D_{j}^{(i)}$ and $G_{j,\ell}^{(i)}$ on $H$ by
$(D_{j}^{(i)})\,e_{\eta}=\begin{cases}f_{i,j}^{(\eta)},&\text{for }\eta\in\{-m,\dots,m\},\\[4.0pt] 0,&\text{else},\end{cases}\qquad(G_{j,\ell}^{(i)})\,e_{\eta}=\begin{cases}h_{i,\ell,j}^{(\eta)},&\text{for }\eta\in\{-m,\dots,m\},\\[4.0pt] 0,&\text{else}.\end{cases}$
By some calculations and arguments similar to those in the proof of [25, Theorem 3.2] and [26, Proposition 2.7], we can conclude that the sequences of operators
$\{D_{j}^{(i)}\}_{i\in\mathbb{N}},\quad\{G_{j,1}^{(i)}\}_{i\in\mathbb{N}},\dots,\{G_{j,N}^{(i)}\}_{i\in\mathbb{N}}$
satisfy the conditions in (2) of Corollary 4.3. for every $j\in\mathbb{Z}$ with $J\leq j\leq J$. ∎
Now we are ready to give an example where the conditions of Corollary 4.3 and Proposition 4.4 are satisfied.
Example 4.5.
Let $N=1$ and $H=L^{2}(\mathbb{R})$. As in [25, Example 3.3], choose some $\{w_{j}\}_{j\in\mathbb{Z}}\subseteq L^{\infty}(\mathbb{R})$ such that $w_{j}>0$, $w_{j}^{-1}\in L^{\infty}(\mathbb{R})$ and $\parallel w_{j}\parallel\leq M$ for all $j\in\mathbb{Z}$ and some $M>0$. Let $\{c_{j}\}_{j\in\mathbb{Z}}\subseteq\mathbb{R}^{+}$ be such that $c_{j}\geq C$ for all $j\in\mathbb{Z}$ and some $C>0$. For each $j\in\mathbb{Z}$ let $\alpha_{j}$ be the translation on $\mathbb{R}$ given by $\alpha_{j}(t)=t-c_{j}$ for all $t\in\mathbb{R}$. If there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}\leq 1-\varepsilon$ for all $j\geq 0$ and $w_{j}\chi_{(-\infty,0)}\geq 1$ for all $j<0$; or such that $w_{j}\chi_{[0,\infty)}\leq 1$ for all $j\geq 0$ and $w_{j}\chi_{(-\infty,0)}\geq 1+\varepsilon$ for all $j<0$, then by the same arguments as in [ ] one can show that the conditions of Proposition 4.4 are satisfied also when $W_{j}(t)=w_{j}\cdot(f\circ\alpha_{j})$ for all $f\in L^{2}(\mathbb{R})$ and $j\in\mathbb{Z}$ with $H_{0}^{(j)}=H_{1}^{(j)}=\dots=H_{N}^{(j)}=C_{c}(\mathbb{R})$ for all $j\in\mathbb{Z},$ where $C_{c}(\mathbb{R})$ stands for the space of all continuous compactly supported functions on $\mathbb{R}.$