### v6

# On $\mathcal{F}$-semi-transitive families of operators

Abstract.
Motivated by the concept of disjoint topological transitivity for supercyclicity given in [32], in this paper, we present the concept of disjoint Furstenberg-semi-transitivity. Then we provide sufficient conditions for cosine operator functions generated by adjoints of weighted composition operators to be Furstenberg-semi-transitive on the space of Radon measures on a locally compact, non-compact Hausdorff space. Also, we characterize disjoint Furstenberg-semi-transitive compositions of a right multiplier and a generalized bilateral weighted shift on the standard Hilbert module over the C*-algebra of compact operators on a separable Hilbert space. As a special case, we consider the case when the corresponding Furstenberg family under consideration is the family of all infinite subsets of the set of natural numbers. Finally, we illustrate our results with concrete examples. We end this paper with an algebraic generalization to the case of non-unital Banach algebras.
Keywords:
Mathematics Subject Classification (2010) Primary MSC 47A16, Secondary MSC 54H20.
1. Introduction
Supercyclicity is a fundamental concept in the dynamics of operators and numerous papers have been written on the topics related to supercyclicity of operators. For example, Hilden and Wallen in [21] proved that any unilateral backward weighted shift is supercyclic. Afterwards, Salas in [37] characterized supercyclic bilateral weighted shift operators on $l^{p}(\mathbb{Z})$ in terms of a supercyclicity criterion. Supercyclicity of several kinds of operators has also been studied in for instance [16, 32, 34, 40]. Moreover, there is a close connection between supercyclicity and semi-Fredholm theory, see [1, 9, 17]. On the other hand, Furstenberg topological transitivity as a generalization of topological transitivity and hypercyclicity has been studied in several papers, see [2, 7, 19, 20, 31]. Motivated by these facts, in this paper we consider the concept of $\mathcal{F}$-semi-transitivity, as a generalization of the concept of topological transitivity for supercyclicity given in [32] and characterize $\mathcal{F}$-semi-transitivity classes of operators on various Banach spaces.
Now, the research on the dynamics of weighted composition operators has started from pioneering papers by Salas on the dynamics on weighted shifts, see [36, 37]. Since then, numerous papers on the dynamics of weighted composition operators on various function spaces have been published. In particular, the dynamics of cosine operator functions generated by weighted composition operators has been studied in for instance [10, 11, 12, 13, 14, 25, 39]. Moreover, the dynamics of cosine operator functions generated by some other operators has been investigated in for instance [24, 29, 30]. Finally, quite recently, topological transitivity of the adjoints of weighted composition operators has been investigated and studied in [24]. These adjoints considered in [24] act on the space of Radon measures on a locally compact, non-compact Hausdorff space. Although there are several published papers on the dynamics of cosine operator functions, supercyclicity of cosine operator functions has not been studied in any of these published papers. All these facts naturally lead us to study in this paper supercyclicity and $\mathcal{F}$-semi-transitivity (as a generalization of supercyclicity) of cosine operator functions generated by the adjoints of weighted composition operators. In Section 3 in this paper, after presenting the concept of $\mathcal{F}$-semi-transitivity, we obtain sufficient conditions for cosine operator functions generated by the adjoints of weighted composition operators to be $\mathcal{F}$-semi-transitive on the space of Radon measures (on a locally compact, non-compact Hausdorff space) equipped with the total variation norm (Theorem 3.1). A special case of this theory is the case when the corresponding Furstenberg family under consideration is the family of all infinite subsets of $\mathbb{N},$ as given in Corollary 3.2. We provide also concrete examples where the sufficient conditions of Theorem 3.1 and Corollary 3.2 in this paper are satisfied, but where the sufficient conditions of [24, Proposition 3.3] are not satisfied, illustrating thus the difference between (ordinary) topological transitivity and $\mathcal{F}$-semi-transitivity of cosine operator functions.
After pioneering papers by Salas on the dynamics on weighted shifts, several papers on the dynamics of various kinds of generalized weighted shifts have been published, see [26, 27, 41]. In particular, the dynamics of generalized bilateral weighted shifts on the standard Hilbert module over the C*-algebra of compact operators on a separable Hilbert space has been studied in [27], however, the concepts such as supercyclicity and disjoint supercyclicity have not been considered in [27]. This motivated us to study disjoint supercyclicity and related concepts for the generalized weighted shift operator considered in [27]. In Section 4 in this paper, we present first the concept of disjoint $\mathcal{F}$-semi-transitivity as a generalization of disjoint topological transitivity for supercyclicity introduced in [32], and then, as a continuation and extension of the results in [27], in Theorem 4.2 we characterize disjoint $\mathcal{F}$-semi-transitive generalized weighted bilateral shift operators on the standard Hilbert module over the C*-algebra of compact operators on a separable Hilbert space in the case when the corresponding $\mathcal{F}$-family is finitely invariant. Once again it turns out that a special case of this theory is the case when the $\mathcal{F}$-family under consideration is the family of all infinite subsets of $\mathbb{N}.$ This special case is treated in Corollary 4.4 and Proposition 4.6 in this paper. Also, we give concrete examples of $\mathcal{F}$-semi-transitive generalized weighted bilateral shift operators (Example 5.10) where we also illustrate the difference between $\mathcal{F}$-semi-transitive (and thus also supercyclic) generalized weighted bilateral shift operators on Hilbert modules and supercyclic classical weighted bilateral shifts on the sequence space $\ell^{2}.$ In addition, we introduce a new concept of $\mathcal{F}$-C*-semi-transitivity for generalized bilateral shift operators on the standard Hilbert C*-module and we characterize disjoint $\mathcal{F}$-C*-semi-transitive generalized weighted bilateral shifts. By our very definition, $\mathcal{F}$-semi-transitivity implies $\mathcal{F}$-C*-semi-transitivity, however, the converse is not true in general, as illustrated in Example 4.9 in this paper.
Next, hypercyclicity and supercyclicity of a composition of a left and a right multiplier on the space of compact operators have been studied in for instance [18, 35]. Motivated by those results, in Proposition 4.10 and Corollary 4.11 we provide sufficient conditions for disjoint $\mathcal{F}$-semi-transitivity of compositions of right multipliers and generalized weighted bilateral shift operators (where the corresponding weights are multiplied from the left) on the standard Hilbert module over the C*-algebra of compact operators on a separable Hilbert space and we illustrate these results with concrete example where these sufficient conditions are satisfied (Example 4.12). Finally, we conclude this paper with a remark (Remark 4.13) where we explain that our results from Section 4 remain valid also if we instead of a separable Hilbert space consider any standard Hilbert module over a unital C*-algebra.
Section 5 provides algebraic generalization of Theorem 4.2 in this paper to the case of more general non-unital Banach algebras.
2. Preliminaries
Throughout this paper, we let $\Omega$ be a locally compact non-compact Hausdorff space and $\alpha$ be an a homeomorphism of $\Omega.$ As usual, $C_{0}(\Omega)$ denotes the space of all continuous functions on $\Omega$ vanishing at infinity, $C_{b}(\Omega)$ denotes the space of bounded, continuous functions on $\Omega$, whereas $C_{c}(\Omega)$ stands for the set of all continuous, compactly supported functions on $\Omega$. Both $C_{0}(\Omega)$ and $C_{b}(\Omega)$ are equipped with the supremum norm. Moreover, we let $w$ be a positive bounded function on $\Omega$ such that also $w^{-1}:=\frac{1}{w}\in C_{b}(\Omega)$ and we put then $T_{\alpha,w}$ to be the weighted composition operator on $C_{0}(\Omega)$ with respect to $\alpha$ and $w,$ that is $T_{\alpha,w}(f)=w\cdot(f\circ\alpha)$ for all $f\in C_{0}(\Omega).$ Easily, one can see that by the above assumptions $T_{\alpha,w}$ is well-defined and $\|T_{\alpha,w}\|\leq\|w\|_{\sup}$. Since $\frac{1}{w}$ is also bounded, then $T_{\alpha,w}$ is invertible and we have
$T_{\alpha,w}^{-1}f=\frac{f\circ\alpha^{-1}}{w\circ\alpha^{-1}},\quad(f\in C_{0}(\Omega)).$
Simply we denote $S_{\alpha,w}:=T_{\alpha,w}^{-1}$.
By some calculation one can see that for each $n\in\mathbb{N}$ and $f\in C_{0}(\Omega)$,
$T^{n}_{\alpha,w}f=\left(\prod_{j=0}^{n-1}(w\circ\alpha^{j})\right)\cdot(f\circ\alpha^{n})\text{ (1)}$
and
$S^{n}_{\alpha,w}f=\left(\prod_{j=1}^{n}(w\circ\alpha^{-j})\right)^{-1}\cdot(f\circ\alpha^{-n}).\text{ (2)}$
The adjoint $T_{\alpha,w}^{\ast}$ is a bounded operator on $M(\Omega)$ where $M(\Omega)$ stands for the Banach space of all Radon measures on $\Omega$ equipped with the total variation norm. It is straightforward to check that
$T_{\alpha,w}^{\ast}(\mu)(E)=\int_{E}w\circ\alpha^{-1}\,d\mu\ \circ\alpha^{-1}$
for every $\mu\in M(\Omega)$ and every measurable subset $E$ od $\Omega.$ Here $\mu\ \circ\alpha^{-1}(E)=\mu(\alpha^{-1}(E))$ for every $\mu\in M(\Omega)$ and every measurable subset $E$ od $\Omega.$ By $(1)$ and $(2)$ it follows then that for every $n\in\mathbb{N},$ $\mu\in M(\Omega)$ and a Borel measurable subset $E\subseteq\Omega$ we have that
${T_{\alpha,w}^{\ast n}}(\mu)(E)=\int_{E}\prod_{j=0}^{n-1}w\circ\alpha^{j-n}\,d\mu\ \circ\alpha^{-n}$
and
${S_{\alpha,w}^{\ast n}}(\mu)(E)=\int_{E}\prod_{j=1}^{n}(w\circ\alpha^{n-j})^{-1}\,d\mu\ \circ\alpha^{n}.$
For every Radon measure $\mu$ on $\Omega,$ we let as usual $|\mu|$ denote the total variation of $\mu.$
In this paper, $\mathcal{F}$ represents a Furstenberg family, which is a specific collection of subsets satisfying the following conditions:
  1. (1)
    The empty set $\emptyset$ is not an element of $\mathcal{F}$.
  2. (2)
    If $A$ belongs to $\mathcal{F}$ and $B$ contains $A$, then $B$ also belongs to $\mathcal{F}$.
3. $\mathcal{F}$-semi-transitive cosine operator functions
Assume that $S$ is a set and $\mathcal{F}$ is a proper Furstenberg family of subsets of $S.$
Let $(\alpha_{z})_{z\in\in\mathcal{S}}$ be a family of homeomorphisms of $\Omega$ and $(w_{z})_{z\in\in\mathcal{S}}$ be a family of positive bounded continuous functions on $\Omega$ such that $\frac{1}{w_{z}}$ is also bounded for all $z\in\mathcal{S}$. For each $z\in\mathcal{S}$, we let $T_{z}^{\ast}:=T_{\alpha_{z},w_{z}}^{\ast}$ and let $C_{z}^{\ast}:=\frac{1}{2}(T_{z}^{\ast}+S_{z}^{\ast})$. We will say that $\{C_{z}^{\ast}\}_{z\in\mathcal{S}}$ is $\mathcal{F}$-semi-transitive if for every non-empty open subsets $U_{1}$ and $U_{2}$ of $M(\Omega)$, the set
$\{z\in\mathcal{S}\text{ : }\exists\lambda_{z}\in\mathbb{R}^{+}\text{ such that }\lambda_{z}C_{z}^{\ast}\ (U_{1})\cap U_{2}\neq\emptyset\}$
is an element of $\mathcal{F}$.
The next theorem gives some sufficient conditions for the family $\{C_{z}^{\ast}\}_{z\in\mathcal{S}}$ to be $\mathcal{F}$-semi-transitive.
Theorem 3.1.
We have $(ii)\Rightarrow(i)$.
  1. $(i)$
    $\{C_{z}^{\ast}\}_{z\in\mathcal{S}}$ is $\mathcal{F}$-semi-transitive;
  2. $(ii)$
    for each compact subset $K$ of $\Omega$, each $\varepsilon>0$, and any two non-zero measures $\mu$ $\nu\in M(\Omega)$ with $|\mu|(K^{c})=|\nu|(K^{c})=0$, there exists some $F\in\mathcal{F}$ and collections of Borel subsets $\{A_{z}\}_{z\in F}$, $\{D_{z}\}_{z\in F}$$\{E_{z}\}_{z\in F}$ of $K$ such that
    $|\mu|(A_{z})<\varepsilon\text{, }|\nu|(A_{z})<\varepsilon\text{, }D_{z}\cap E_{z}=\emptyset\text{, }A^{c}_{z}\cap K=D_{z}\cup E_{z},$
    for all $z\in F$, and
    $\left(\sup_{t\in A^{c}_{z}\cap K}w_{z}(t)\right)\cdot\left(\sup_{t\in D_{z}}w_{z}(t)\right)<\varepsilon,$
    $\left(\sup_{t\in A^{c}_{z}\cap K}w_{z}(t)\right)\cdot\left(\sup_{t\in E_{z}}\left(w_{z}\circ\alpha^{-1}\right)^{-1}(t)\right)<\varepsilon,$
    $\left(\sup_{t\in A^{c}_{z}\cap K}\left(w_{z}\circ\alpha^{-1}\right)^{-1}(t)\right)\cdot\left(\sup_{t\in D_{z}}w_{z}(t)\right)<\varepsilon,$
    $\left(\sup_{t\in A^{c}_{z}\cap K}\left(w_{z}\circ\alpha^{-1}\right)^{-1}(t)\right)\cdot\left(\sup_{t\in E_{z}}\left(w_{z}\circ\alpha^{-1}\right)^{-1}(t)\right)<\varepsilon,$
    $\sup_{t\in D_{z}}\left(\prod_{j=0}^{1}\left(w_{z}\circ\alpha_{z}^{j}\right)(t)\right)<\varepsilon,$
    $\sup_{t\in E_{z}}\left(\prod_{j=1}^{2}\left(w_{z}\circ\alpha_{z}^{-j}\right)^{-1}(t)\right)<\varepsilon\ \ \text{ for all }z\in F.$
Proof.
Let $U_{1}$ and $U_{2}$ be two open non-empty subsets of $M(\Omega)$. Then $U_{1}\setminus\{0\}$ and $U_{2}\setminus\{0\}$ are also open and non-empty subsets of $M(\Omega)$. As in the proof of [24, Proposition 3.1] part $(ii)\Rightarrow(i)$, we can find two measures $\mu\in U_{1}\setminus\{0\}$, $\nu\in U_{2}\setminus\{0\}$, and a compact subset $K$ of $\Omega$ such that $|\mu|(K^{c})=|\nu|(K^{c})=0$. Also, we can find some $\delta>0$ such that $B(\mu,\delta)\subset U_{1}\setminus\{0\}$ and $B(\nu,\delta)\subset U_{2}\setminus\{0\}$. Choose $F\in\mathcal{F}$ and collections of Borel subsets $\{A_{z}\}_{z\in F}$, $\{D_{z}\}_{z\in F}$$\{E_{z}\}_{z\in F}$ satisfying the assumptions of $(ii)$ with respect to $\mu$, $\nu$, and
$\varepsilon:=\min\left\{\frac{\delta}{4}\text{, }\frac{\delta^{2}}{64\|\mu\|\|\nu\|}\text{, }\frac{\delta}{8\|\nu\|}\right\}.$
For each $z\in F$, let $\tilde{\mu_{z}}$, $\tilde{\nu_{z}}$, $\tilde{\tilde{\nu_{z}}}$ be the elements of $M(\Omega)$ given by $\tilde{\mu_{z}}(B)=\mu(B\cap A_{z}^{c}\cap K)$, $\tilde{\nu_{z}}(B)=\nu(B\cap D_{z})$, and $\tilde{\tilde{\nu_{z}}}(B)=\nu(B\cap E_{z})$ for every $z\in F$ and any Borel subset $B$ of $\Omega$. Then
$|\mu-\tilde{\mu_{z}}|(\Omega)=|\mu|(A_{z})<\varepsilon<\delta\ \ \text{ and }\ \ |\mu-\tilde{\nu_{z}}-\tilde{\tilde{\nu_{z}}}|(\Omega)=|\nu|(A_{z})<\varepsilon<\delta,$
[…]
$\|T^{\ast}_{z}(\tilde{\mu_{z}})\|+\|S^{\ast}_{z}(\tilde{\mu_{z}})\|>0\ \ \text{ and }\|T^{\ast}_{z}(\tilde{\nu_{z}})\|+\|S^{\ast}_{z}(\tilde{\tilde{\nu_{z}}})\|>0.$
For each $z\in F$, put
[…]
Then
$\|\phi_{z}-\mu\|\leq\|\mu-\tilde{\mu_{z}}\|+2\sqrt{\|T^{\ast}_{z}(\tilde{\mu_{z}})\|+\|S^{\ast}_{z}(\tilde{\mu_{z}})\|}\sqrt{\|T^{\ast}_{z}(\tilde{\nu_{z}})\|+\|S^{\ast}_{z}(\tilde{\tilde{\nu_{z}}})\|}$
and
$\left\|\frac{\sqrt{\|T^{\ast}_{z}(\tilde{\nu_{z}})\|+\|S^{\ast}_{z}(\tilde{\tilde{\nu_{z}}})\|}}{\sqrt{\|T^{\ast}_{z}(\tilde{\mu_{z}})\|+\|S^{\ast}_{z}(\tilde{\mu_{z}})\|}}C^{\ast}_{z}(\phi_{z})-\nu\right\|$ $\leq\frac{1}{2}\sqrt{\|T^{\ast}_{z}(\tilde{\nu_{z}})\|+\|S^{\ast}_{z}(\tilde{\tilde{\nu_{z}}})\|}\ \sqrt{\|T^{\ast}_{z}(\tilde{\mu_{z}})\|+\|S^{\ast}_{z}(\tilde{\mu_{z}})\|}$
$+\|(T^{\ast}_{z})^{2}(\tilde{\nu_{z}})\|+\|(S^{\ast}_{z})^{2}(\tilde{\tilde{\nu_{z}}})\|+\|\tilde{\nu_{z}}+\tilde{\tilde{\nu_{z}}}-\nu\|,$
for all $z\in F$. Now, by exactly the same arguments as in the proof of [24, Proposition 3.1] part $(ii)\Rightarrow(i)$, one can show that for each $z\in F$ we have
$\left\|T^{\ast}_{z}(\tilde{\mu_{z}})\right\|\leq\sup_{t\in A^{c}_{z}\cap K}\left\{w_{z}(t)\|\mu\|\right\}\text{,}$
$\left\|S^{\ast}_{z}(\tilde{\mu_{z}})\right\|\leq\sup_{t\in A^{c}_{z}\cap K}\left\{(w_{z}\circ\alpha_{z}^{-1})^{-1}(t)\|\mu\|\right\}\text{,}$
$\left\|T^{\ast}_{z}(\tilde{\nu_{z}})\right\|\leq\sup_{t\in D_{z}}\left\{w_{z}(t)\|\nu\|\right\}\text{,}$
$\left\|S^{\ast}_{z}(\tilde{\tilde{\nu_{z}}})\right\|\leq\sup_{t\in E_{z}}\left\{(w_{z}\circ\alpha_{z}^{-1})^{-1}(t)\|\nu\|\right\}\text{,}$
$\left\|(T^{\ast}_{z})^{2}(\tilde{\nu_{z}})\right\|\leq\sup_{t\in D_{z}}\left\{\prod_{j=0}^{1}(w_{z}\circ\alpha_{z}^{j})(t)\|\nu\|\right\}\text{,}$
and
$\left\|(S^{\ast}_{z})^{2}(\tilde{\tilde{\nu_{z}}})\right\|\leq\sup_{t\in E_{z}}\left\{\prod_{j=1}^{2}(w_{z}\circ\alpha_{z}^{-j})^{-1}(t)\|\nu\|\right\}\text{.}$
Utilizing this fact together with assumptions $(ii)$ w. r. to $\varepsilon=\min\left\{\frac{\delta}{4}\text{, }\frac{\delta^{2}}{64\|\mu\|\|\nu\|}\text{, }\frac{\delta}{8\|\nu\|}\right\}$ and the fact that $B(\mu,\delta)\subset U_{1}$ and $B(\nu,\delta)\subset U_{2},$ we deduce that
$\frac{\sqrt{\|T^{\ast}_{z}(\tilde{\nu_{z}})\|+\|S^{\ast}_{z}(\tilde{\tilde{\nu_{z}}})\|}}{\sqrt{\|T^{\ast}_{z}(\tilde{\mu_{z}})\|+\|S^{\ast}_{z}(\tilde{\mu_{z}})\|}}C^{\ast}_{z}(\phi_{z})\in U_{1}$
for all $z\in F$. Hence
$F\subseteq\{z\in\mathcal{S}\text{ : }\exists\lambda_{z}\in\mathbb{R}^{+}\text{ such that }\lambda_{z}C^{\ast}(U_{1})\cap U_{2}\neq\emptyset\}.$
Since $\mathcal{F}$ is Furstenberg , we obtain that
$\{z\in\mathcal{S}\text{ : }\exists\lambda_{z}\in\mathbb{R}^{+}\text{ such that }\lambda_{z}C^{\ast}(U_{1})\cap U_{2}\neq\emptyset\}\in\mathcal{F}.$
Hence, $\{C_{z}^{\ast}\}_{z\in\mathcal{S}}$ is $\mathcal{F}$-semi-transitive since $U_{1}$ and $U_{2}$ were arbitrary. ∎
Corollary 3.2.
Let $w$ be a continuous positive function on $\Omega$ such that $w$ and $\frac{1}{w}$ are bounded and let $\alpha$ be a homeomorphism of $\Omega$. Suppose that for each compact subset $K$ of $\Omega$ it holds that
$\lim_{n\longrightarrow\infty}\left(\sup_{t\in K}\prod_{j=0}^{n-1}\left(w\circ\alpha^{j}\right)(t)\right)\left(\sup_{t\in K}\prod_{j=1}^{n}\left(w\circ\alpha^{-j}\right)^{-1}(t)\right)=0$
and either
$\lim_{n\longrightarrow\infty}\left(\sup_{t\in K}\prod_{j=0}^{2n-1}\left(w\circ\alpha^{j}\right)(t)\right)=0$
or
$\lim_{n\longrightarrow\infty}\left(\sup_{t\in K}\prod_{j=1}^{2n}\left(w\circ\alpha^{-j}\right)^{-1}(t)\right)=0.$
Then, for each non-empty open subsets $U_{1}$ and $U_{2}$ of $M(\Omega)$, there exists some $N\in\mathbb{N}$ and some sequence $\{\lambda_{n}\}_{n=N}^{\infty}$ of positive numbers such that $\lambda_{n}C_{n}^{\ast}(U_{1})\cap U_{2}\neq\emptyset\text{ for all }n\geq N,$ where $C_{n}^{\ast}=\frac{1}{2}\left(T^{\ast n}_{\alpha,w}+S^{\ast n}_{\alpha,w}\right)$.
Proof.
Let $S=\mathbb{N}$ and $\mathcal{F}$ be the family of all infinite subset of $\mathbb{N}$. For each $n\in\mathbb{N}$, we let $\alpha_{n}:=\alpha^{n}$ and $w_{n}:=\prod_{j=0}^{n-1}\left(w\circ\alpha^{j}\right)$. Then we consider the case when $A_{n}=E_{n}=\emptyset$ and $D_{n}=K$ (such that $\tilde{\tilde{\nu_{n}}}=0$) for all $n\in\mathbb{N}$, and the case when $A_{n}=D_{n}=\emptyset$ and $E_{n}=K$ for all $n\in\mathbb{N}$, (such that $\tilde{\nu_{n}}=0$ for all $n\in\mathbb{N}$). By adapting the proof of Theorem 3.1 for these special cases, we can deduce the statement of corollary. ∎
Now we will provide a concrete example where the conditions of Corollary 3.2 are satisfied.
Example 3.3.
Let $X=\mathbb{R}$ and $M,\delta$ be positive constants such that $M\geq 2+2\delta$ and $\delta\geq 1.$ Put $\alpha$ to be the function on $\mathbb{R}$ given by $\alpha(t)=t+1$ for all $t\in\mathbb{R}.$ Set
$w(t)=\begin{cases}M\text{ for }t\leq-1,\\ M+\frac{t+1}{2}(1+\delta-M)\text{ for }t\in[-1,1],\\ 1+\delta\text{ for }t\geq 1.\\ \end{cases}.$
By some calculations one can check that for each compact subset $K$ of $\Omega$ it holds that
$\lim_{n\longrightarrow\infty}\left(\sup_{t\in K}\prod_{j=0}^{n-1}\left(w\circ\alpha^{j}\right)(t)\right)\left(\sup_{t\in K}\prod_{j=1}^{n}\left(w\circ\alpha^{-j}\right)^{-1}(t)\right)=0$
and
$\lim_{n\longrightarrow\infty}\left(\sup_{t\in K}\prod_{j=1}^{2n}\left(w\circ\alpha^{-j}\right)^{-1}(t)\right)=0,$
so the conditions of Corollary 3.2 are satisfied. However, it is not hard to check by some calculations that for every strictly increasing sequence $\{n_{k}\}\subseteq\mathbb{N}$ and every $t\in\mathbb{R}$ it holds that
$\lim_{k\rightarrow\infty}\text{ }\prod_{j=0}^{n_{k}-1}(w\circ\alpha^{j})(t)=\infty,$
hence the conditions of [24, Proposition 3.3] are not satisfied. This illustrates the difference between the sufficient conditions of Corollary 3.2 and the sufficient conditions of [24, Proposition 3.3].
4. Disjoint $\mathcal{F}$-semi-transitive generalized weighted shifts
We start with the following definition.
Definition 4.1.
Let $\mathcal{S}$ be a set, $\mathcal{F}$ be a family of subsets of $\mathcal{S}.$ and $\{T_{t,1}\}_{t\in\mathcal{S}},\dots,\\ \{T_{t,N}\}_{t\in\mathcal{S}}$ be families of bounded linear operators on a Banach space $X$. We say that $\{T_{t,1}\}_{t\in\mathcal{S}},\dots,\{T_{t,N}\}_{t\in\mathcal{S}}$ are disjoint $\mathcal{F}$-semi-transitive, or shortly $d\mathcal{F}$-semi-transitive, if for every collection of non-empty open subsets $\mathcal{O},V_{1},\dots,V_{N}$ of $X$, there exists some $F\in\mathcal{F}$ such that for all $t\in F$ there exists some $\lambda_{t}\in\mathbb{R}^{+}$ satisfying that
$\mathcal{O}\cap\lambda_{t}T_{t,1}^{-1}(V_{1})\cap\dots\cap\lambda_{t}T_{t,N}^{-1}(V_{N})\neq\varnothing.$
We recall also that the families $\{T_{t,1}\}_{t\in S},\dots,\{T_{t,N}\}_{t\in S}$ are said to be disjoint $\mathcal{F}$-transitive, or shortly $d\mathcal{F}$-transitive, if for every collection of non-empty open subsets $\mathcal{O},V_{1},\dots,V_{N}$ of $X$, there exists some $F\in\mathcal{F}$ such that for all $t\in F$ we have
$\mathcal{O}\cap T_{t,1}^{-1}(V_{1})\cap\dots\cap T_{t,N}^{-1}(V_{N})\neq\varnothing.$
In this section, for a separable Hilbert space $H$ we let $B(H)$ be the space of all bounded linear operators on $H$, and $\mathcal{C}$ be the $C^{*}$-algebra of compact operators on $H$. For an orthonormal basis $\{e_{j}\}_{j\in\mathbb{Z}}$ for $H$, we let for each $m\in\mathbb{N}$, $P_{m}$ be the orthogonal projection onto $\text{Span}\{e_{-m},\dots,e_{m}\}$. We will denote by $\ell_{2}(\mathcal{C})$ the standard (right) Hilbert module over $\mathcal{C}$, see [33, Example 1.3.5]. Notice that $\ell_{2}(\mathcal{C})$ is a Banach algebra. Indeed, we can define multiplication on $\ell_{2}(\mathcal{C})$ as pointwise multiplication, i.e., if $\{x_{j}\}_{j\in\mathbb{Z}},\{y_{j}\}_{j\in\mathbb{Z}}\in\ell_{2}(\mathcal{C})$, then
$\{x_{j}\}_{j\in\mathbb{Z}}\cdot\{y_{j}\}_{j\in\mathbb{Z}}=\{x_{j}y_{j}\}_{j\in\mathbb{Z}}.$
To see that $\{x_{j}y_{j}\}_{j\in\mathbb{Z}}$ belongs to $\ell_{2}(\mathcal{C})$, it suffices to observe that
$\sum_{j}y_{j}^{*}x_{j}^{*}x_{j}y_{j}\leq\sum_{j}y_{j}^{*}\|x_{j}\|^{2}y_{j}\leq\|x\|^{2}\sum_{j}y_{j}^{*}y_{j}\leq\parallel\{x_{j}\}_{j\in\mathbb{Z}}\parallel^{2}\sum y_{j}^{*}y_{j}.$
For each $m,J\in\mathbb{N},$ we let $\tilde{p}_{J,m}\in\ell_{2}(\mathcal{C})$ be given by
[…]
As shown in the proof of [27, Proposition 3.1], the set $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ forms a left approximate unit for $\ell_{2}(\mathcal{C})$.
For $N\in\mathbb{N},$ let $r_{1},\dots,r_{N}\in\mathbb{N}$ with $r_{1}<r_{2}<\dots<r_{N}$. Further, for each $n\in\mathbb{N}$ and $s\in\{1,\dots,N\},$ let $\{W_{n,j}^{(s)}\}_{j\in\mathbb{Z}}$ be a family of operators in $B(H)$ which is uniformly bounded in norm and such that each $W_{n,j}^{(s)}$ has a bounded inverse. Also, assume that $\mathcal{U}_{1},\dots,\mathcal{U}_{N}$ are unitary operators on $H$. Then, for each $n\in\mathbb{N}$ and $s\in\{1,\dots,N\},$ we let $T_{n,s}$ be the operator on $\ell_{2}(\mathcal{\mathcal{C}})$ defined as
$\big(T_{n,s}(\{x_{i}\}_{i\in\mathbb{Z}})\big)=\{W_{n,j}^{(s)}\,x_{j-r_{s}n}\mathcal{U}_{s}^{r_{s}n}\}_{j\in\mathbb{Z}}.$
By [27], $T_{n,s}$ is a bounded linear operator on $\ell_{2}(\mathcal{\mathcal{C}})$ for each $n\in\mathbb{N}$ and $s\in\{1,\dots,N\}$.
In what follows, we will assume that $\mathcal{F}$ is a family of subsets of $\mathbb{N}$ which is finitely invariant, that is, for each $n\in\mathbb{N}$ and $F\in\mathcal{F}$ it holds that $F\setminus\{1,\dots,n\}\in\mathcal{F}$ also.
Now we are ready to give the first result of this section.
Theorem 4.2.
Under the above notation and assumptions, the following statements are equivalent.
(1) The families of operators $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
(2) For every $J,m\in\mathbb{N}$ and $\varepsilon>0$, there exists some $F\in\mathcal{F}$ such that for each $j\in\mathbb{Z}$ with $J\leq j\leq J$ we can find families $\{D_{n,j}\}_{n\in F},\quad\{G_{n,j}^{(1)}\}_{n\in F},\dots,\{G_{n,j}^{(N)}\}_{n\in F}$ of compact operators on $H$ satisfying
$\|D_{n,j}-P_{m}\|<\varepsilon,\quad\|G_{n,j}^{(s)}-P_{m}\|<\varepsilon,$
and
$\|W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|\,\|W_{n,j}^{(s)\,-1}G_{n,j}^{(s)}\|<\varepsilon^{2}\quad\text{for all }n\in F\text{ and }\ell,s\in\{1,\dots,N\},$
and satisfying in addition that for each distinct $s,\ell\in\{1,\dots,N\}$ and all $n\in F$
$\|W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)-1}\,G_{n,j}^{(s)}\|<\varepsilon.$
Proof.
$(1)\Rightarrow(2)$ Let $J,m\in\mathbb{N}$ be given. Since the families $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ are disjoint $\mathcal{F}$-semi-transitive and $\mathcal{F}$ is finitely invariant. given $\varepsilon>0$, we can find some $F\subseteq\mathcal{F}$ with $\{1,\dots,2J\}\cap F=\varnothing$ such that for each $n\in F$, there exists some $x_{n}\in\ell_{2}(\mathcal{C})$ and some $\lambda_{n}\in\mathbb{R}^{+}$ with
$\|(x_{n})-\tilde{p}_{J,m}\|<\varepsilon,$
$\|\lambda_{n}(T_{n,\ell}x_{n})-\tilde{p}_{J,m}\|<\varepsilon\text{ for all }\ell\in\{1,\dots,N\}.$
This gives further that
$\|(x_{n})_{j}-P_{m}\|<\varepsilon,$
and
$\|\lambda_{n}W_{n,j+r_{\ell}n}^{(\ell)}(x_{n})_{j}\,\mathcal{U}_{\ell}^{r_{\ell}n}\|<\varepsilon.$
for all $j\in\mathbb{Z}$ with $J\leq j\leq J$ and $\ell\in\{1,\dots,N\}$, because
$(T_{n,\ell}(x_{n}))_{j+r_{\ell}n}=W_{n,j+r_{\ell}n}^{(\ell)}\,(x_{n})_{j}\,\mathcal{U}_{\ell}^{r_{\ell}n},$
and $j+r_{\ell}n>J$ for all $j\in\mathbb{Z}$ with $J\leq j\leq J$ and $n\in F$.
Hence
$|\lambda_{n}|\,\|W_{n,j+r_{\ell}n}^{(\ell)}(x_{n})_{j}\|=\|\lambda_{n}W_{n,j+r_{\ell}n}^{(\ell)}(x_{n})_{j}\|=$
$\|\lambda_{n}W_{n,j+r_{\ell}n}^{(\ell)}(x_{n})_{j}\,\mathcal{U}_{\ell}^{r_{\ell}n}\|<\varepsilon$
for all $n\in F$, $j\in\mathbb{Z}$ with $J\leq j\leq J$ and $\ell\in\{1,\dots,N\}$. Further, since
$\|\lambda_{n}(T_{n,s}\,x_{n})_{j}-P_{m}\|<\varepsilon\quad\text{for all }j\in\mathbb{Z}\text{ with }-J\leq j\leq J\text{ and }s\in\{1,\dots,N\},$
it follows that
$\|\lambda_{n}W_{n,j}^{(s)}(x_{n})_{j-r_{s}n}\,\mathcal{U}_{s}^{r_{s}n}-P_{m}\|<\varepsilon,$
however,
$\frac{1}{|\lambda_{n}|}\,\|W_{n,j}^{(s)-1}(\lambda_{n}W_{n,j}^{(s)}(x_{n})_{j-r_{s}n}\,\mathcal{U}_{s}^{r_{s}n})\|=\,\|W_{n,j}^{(s)-1}\frac{1}{\lambda_{n}}(\lambda_{n}W_{n,j}^{(s)}(x_{n})_{j-r_{s}n}\,\mathcal{U}_{s}^{r_{s}n})\|$
$\|(x_{n})_{j-r_{s}n}\,\mathcal{U}_{s}^{r_{s}n}\|=\|(x_{n})_{j-r_{s}n}\|\leq\|x_{n}-\tilde{p}_{J,m}\|<\varepsilon,$
once again because $j-r_{s}n<J-2J\leq-J.$ For each $n\in F,$ $\ell,s\in\{1,\dots,N\}$ and $j\in\mathbb{Z}$ with $J\leq j\leq J$, define
$D_{n,j}:=(x_{n})_{j}\text{ and }\quad G_{n,j}^{(s)}=\lambda_{n}W_{n,j}^{(s)}(x_{n})_{j-r_{s}n}\,\mathcal{U}_{s}^{r_{s}n}.$
Then
$\|D_{n,j}-P_{m}\|<\varepsilon,\quad\|G_{n,j}^{(s)}-P_{m}\|<\varepsilon,$
and
$|\lambda_{n}|\,\|W_{n,j+r_{\ell}n}^{(\ell)\,}D_{n,j}\|=\|\lambda_{n}W_{n,j+r_{\ell}n}^{(\ell)\,}D_{n,j}\|<\varepsilon,$
$\tfrac{1}{|\lambda_{n}|}\,\|W_{n,j}^{(s)\,-1}G_{n,j}^{(s)}\|=\|W_{n,j}^{(s)\,-1}\tfrac{1}{\lambda_{n}}G_{n,j}^{(s)}\|<\varepsilon,$
so
$\|W_{n,j+r_{\ell}n}^{(\ell)\,}D_{n,j}\|\|W_{n,j}^{(s)\,-1}G_{n,j}^{(s)}\|<\varepsilon^{2}.$
Finally, if $\ell,s\in\{1,\dots,N\}$ with $\ell\neq s$, then $|r_{\ell}-r_{s}|n>2J$ for all $n\in F$. Hence, either $j-r_{s}n+r_{\ell}n<-J$ or $j-r_{s}n+r_{\ell}n>J$ for any $j\in\mathbb{Z}$ with $J\leq j\leq J$. Therefore, $(\tilde{p}_{J,m})_{j-r_{s}n+r_{\ell}n}=0,$ hence
$\|W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)\,-1}\,G_{n,j}^{(s)}\|=$
$\|\lambda_{n}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)\,-1}\,\tfrac{1}{\lambda_{n}}G_{n,j}^{(s)}\|=\big\|\lambda_{n}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}(x_{n})_{j-r_{s}n}\mathcal{U}_{s}^{r_{s}n}\big\|$
$\big\|\lambda_{n}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}(x_{n})_{j-r_{s}n}\big\|=\big\|\lambda_{n}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}(x_{n})_{j-r_{s}n}\mathcal{U}_{\ell}^{r_{\ell}n}\big\|$
$\big\|\lambda_{n}(T_{n,\ell}(x_{n}))_{j-r_{s}n+r_{\ell}n}\big\|=\big\|(\lambda_{n}T_{n,\ell}(x_{n})-\tilde{p}_{J,m})_{j-r_{s}n+r_{\ell}n}\big\|<\varepsilon.$
for all $n\in F$, $j\in\mathbb{Z}$ with $J\leq j\leq J$, and $s,\ell\in\{1,\dots,N\}$ with $s\neq\ell$.
$(2)\Rightarrow(1)$ Let $\mathcal{O},V_{1},\dots,V_{N}$ be non-empty open subsets of $\ell_{2}(\mathcal{C})$. Then there exist some $J,m\in\mathbb{N}$ and some $x\in\mathcal{O}$, $y^{(1)}\in V_{1},\dots,y^{(N)}\in V_{N}$ such that $\tilde{p}_{J,m}x=x$ and $\tilde{p}_{J,m}y^{(\ell)}=y^{(\ell)}$ for each $\ell\in\{1,\dots,N\}$ since the set $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ forms a left approximate unit for $\ell_{2}(\mathcal{C})$.
For each $a\in\ell_{2}(\mathcal{C}),$ let $B_{\delta}(a)$ denote the open ball with centre in $a$ and radius $\delta$ . We can find some $\delta>0$ such that $B_{\delta}(x)\subseteq\mathcal{O}$ and $B_{\delta}(y^{(\ell)})\subseteq V_{\ell}$ for each $\ell\in\{1,\dots,N\}.$ Put $C_{y}=\sum_{s=1}^{N}\|y^{(s)}\|$ and
$\varepsilon:=\frac{\delta}{2JN(\|x\|+2C_{y}+\delta)}.$
Choose $F\in\mathcal{F}$ and families
$\{D_{n,j}\}_{n\in F},\quad\{G_{n,j}^{(1)}\}_{n\in F},\dots,\{G_{n,j}^{(N)}\}_{n\in F}$
for each $j\in\mathbb{Z}$ with $J\leq j\leq J$, satisfying the assumptions of (2) with respect to $J$, $m$, and $\varepsilon$.
For each $n\in F$, let $u_{n},v_{n}^{(1)},\dots,v_{n}^{(N)}\in\ell_{2}(\mathcal{C})$ be given by
$(u_{n})_{j}=\begin{cases}D_{n,j}\,x_{j},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else.}\end{cases}$
and
$(v_{n}^{(\ell)})_{j-r_{\ell}n}=\begin{cases}W_{n,j}^{(\ell)\,-1}\,G_{n,j}^{(\ell)}\,(y^{(\ell)})_{j}\,\mathcal{U}_{\ell}^{*r_{\ell}n},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else.}\end{cases}$
Since $\|D_{n,j}-P_{m}\|<\varepsilon$ and $\|G_{n,j}^{(\ell)}-P_{m}\|<\varepsilon<1$ for all $n\in F$, $\ell\in\{1,\dots,N\}$ and $j\in\mathbb{Z}$ with $J\leq j\leq J$, it follows that $D_{n,j}\neq 0$ and $G_{n,j}^{(\ell)}\neq 0,$ hence $W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\neq 0$, and $W_{n,j}^{(\ell)-1}G_{n,j}^{(\ell)}\neq 0$ for all $n\in F$, $\ell\in\{1,\dots,N\}$ and $j\in\mathbb{Z}$ with $J\leq j\leq J$. Now we notice that
$\|T_{n,\ell}(u_{n})\|\leq\sum_{j=-J}^{J}\|W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|\,\|x\|$
because
$(T_{n,\ell}(u_{n}))_{j+r_{\ell}n}=\begin{cases}W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\,x_{j}\,\mathcal{U}_{\ell}^{r_{\ell}n},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else}.\end{cases}$
Moreover, since for each distinct $s,\ell\in\{1,\dots,N\}$ and all $n\in F$,
$(T_{n,\ell}(v_{n}^{(s)}))_{j-r_{s}n+r_{\ell}n}=$
$\begin{cases}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)-1}\,G_{n,j}^{(s)}(y^{(s)})_{j}\,\mathcal{U}_{s}^{*\,r_{s}n}\,\mathcal{U}_{\ell}^{r_{\ell}n},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else.}\end{cases}$
it follows that
$\|T_{n,\ell}(v_{n}^{(s)})\|\leq\sum_{j=-J}^{J}\|W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)\,-1}\,G_{n,j}^{(s)}\|\,\|y^{(s)}\|.$
Finally, it is easy to see that by the construction of $v_{n}^{(\ell)}$, we have
$T_{n,\ell}(v_{n}^{(\ell)})_{j}=G_{n,j}^{(\ell)}(y^{(\ell)})_{j}\quad\text{if }-J\leq j\leq J,\quad\text{and }0\text{ else.}$
For each $n\in F$, set
$x_{n}:=u_{n}+\frac{\sqrt{\sum_{\ell=1}^{N}\sum_{j=-J}^{J}\|W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|}}{\sqrt{\sum_{\ell=1}^{N}\sum_{j=-J}^{J}\|W_{n,j}^{(\ell)\,-1}G_{n,j}^{(\ell)}\|}}\sum_{\ell=1}^{N}v_{n}^{(\ell)}.$
By the triangle inequality and the choice of $\varepsilon$, it is not hard to deduce by some calculations that
$x_{n}\in\mathcal{O}\quad\text{and}\quad\frac{\sqrt{\sum_{\ell=1}^{N}\sum_{j=-J}^{J}\|W_{n,j}^{(\ell)\,-1}G_{n,j}^{(\ell)}\|}}{\sqrt{\sum_{\ell=1}^{N}\sum_{j=-J}^{J}\|W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|}}T_{n,\ell}(x_{n})\in V_{\ell}$
for all $n\in F$ and $\ell\in\{1,\dots,N\}$. ∎
Next, let $\{S_{n}\}_{n\in\mathbb{N}}$ be a sequence of bounded invertible linear operators on $H$ and $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ be families of generalized weighted bilateral shifts on $\ell_{2}(\mathcal{C})$ as defined in the beginning of this section. For each $n\in F$ and $s\in\{1,\dots,N\}$ we let $S_{n}T_{n,s}$ be the operator on $\ell_{2}(\mathcal{C})$ given by
$\big(S_{n}T_{n,s}(\{x_{i}\}_{i\in\mathbb{Z}})\big):=\{S_{n}W_{n,j}^{(s)}\,x_{j-r_{s}n}\mathcal{U}_{s}^{r_{s}n}\}_{j\in\mathbb{Z}}.$
It is not hard to check that the operator $S_{n}T_{n,s}$ is a bounded linear operator on $\ell_{2}(\mathcal{C}).$ In fact, $S_{n}T_{n,s}$ is again a generalized weighted bilateral shift operator on $\ell_{2}(\mathcal{C})$ with respective weights $\{S_{n}W_{n,j}^{(s)}\}_{j\in\mathbb{Z}}$ for every $n\in F$ and $s\in\{1,\dots,N\}.$
We will say that $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ are disjoint $\mathcal{F}$-C*-semi-transitive , or shortly $d\mathcal{F}$-C*-semi-transitive , if for every collection of non-empty open subsets $\mathcal{O},V_{1},\dots,V_{N}$ of $X$, there exists some $F\in\mathcal{F}$ such that for all $n\in F$ there exists some invertible bounded linear operator $S_{n}$ on $H$ satisfying that
$\mathcal{O}\cap(S_{n}T_{n,1})^{-1}(V_{1})\cap\dots\cap(S_{n}T_{n,N})^{-1}(V_{N})\neq\varnothing.$
The next theorem can easily be deduced from Theorem 4.2.
Theorem 4.3.
Under the above notation and assumptions, the following statements are equivalent.
(1) The families of operators $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-C*-semi-transitive.
(2) For every $J,m\in\mathbb{N}$ and $\varepsilon>0$, there exists some $F\in\mathcal{F}$ and a sequence $\{S_{n}\}_{n\in F}$ of bounded invertible linear operators on $H$ such that for each $j\in\mathbb{Z}$ with $J\leq j\leq J$ we can find families $\{D_{n,j}\}_{n\in F},\quad\{G_{n,j}^{(1)}\}_{n\in F},\dots,\{G_{n,j}^{(N)}\}_{n\in F}$ of compact operators on $H$ satisfying for all $n\in F$ and $\ell,s\in\{1,\dots,N\}$ that
$\|D_{n,j}-P_{m}\|<\varepsilon,\quad\|G_{n,j}^{(s)}-P_{m}\|<\varepsilon,$
$\|S_{n}W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|<\varepsilon,\text{ }\|W_{n,j}^{(s)\,-1}S_{n}^{-1}G_{n,j}^{(s)}\|<\varepsilon,$
and satisfying in addition that for each distinct $s,\ell\in\{1,\dots,N\}$ and all $n\in F$
$\|S_{n}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)-1}S_{n}^{-1}\,G_{n,j}^{(s)}\|<\varepsilon.$
Proof.
For the proof of $(1)\Rightarrow(2),$ we can proceed in exactly the same way as in the proof of Theorem 4.2 part $(1)\Rightarrow(2)$ by considering $S_{n}$ and $S_{n}^{-1}$ instead of $\lambda_{n}$ and $\tfrac{1}{\lambda_{n}},$ respectively.
For the proof of $(2)\Rightarrow(1),$ we can also proceed similarly as in the proof of Theorem 4.2 part $(2)\Rightarrow(1).$ However, the only difference is that in this case we let now for each $n\in F$, $v_{n}^{(1)},\dots,v_{n}^{(N)}\in\ell_{2}(\mathcal{C})$ be given by
$(v_{n}^{(\ell)})_{j-r_{\ell}n}=\begin{cases}W_{n,j}^{(\ell)\,-1}\,S_{n}^{-1}G_{n,j}^{(\ell)}\,(y^{(\ell)})_{j}\,\mathcal{U}_{\ell}^{*r_{\ell}n},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else,}\end{cases}$
which clearly gives that
$\|(v_{n}^{(\ell)})\|\leq\sum_{j=-J}^{J}\|W_{n,j}^{(\ell)\,-1}\,S_{n}^{-1}G_{n,j}^{(\ell)}\|\|y^{(\ell)}\|.$
Then it suffices to also notice in this case that for each $n\in F$ and $\ell\in\{1,\dots,N\}$ it holds that
$\|S_{n}T_{n,\ell}(u_{n})\|\leq\sum_{j=-J}^{J}\|S_{n}W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|\,\|x\|$
(where $u_{n}$ is the same as in the proof of Theorem 4.2 in the part $(2)\Rightarrow(1)$) and and moreover, for each distinct $s,\ell\in\{1,\dots,N\}$ and $n\in F$ we have
$\|S_{n}T_{n,\ell}(v_{n}^{(s)})\|\leq\sum_{j=-J}^{J}\|S_{n}W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)\,-1}S_{n}^{-1}\,G_{n,j}^{(s)}\|\,\|y^{(s)}\|.$
Finally, by the construction of $v_{n}^{(\ell)}$, it follows that
$S_{n}T_{n,\ell}(v_{n}^{(\ell)})_{j}=G_{n,j}^{(\ell)}(y^{(\ell)})_{j}\quad\text{if }-J\leq j\leq J,\quad\text{and }0\text{ else,}$
for each $n\in F$ and $\ell\in\{1,\dots,N\}.$
For each $n\in F$, set
$x_{n}:=u_{n}+\sum_{\ell=1}^{N}v_{n}^{(\ell)}.$
By the triangle inequality and the choice of $\varepsilon$, it is not hard to deduce by some calculations that
$x_{n}\in\mathcal{O}\quad\text{and}\quad S_{n}T_{n,\ell}(x_{n})\in V_{\ell}$
for all $n\in F$ and $\ell\in\{1,\dots,N\}.$
Let now $\mathcal{F}$ be the family of all infinite subsets of $\mathbb{N}$ and
$\{W_{j}^{(1)}\}_{j\in\mathbb{Z}},\dots,\{W_{j}^{(N)}\}_{j\in\mathbb{Z}}\subset B(H)$
be such that each $W_{j}^{(\ell)}$ $(j\in\mathbb{Z},\,1\leq\ell\leq N)$ is invertible, and such that
$\|W_{j}^{(\ell)}\|\leq M\quad\text{for all }j\in\mathbb{Z},\,\ell\in\{1,\dots,N\},$
and some fixed $M>0$. For each $n\in\mathbb{N},i\in\mathbb{Z}$ and $\ell\in\{1,\dots,N\}$, we let $W_{n,i}^{(l)}\in B(H)$ be given by
$W_{n,i}^{(l)}=W_{i}^{(\ell)}W_{i-1}^{(\ell)}\cdots W_{i-r_{\ell}n+1}^{(\ell)}.$
Finally, for each $\ell\in\{1,\dots,N\}$ we let $\tilde{T}_{\ell}$ be the generalized bilateral weighted shift operator on $\ell_{2}(\mathcal{C})$ (introduced in [27]) with respect to the sequence $\{W_{j}^{(\ell)}\}_{j\in\mathbb{Z}}$ and unitary $\mathcal{U}_{\ell}$, i.e.
[…]
Then it is not hard to check that for each $\ell\in\{1,\dots,N\}$ it holds that
[…]
We will now present an application of Theorem 4.2 and Theorem 4.3 in this concrete case.
Corollary 4.4.
Under the above notation and assumptions, the following statements are equivalent.
(1) The sequences of operators $\{\tilde{T}_{1}^{r_{1}n}\}_{n\in\mathbb{N}},\dots,\{\tilde{T}_{N}^{r_{N}n}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
(2) For every $J,m\in\mathbb{N}$ there exist a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ and sequences of operators in $\mathcal{C}$
$\{D_{-J}^{(k)}\}_{k\in\mathbb{N}},\quad\{G_{-J,1}^{(k)}\}_{k\in\mathbb{N}},\dots,\{G_{-J,N}^{(k)}\}_{k\in\mathbb{N}},$
$\{D_{-J+1}^{(k)}\}_{k\in\mathbb{N}},\quad\{G_{-J+1,1}^{(k)}\}_{k\in\mathbb{N}},\dots,\{G_{-J+1,N}^{(k)}\}_{k\in\mathbb{N}},$
$\vdots$
$\{D_{J}^{(k)}\}_{k\in\mathbb{N}},\quad\{G_{J,1}^{(k)}\}_{k\in\mathbb{N}},\dots,\{G_{J,N}^{(k)}\}_{k\in\mathbb{N}},$
such that
$\lim_{k\to\infty}\|D_{j}^{(k)}-P_{m}\|=\lim_{k\to\infty}\|G_{j,\ell}^{(k)}-P_{m}\|=0,$
and
$\lim_{k\to\infty}\|W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}G_{j,s}^{(k)}\|\|W_{j+r_{\ell}n_{k}}^{(\ell)}\cdots W_{j+1}^{(\ell)}D_{j}^{(k)}\|=0$
for all $J\in\mathbb{Z}$ with $J\leq j\leq J$ and $\ell,s\in\{1,\dots,N\}$. Moreover, for each distinct $\ell,s\in\{1,\dots,N\}$ and all $j\in\mathbb{Z}$ with $J\leq j\leq J$, it holds that
$\lim_{k\to\infty}\|W_{j+r_{\ell}n_{k}-r_{s}n_{k}}^{(\ell)}\cdots W_{j-r_{s}n_{k}+1}^{(\ell)}W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}G_{j,s}^{(k)}\|=0.$
Corollary 4.5.
Under the above notation and assumptions, the following statements are equivalent.
(1) The sequences of operators $\{\tilde{T}_{1}^{r_{1}n}\}_{n\in\mathbb{N}},\dots,\{\tilde{T}_{N}^{r_{N}n}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-C*-semi-transitive.
(2) For every $J,m\in\mathbb{N}$ there exist a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N},$ a sequence $\{S_{n_{k}}\}_{k\in\mathbb{N}}$ of bounded invertible linear operators on $H$ and sequences of operators in $\mathcal{C}$
$\{D_{-J}^{(k)}\}_{k\in\mathbb{N}},\quad\{G_{-J,1}^{(k)}\}_{k\in\mathbb{N}},\dots,\{G_{-J,N}^{(k)}\}_{k\in\mathbb{N}},$
$\{D_{-J+1}^{(k)}\}_{k\in\mathbb{N}},\quad\{G_{-J+1,1}^{(k)}\}_{k\in\mathbb{N}},\dots,\{G_{-J+1,N}^{(k)}\}_{k\in\mathbb{N}},$
$\vdots$
$\{D_{J}^{(k)}\}_{k\in\mathbb{N}},\quad\{G_{J,1}^{(k)}\}_{k\in\mathbb{N}},\dots,\{G_{J,N}^{(k)}\}_{k\in\mathbb{N}},$
such that
$\lim_{k\to\infty}\|D_{j}^{(k)}-P_{m}\|=\lim_{k\to\infty}\|G_{j,\ell}^{(k)}-P_{m}\|=0,$
and
$\lim_{k\to\infty}\|W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}S_{n_{k}}^{-1}G_{j,s}^{(k)}\|=0,$
$\lim_{k\to\infty}\|S_{n_{k}}W_{j+r_{\ell}n_{k}}^{(\ell)}\cdots W_{j+1}^{(\ell)}D_{j}^{(k)}\|=0,$
for all $J\in\mathbb{Z}$ with $J\leq j\leq J$ and $\ell,s\in\{1,\dots,N\}$ and satisfying in addition that for each distinct $\ell,s\in\{1,\dots,N\}$ and all $j\in\mathbb{Z}$ with $J\leq j\leq J$, it holds that
$\lim_{k\to\infty}\|S_{n_{k}}W_{j+r_{\ell}n_{k}-r_{s}n_{k}}^{(\ell)}\cdots W_{j-r_{s}n_{k}+1}^{(\ell)}W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}S_{n_{k}}^{-1}G_{j,s}^{(k)}\|=0.$
In order to construct concrete examples satisfying the conditions in $(2)$ of Corollary 4.4 and Corollary 4.5, we need first the following technical results.
Proposition 4.6.
Under the above notation and assumptions, if for each $j\in\mathbb{Z}$ there exist dense subsets $H_{0}^{(j)},H_{1}^{(j)},\dots,H_{N}^{(j)}$ of $H$ and a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ such that for each $s,\ell\in\{1,\dots,N\}$ and every $x^{(j)}\in H_{0}^{(j)}$, $y_{s}^{(j)}\in H_{s}^{(j)}$ we have
$\lim_{k\to\infty}\big\|W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}\,y_{s}^{(j)}\big\|\big\|W_{j+r_{\ell}n_{k}}^{(\ell)}\cdots W_{j+1}^{(\ell)}\,x^{(j)}\big\|=0,$
and for each distinct $s,\ell\in\{1,\dots,N\}$ and every $y_{s}^{(j)}\in H_{s}^{(j)}$,we have
$\lim_{k\to\infty}\|W_{j+r_{\ell}n_{k}-r_{s}n_{k}}^{(\ell)}\cdots W_{j-r_{s}n_{k}+1}^{(\ell)}W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}y_{s}^{(j)}\|=0,$
then the sequences of operators $\{\tilde{T}_{1}^{r_{1}n}\}_{n\in\mathbb{N}},\dots,\{\tilde{T}_{N}^{r_{N}n}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
Proof.
The main idea in this proof is motivated by the proofs of [27, Theorem 3.2] and [28, Proposition 2.7]. Given $m\in\mathbb{N}$ and $j\in\mathbb{Z}$ with $J\leq j\leq J$, for every $\eta\in\{-m,\dots,m\}$ we can find sequences
$\{f_{i,j}^{(\eta)}\}_{i\in\mathbb{N}},\subseteq H_{0}^{(j)},\{h_{i,l,j}^{(\eta)}\}_{i\in\mathbb{N}}\subseteq H_{l}^{(j)}\text{ for each }\ell\in\{1,\dots,N\}$
such that
$f_{i,j}^{(\eta)}\to e_{\eta},\text{ and }h_{i,\ell,j}^{(\eta)}\to e_{\eta}\quad\text{as }i\to\infty,$
for all $\eta\in\{-m,\dots,m\}$ and $\ell\in\{1,\dots,N\}$. Since by the assumption we have for all $\ell,s\in\{1,\dots,N\}$ and $\eta\in\{-m,\dots,m\}$ that
$\lim_{k\to\infty}\big\|W_{j-r_{s}n_{k}+1}^{(s)^{-1}}\cdots W_{j}^{(s)^{-1}}h_{i,s,j}^{(\eta)}\big\|\big\|W_{j+r_{\ell}n_{k}}^{(\ell)}\cdots W_{j+1}^{(\ell)}f_{i,j}^{(\eta)}\big\|=0,$
and for each distinct $s,\ell\in\{1,\dots,N\}$ and $\eta\in\{-m,\dots,m\}$ we have
$\lim_{k\to\infty}\big\|W_{j+r_{\ell}n_{k}-r_{s}n_{k}}^{(\ell)}\dots W_{j-r_{\ell}n_{k}+1}^{(\ell)}W_{j-r_{s}n_{k}+1}^{(s)^{-1}}\dots W_{j}^{(s)^{-1}}\,h_{i,s,j}^{(\eta)}\big\|=0,$
we can inductively easily construct a subsequence $\{n_{k_{i}}\}_{i\in\mathbb{N}}$ such that
$\big\|W_{j-r_{s}n_{k_{i}}}^{(s)^{-1}}\cdots W_{j}^{(s)^{-1}}\,h_{i,s,j}^{(\eta)}\big\|\big\|W_{j+r_{\ell}n_{k_{i}}}^{(\ell)}\cdots W_{j+1}^{(\ell)}\,f_{i,j}^{(\eta)}\big\|<\tfrac{1}{4m^{2}4^{i}},$
and
$\big\|W_{j+r_{\ell}n_{k_{i}}-r_{s}n_{k_{i}}}^{(\ell)}\cdots W_{j-r_{s}n_{k_{i}}+1}^{(\ell)}W_{j-r_{s}n_{k_{i}}+1}^{(s)^{-1}}\cdots W_{j}^{(s)^{-1}}\,h_{i,s,j}^{(\eta)}\big\|<\tfrac{1}{2_{m}2^{i}}.$
For each $i\in\mathbb{N}$ and $\ell\in\{1,\dots,N\}$ we define operators $D_{j}^{(i)}$ and $G_{j,\ell}^{(i)}$ on $H$ by
$(D_{j}^{(i)})\,e_{\eta}=\begin{cases}f_{i,j}^{(\eta)},&\text{for }\eta\in\{-m,\dots,m\},\\[4.0pt] 0,&\text{else},\end{cases}\qquad(G_{j,\ell}^{(i)})\,e_{\eta}=\begin{cases}h_{i,\ell,j}^{(\eta)},&\text{for }\eta\in\{-m,\dots,m\},\\[4.0pt] 0,&\text{else}.\end{cases}$
By some calculations and arguments similar to those in the proof of [27, Theorem 3.2] and [28, Proposition 2.7], we can conclude that the sequences of operators
$\{D_{j}^{(i)}\}_{i\in\mathbb{N}},\quad\{G_{j,1}^{(i)}\}_{i\in\mathbb{N}},\dots,\{G_{j,N}^{(i)}\}_{i\in\mathbb{N}}$
satisfy the conditions in (2) of Corollary 4.4. for every $j\in\mathbb{Z}$ with $J\leq j\leq J$. ∎
Similarly, based on Theorem 4.3 and Corollary 4.5, we can prove the following proposition.
Proposition 4.7.
Under the above notation and assumptions, if for each $j\in\mathbb{Z}$ there exist dense subsets $H_{0}^{(j)},H_{1}^{(j)},\dots,H_{N}^{(j)}$ of $H,$ a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ and a sequence $\{S_{n_{k}}\}_{k\in\mathbb{N}}$ of bounded invertible linear operators on $H$ such that for each $s,\ell\in\{1,\dots,N\}$ and every $x^{(j)}\in H_{0}^{(j)}$, $y_{s}^{(j)}\in H_{s}^{(j)}$ we have
$\lim_{k\to\infty}\big\|W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}S_{n_{k}}^{-1}\,y_{s}^{(j)}\big\|=0,$
$\lim_{k\to\infty}\big\|S_{n_{k}}W_{j+r_{\ell}n_{k}}^{(\ell)}\cdots W_{j+1}^{(\ell)}\,x^{(j)}\big\|=0,$
and for each distinct $s,\ell\in\{1,\dots,N\}$ and every $y_{s}^{(j)}\in H_{s}^{(j)}$,we have
$\lim_{k\to\infty}\|S_{n_{k}}W_{j+r_{\ell}n_{k}-r_{s}n_{k}}^{(\ell)}\cdots W_{j-r_{s}n_{k}+1}^{(\ell)}W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}S_{n_{k}}^{-1}y_{s}^{(j)}\|=0,$
then the sequences of operators $\{\tilde{T}_{1}^{r_{1}n}\}_{n\in\mathbb{N}},\dots,\{\tilde{T}_{N}^{r_{N}n}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-C*-semi-transitive.
Thanks to Proposition 4.6 we will now give an example where we will show that it can happen that $\|W_{j}\|,\|W_{j}^{-1}\|\geq 1$ for all $j\in\mathbb{Z}$ and that at a same time the sequence of operators $\{\tilde{T}^{n}\}_{n\in\mathbb{N}}$ is $\mathcal{F}$-semi-transitive, which illustrates the important difference between $\mathcal{F}$-semi-transitivity (and thus supercyclicity) of generalized bilateral weighted shifts on Hilbert C*-modules and supercyclicity of the classical bilateral weighted shifts on the sequence space $\ell^{2}.$
Example 4.8.
Let $N=1,r_{1}=1$ and $H=L^{2}(\mathbb{R})$. As in [27, Example 3.3], choose some $\{w_{j}\}_{j\in\mathbb{Z}}\subseteq L^{\infty}(\mathbb{R})$ such that $w_{j}>0$, $w_{j}^{-1}\in L^{\infty}(\mathbb{R})$ and $\parallel w_{j}\parallel\leq M$ for all $j\in\mathbb{Z}$ and some $M>1$. Let $\{c_{j}\}_{j\in\mathbb{Z}}\subseteq\mathbb{R}^{+}$ be such that $c_{j}\geq C$ for all $j\in\mathbb{Z}$ and some $C>0$. For each $j\in\mathbb{Z}$ let $\alpha_{j}$ be the translation on $\mathbb{R}$ given by $\alpha_{j}(t)=t-c_{j}$ for all $t\in\mathbb{R}$. If there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}\leq 1-\varepsilon,w_{j}\chi_{(-\infty,0)}=1$ for all $j\geq 0$ and $w_{j}=1$ for all $j<0$; then by the same arguments as in [27, Example 3.3] one can show that the conditions of Proposition 4.6 are satisfied when $W_{j}(t)=w_{j}\cdot(f\circ\alpha_{j})$ for all $f\in L^{2}(\mathbb{R}),j\in\mathbb{Z}$ and $H_{0}^{(j)}=H_{1}^{(j)}=C_{c}(\mathbb{R})$ for all $j\in\mathbb{Z}.$ However, by the construction of the sequence $\{w_{j}\}$, it follows that $\|W_{j}\|,\|W_{j}^{-1}\|\geq 1$ for all $j\in\mathbb{Z}.$ On the other hand, if there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}=1,w_{j}\chi_{(-\infty,0)}\geq 1+\varepsilon$ for all $j<0$ and $w_{j}=1$ for all $j\geq 0,$ then again the conditions of Proposition 4.6 will be satisfied, however, in this case we would again have by the construction of the sequence $\{w_{j}\}$ that $\|W_{j}\|,\|W_{j}^{-1}\|\geq 1$ for all $j\in\mathbb{Z}.$
By our definition, $d\mathcal{F}$-semi-transitivity implies $d\mathcal{F}$-C*-semi-transitivity, however, the converse is not true in general, as the next example illustrates.
Example 4.9.
Let again $N=1,r_{1}=1,H=L^{2}(\mathbb{R})$ and $\alpha$ be the function on $\mathbb{R}$ given by $\alpha(t)=t-1$ for all $t\in\mathbb{R}.$ Let $v=2\chi_{(-\infty,0)}+\frac{1}{2}\chi_{[0,\infty)}$ and $\tilde{S}$ be the weighted composition operator on $H$ with respect to $v$ and $\alpha,$ that is $\tilde{S}(f)=v\cdot(f\circ\alpha)$ for all $f\in H.$ For each $n\in\mathbb{N},$ put $S_{n}:=\tilde{S}^{n}.$ If we let for each $j\in\mathbb{Z},$ $W_{j}$ be the identity operator on $H,$ then it is not hard to check that the conditions of Proposition 4.7 are satisfied with $H_{0}^{(j)}=H_{1}^{(j)}=C_{c}(\mathbb{R})$ for all $j\in\mathbb{Z}.$ Thus, the generalized non-weighted bilateral shift operator on $\ell_{2}(\mathcal{C})$ is $\mathcal{F}$-C*-semi-transitive, however, by Theorem 4.2 it can not be $\mathcal{F}$-semi-transitive on $\ell_{2}(\mathcal{C}).$
Hypercyclicity and supercyclicity of a composition of a left and a right multiplier on the space of compact operators have been studied in for instance [18, 35]. Motivated by those results, in the rest of this paper, we shall study the dynamics of a composition of a right multiplier and generalized weighted bilateral shift operator on $\ell_{2}(\mathcal{C})$ defined in the beginning of this section. To this end, let $\tilde{W_{1}},\dots,\tilde{W_{N}}$ be invertible, bounded linear operators on $H.$ For each $s\in\{1,\dots,N\}$ and $n\in\mathbb{N},$ we let $R_{s,n}$ be the operator on $\ell_{2}(\mathcal{C})$ given by
$R_{s,n}(\{x_{i}\}_{i\in\mathbb{Z}})=(\{x_{i}\cdot\mathcal{U}_{s}^{*\,r_{s}n}\tilde{W_{s}}^{r_{s}n}\}_{i\in\mathbb{Z}})$
for all $\{x_{i}\}_{i\in\mathbb{Z}}\in\ell_{2}(\mathcal{C}).$ By [33, Proposition 1.2.4 (i)], the operator $R_{s,n}$ is a bounded linear operator on $\ell_{2}(\mathcal{C})$ for each $s\in\{1,\dots,N\}$ and $n\in\mathbb{N}.$ Let $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ be the families of operators on $\ell_{2}(\mathcal{C})$ as defined in the beginning of this section. For each $s\in\{1,\dots,N\}$ and $n\in\mathbb{N},$ we let $T_{n,s}\circ R_{s,n}$ be the composition of $T_{n,s}$ and $R_{s,n},$ that is
$\big((T_{n,s}\circ R_{s,n})(\{x_{i}\}_{i\in\mathbb{Z}})\big)=\{W_{n,j}^{(s)}\,x_{j-r_{s}n}\tilde{W_{s}}^{r_{s}n}\}_{j\in\mathbb{Z}}.$
We will now present an extension of Theorem 4.2 in this setting.
Proposition 4.10.
Under the above notation and assumptions, we have $(2)\Rightarrow(1).$
(1) The families of operators $\{T_{n,1}\circ R_{1,n}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\circ R_{N,n}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
(2) For every $J,m\in\mathbb{N}$ and $\varepsilon>0$, there exists some $F\in\mathcal{F}$ such that we can find families
$\{\tilde{D_{n}}\}_{n\in F},\{\tilde{G_{n}}^{(1)}\}_{n\in F},\dots,\{\tilde{G_{n}}^{(N)}\}_{n\in F},$
$\{D_{n,-J}\}_{n\in F},\{G_{n,-J}^{(1)}\}_{n\in F},\dots,\{G_{n,-J}^{(N)}\}_{n\in F},$
$\{D_{n,-J+1}\}_{n\in F},\{G_{n,-J+1}^{(1)}\}_{n\in F},\dots,\{G_{n,-J+1}^{(N)}\}_{n\in F},$
$\vdots$
$\{D_{n,J}\}_{n\in F},\{G_{n,J}^{(1)}\}_{n\in F},\dots,\{G_{n,J}^{(N)}\}_{n\in F},$
of compact operators on $H$ satisfying
$\|\tilde{D_{n}}-P_{m}\|<\varepsilon,\|D_{n,j}-P_{m}\|<\varepsilon,\|\tilde{G_{n}}^{(s)}-P_{m}\|<\varepsilon,\|G_{n,j}^{(s)}-P_{m}\|<\varepsilon,$
and
$\|W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|\|\tilde{D_{n}}\tilde{W_{\ell}}^{r_{\ell}n}\|\,\|W_{n,j}^{(s)\,-1}G_{n,j}^{(s)}\|\|\tilde{G_{n}}^{(s)}\tilde{W_{s}}^{-r_{s}n}\|<\varepsilon^{2}$
for all $n\in F,j\in\mathbb{Z}$ with $J\leq j\leq J$ and $\ell,s\in\{1,\dots,N\};$ and satisfying in addition that for each distinct $s,\ell\in\{1,\dots,N\}$ and all $n\in F,j\in\mathbb{Z}$ with $J\leq j\leq J$ it holds that
$\|W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)-1}\,G_{n,j}^{(s)}\|\|\tilde{G_{n}}^{(s)}\tilde{W_{s}}^{-r_{s}n}\tilde{W_{\ell}}^{r_{\ell}n}\|<\varepsilon.$
Proof.
Let $\mathcal{O},V_{1},\dots,V_{N}$ be non-empty open subsets of $\ell_{2}(\mathcal{C})$. Then there exist some $J,m_{1}\in\mathbb{N}$ and some $x\in\mathcal{O}$, $y^{(1)}\in V_{1},\dots,y^{(N)}\in V_{N}$ such that $\tilde{p}_{J,m_{1}}x=x$ and $\tilde{p}_{J,m_{1}}y^{(\ell)}=y^{(\ell)}$ for each $\ell\in\{1,\dots,N\}.$ Since $\{P_{m}\}_{m\in\mathbb{N}}$ forms also a right approximate unit for $\mathcal{C}$ by [33, Proposition 2.2.1], it follows that we can find some $m_{2}\in\mathbb{N}$ such that $\tilde{p}_{J,m_{1}}x\tilde{p}_{J,m_{2}}\in\mathcal{O}$ and $\tilde{p}_{J,m_{1}}y^{(\ell)}\tilde{p}_{J,m_{2}}\in V_{l}$ for each $\ell\in\{1,\dots,N\}.$ Set $m:=\text{max}\{m_{1},m_{2}\}.$ Then $\tilde{p}_{J,m}x\tilde{p}_{J,m}\in\mathcal{O}$ and $\tilde{p}_{J,m}y^{(\ell)}\tilde{p}_{J,m}\in V_{l}$ for each $\ell\in\{1,\dots,N\}.$ Put $C_{y}=\sum_{s=1}^{N}\|y^{(s)}\|$ and
$\varepsilon:=\frac{\delta}{(2JN+3)(\|x\|+2C_{y}+\delta)}.$
Choose $F\in\mathcal{F}$ and families
$\{\tilde{D_{n}}\}_{n\in F},\{D_{n,j}\}_{n\in F},\{\tilde{G_{n}}^{(1)}\}_{n\in F},\{G_{n,j}^{(1)}\}_{n\in F},\dots,\{\tilde{G_{n}}^{(N)}\}_{n\in F},\{G_{n,j}^{(N)}\}_{n\in F}$
for each $j\in\mathbb{Z}$ with $J\leq j\leq J$, satisfying the assumptions of (2) with respect to $J$, $m$, and $\varepsilon$. For each $n\in F$, let $u_{n},v_{n}^{(1)},\dots,v_{n}^{(N)}\in\ell_{2}(\mathcal{C})$ be given by
$(u_{n})_{j}=\begin{cases}D_{n,j}\,x_{j}\tilde{D_{n}},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else.}\end{cases}$
and
$(v_{n}^{(\ell)})_{j-r_{\ell}n}=\begin{cases}W_{n,j}^{(\ell)\,-1}\,G_{n,j}^{(\ell)}\,(y^{(\ell)})_{j}\,\tilde{G_{n}}^{(\ell)}\tilde{W_{\ell}}^{-r_{\ell}n},&\text{if }-J\leq j\leq J,\\[6.0pt] 0,&\text{else.}\end{cases}$
By similar arguments as in the proof of Theorem 4.2 part $(2)\Rightarrow(1)$, we can deduce that $\tilde{D_{n}}\neq 0$ and $\tilde{G_{n}}^{(s)}\neq 0$ for all $n\in F,s\in\{1,\dots,N\},$ and moreover
$\|T_{n,\ell}(u_{n})\|\leq\sum_{j=-J}^{J}\|W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|\|\tilde{D_{n}}\tilde{W_{\ell}}^{r_{\ell}n}\|\,\|x\|$
and
$\|T_{n,\ell}(v_{n}^{(s)})\|\leq\sum_{j=-J}^{J}\|W_{n,j-r_{s}n+r_{\ell}n}^{(\ell)}\,W_{n,j}^{(s)\,-1}\,G_{n,j}^{(s)}\|\|\tilde{G_{n}}^{(s)}\tilde{W_{s}}^{-r_{s}n}\tilde{W_{\ell}}^{r_{\ell}n}\|\,\|y^{(s)}\|.$
For each $n\in F$, set
$x_{n}:=u_{n}+\frac{\sqrt{\sum_{\ell=1}^{N}\sum_{j=-J}^{J}\|W_{n,j+r_{\ell}n}^{(\ell)}D_{n,j}\|\|\tilde{D_{n}}\tilde{W_{\ell}}^{r_{\ell}n}\|}}{\sqrt{\sum_{\ell=1}^{N}\sum_{j=-J}^{J}\|W_{n,j}^{(\ell)\,-1}G_{n,j}^{(\ell)}\|\|\tilde{G_{n}}^{(\ell)}\tilde{W_{\ell}}^{-r_{\ell}n}\|}}\sum_{\ell=1}^{N}v_{n}^{(\ell)}.$
and proceed further as in the proof of Theorem 4.2 part $(2)\Rightarrow(1).$
Let again, for each $\ell\in\{1,\dots,N\},$ $\tilde{T}_{\ell}$ be the generalized bilateral weighted shift operator on $\ell_{2}(\mathcal{C})$ with respect to the sequence $\{W_{j}^{(\ell)}\}_{j\in\mathbb{Z}}$ and unitary $\mathcal{U}_{\ell},$ as defined in the beginning of this section. Recalling that the involution on $\mathcal{C}$ is an isometry, by similar arguments as in the proof of Proposition 4.6, we easily obtain the following corollary of Proposition 4.10 in the special case when $\mathcal{F}$ is the family of all infinite subsets of $\mathbb{N}.$
Corollary 4.11.
Under the above notation and assumptions, if there exist dense subsets
$\tilde{H_{0}},\tilde{H_{1}},\dots,\tilde{H_{N}}$,
$H_{0}^{(-J)},H_{1}^{(-J)},\dots,H_{N}^{(-J)},$
$\vdots$
$H_{0}^{(J)},H_{1}^{(J)},\dots,H_{N}^{(J)}$
of $H$ and a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ such that for each $s,\ell\in\{1,\dots,N\},$ each $j\in\mathbb{Z}$ with $J\leq j\leq J$ and every $\tilde{x}\in\tilde{H_{0}},x^{(j)}\in H_{0}^{(j)}$, $\tilde{y_{s}}\in\tilde{H_{s}},y_{s}^{(j)}\in H_{s}^{(j)}$ we have
$\lim_{k\to\infty}\big\|W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}\,y_{s}^{(j)}\big\|\|((\tilde{W_{s}}^{-r_{s}n_{k}})^{*}\tilde{y_{s}}\|\big\|W_{j+r_{\ell}n_{k}}^{(\ell)}\cdots W_{j+1}^{(\ell)}\,x^{(j)}\big\|\|(\tilde{W_{\ell}}^{r_{\ell}n_{k}})^{*}\tilde{x}\|=0,$
and for each distinct $s,\ell\in\{1,\dots,N\}$ and every $y_{s}^{(j)}\in H_{s}^{(j)}$,we have
$\lim_{k\to\infty}\|W_{j+r_{\ell}n_{k}-r_{s}n_{k}}^{(\ell)}\cdots W_{j-r_{s}n_{k}+1}^{(\ell)}W_{j-r_{s}n_{k}+1}^{(s)\,-1}\cdots W_{j}^{(s)\,-1}y_{s}^{(j)}\|\|(\tilde{W_{\ell}}^{r_{\ell}n_{k}})^{*}(\tilde{W_{s}}^{-r_{s}n_{k}})^{*}\tilde{y_{s}}\|=0,$
then the sequences of operators $\{\tilde{T}_{1}^{r_{1}n}\circ R_{1,n}\}_{n\in\mathbb{N}},\dots,\{\tilde{T}_{N}^{r_{N}n}\circ R_{N,n}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive on $\ell_{2}(\mathcal{C}).$
We end this paper with an example where the conditions of Corollary 4.11 are satisfied.
Example 4.12.
Let $N=1$ and $H=L^{2}(\mathbb{R})$. For each $j\in\mathbb{Z},$ let $W_{j}$ be the identity operator on $H.$ Put $\tilde{\alpha}$ to be the function on $\mathbb{R}$ given by $\tilde{\alpha}(t)=t-1$ for all $t\in\mathbb{R}$ and $\tilde{w}==2\chi_{(-\infty,0)}+\frac{1}{2}\chi_{[0,\infty)}.$ If we let $\tilde{W}$ be the weighted composition operator on $H$ with respect to $\tilde{\alpha}$ and $\tilde{w}$, then by some calculations it can be checked that the the conditions of Corollary 4.11 are satisfied with $\tilde{H_{0}}=\tilde{H_{1}}=H_{0}^{(j)}=H_{1}^{(j)}=C_{c}(\mathbb{R})$ for all $j\in\mathbb{Z}.$
Remark 4.13.
We notice that due to [33, Proposition 2.2.1] our results in this section remain valid if we instead of a separable Hilbert space $H$ consider any standard Hilbert module $\ell_{2}(\mathcal{A})$ over a unital C*-algebra $\mathcal{A}.$
5. Algebraic generalizations and axiomatic approach
We recall first the following definition.
Let now $\mathcal{A}$ be a non-unital Banach algebra such that $\mathcal{A}$ is a left ideal in a unital normed algebra $\mathcal{A}_{1}$. We do not assume that the norm on $\mathcal{A}$ extends to $\mathcal{A}_{1}$. We will say that $\mathcal{A}$ satisfies the condition (E) with respect to $\mathcal{A}_{1}$ if
$\|ba\|\leq\|b\|_{1}\|a\|$
for all $a\in\mathcal{A}$ and $b\in\mathcal{A}_{1},$ where  $\|\text{ }\|_{1}$ denotes the norm on $\mathcal{A}_{1}.$
Let $\{p_{\alpha}\}_{\alpha}$ be a set in $\mathcal{A}$ satisfying that given any open subset $O$ of $\mathcal{A}$ and $x\in O$ there exists some $p_{\alpha_{0}}\in\{p_{\alpha}\}_{\alpha}$ such that $p_{\alpha_{0}}^{3}x\in O$ as well. This condition will be called the condition $(\mathcal{P}.)$
For a set $\mathcal{S},$ let $\mathcal{F}$ be a family of subsets of $\mathcal{S}$, and let
$\{\Phi_{t,1}\}_{t\in\mathcal{S}},\ \{\Phi_{t,2}\}_{t\in\mathcal{S}},\ \ldots,\ \{\Phi_{t,N}\}_{t\in\mathcal{S}}$
be families of isometric algebra isomorphism of $\mathcal{A}_{1}$ such that
$\Phi_{t,j}(\mathcal{A})=\mathcal{A}\qquad(t\in\mathcal{S},\ 1\leq j\leq N).$
and such that $\Phi_{t,j|\mathcal{A}}$ is an isometry for all $j\in\{1,\dots,N\}.$ We will assume that the system
$\{\Phi_{t,1}\}_{t\in\mathcal{S}},\ \{\Phi_{t,2}\}_{t\in\mathcal{S}},\ \ldots,\ \{\Phi_{t,N}\}_{t\in\mathcal{S}}$
is disjoint aperiodic with respect to $\{p_{\alpha}\}_{\alpha}$, that is for each fixed $p_{\alpha}$ and every $H\in\mathcal{F}$ there exist some $F\subseteq H$ with $F\in\mathcal{F}$ such that
$p_{\alpha}\,\Phi_{t,\ell}(p_{\alpha})=0,\quad p_{\alpha}\,\Phi_{t,\ell}^{-1}(p_{\alpha})=0,\quad\text{and}\quad p_{\alpha}\,\Phi_{t,\ell}\!\big(\Phi_{t,r}^{-1}(p_{\alpha})\big)=0\quad\text{for all }t\in F$
$\text{ and all }r,\ell\in\{1,\dots,N\}\text{ with }r\neq\ell.$ Further, we will assume that for each $\alpha$ and all $t\in\mathcal{S}$, $r,\ell\in\{1,\dots,N\}$, and $a\in\mathcal{A}$, it holds that
$\|a\,p_{\alpha}\|\leq\|a\|,\quad\|a\,\Phi_{t,\ell}(p_{\alpha})\|\leq\|a\|,$
$\|a\,\Phi_{t,\ell}^{-1}(p_{\alpha})\|\leq\|a\|,\quad\quad\|a\,\Phi_{t,\ell}\big(\Phi_{t,r}^{-1}(p_{\alpha})\big)\|\leq\|a\|.$
This condition will be called the condition (R) throughout this section.
We now set up families $\{b_{t,1}\}_{t\in\mathcal{S}},\cdots,\{b_{t,N}\}_{t\in\mathcal{S}}$ of elements in $\mathcal{A}_{1}$, and for each $t\in\mathcal{S}$ and $\ell\in\{1,\dots,N\}$, we define the operator
$T_{t,\ell}:\mathcal{A}\to\mathcal{A}\quad\text{by}\quad T_{t,\ell}(a)=b_{t,\ell}\,\Phi_{t,\ell}(a).$
Due to the condition $(E)$, it is not hard to check that $T_{t,\ell}$ is a bounded linear operator on $\mathcal{A}$. Further, we will assume that for each $\alpha$ there exists so called $\alpha$-inverse” for $b_{t,\ell}$ denoted by $b_{t,\ell,\alpha}^{-1},$ that is for each $\alpha$ there exists an element $b_{t,\ell,\alpha}^{-1}\in\mathcal{A}_{1}$ satisfying that $b_{t,\ell}b_{t,\ell,\alpha}^{-1}ap_{\alpha}=b_{t,\ell,\alpha}^{-1}b_{t,\ell}ap_{\alpha}=ap_{\alpha}$ for all $a\in\mathcal{A}.$
Also, we will assume that for each $t\in\mathcal{S}$ and $\ell\in\{1,\dots,N\}$ it holds that $b_{t,\ell}a=0$ if and only if $a=0.$ Such condition will be called the condition (C) throughout the manuscript.
5.1. Main results
We have the following theorem.
Theorem 5.1.
Under the above notation and assumptions, the following statements are equivalent.
  1. (1)
    The families $\{T_{t,1}\}_{t\in\mathcal{S}},\cdots,\{T_{t,N}\}_{t\in\mathcal{S}}$ are $d\mathcal{F}$-semi-transitive.
  2. (2)
    For each fixed $p_{\alpha}$ and $\varepsilon>0$ there exists some $F\in\mathcal{F}$ and families $\{d_{t}\}_{t\in F}$, $\{g_{t,1}\}_{t\in F}\dots,\{g_{t,N}\}_{t\in F}$ in $\mathcal{A}$ such that $\|d_{t}-p_{\alpha}^{2}\|<\varepsilon,\quad\|g_{t,l}-p_{\alpha}^{2}\|<\varepsilon,$ and
    $\|b_{t,r,\alpha}^{-1}g_{t,r}\|\|\Phi_{t,l}^{-1}(b_{t,l})d_{t}\|<\varepsilon^{2}\quad\text{for all }t\in F\text{ and }r,l\in\{1,\ldots,N\}.$
    Moreover, for each distinct $r,l\in\{1,\ldots,N\}$ and $t\in F$ it holds that
    $\|\Phi_{t,r}^{-1}(b_{t,r})\Phi_{t,l}^{-1}(b_{t,l,\alpha}^{-1}g_{t,l})\|<\varepsilon.$
Proof.
We prove first (1) $\Rightarrow$ (2). Since the families $\{T_{\Phi_{t,1},b_{t,1}}\}_{t\in S},\dots,\{T_{\Phi_{t,N},b_{t,N}}\}_{t\in S}$ are $d\mathcal{F}$-semi-transitive, given $\varepsilon>0$ and some fixed $p_{\alpha}$ we can find $a_{t}\in\mathcal{A}$ and $\lambda_{t}\in\mathbb{R}^{+}$ such that $\|a_{t}-p_{\alpha}\|<\varepsilon$ and $\|\lambda_{t}T_{\Phi_{t,l}},b_{t,l}(a_{t})-p_{\alpha}\|<\varepsilon\text{ for all }t\in H\text{ and some }H\in\mathcal{F}.$ This gives $\bigl\|\,(a_{t}-p_{\alpha})\,p_{\alpha}\,\bigr\|\;\leq\;\bigl\|a_{t}-p_{\alpha}\bigr\|\;<\;\varepsilon.$
Moreover, since $\bigl\{\ {\Phi_{t,l}}\bigr\}_{t\in S,\;1\leq l\leq N}$ is a disjoint aperiodic system, we can find some $F\in\mathcal{F}$ with $F\subseteq H$ such that $p_{\alpha}\!\Phi_{t,l}^{-1}(p_{\alpha})=0,\;p_{\alpha}\!\Phi_{t,l}(p_{\alpha})=0,\;$ and $\Phi_{t,r}(\Phi_{t,l}^{-1}(p_{\alpha}))=0,\;\text{for all }\;t\in F\;\text{ and }r,l\in\{1,\dots N\}\text{ with }r\neq l.$ Hence, we get
$\qquad\frac{1}{|\lambda_{t}|}\,\left\|\Phi_{t,r}^{-1}\!\bigl(b_{t,r,\alpha}^{-1}\bigr)\,\Phi_{t,r}^{-1}\!\bigl(b_{t,r}\bigr)\,a_{t}\,\Phi_{t,r}^{-1}(p_{\alpha})\,\lambda_{t}\right\|=\left\|\!b_{t,r,\alpha}^{-1}\,\!\bigl(b_{t,r}\,\Phi_{t,r}(a_{t})\,(p_{\alpha})\,\bigr)\lambda_{t}\right\|$
$\left\|a_{t}\,\Phi_{t,r}^{-1}(p_{\alpha})\right\|=\left\|(a_{t}-p_{\alpha})\,\Phi_{t,r}^{-1}(p_{\alpha})\right\|\leq\left\|a_{t}-p_{\alpha}\right\|<\varepsilon.$
and
$\left\|\lambda_{t}\,b_{t,r}\,\Phi_{t,r}(a_{t})\,p_{\alpha}\;-\;p_{\alpha}^{2}\right\|\leq\left\|\lambda_{t}\,b_{t,r}\,\Phi_{t,r}(a_{t})-p_{\alpha}\right\|=\left\|\lambda_{t}\,T_{t,r}(a_{t})-p_{\alpha}\right\|<\epsilon,\qquad$
for all $t\in F$ and $r\in\{1,\dots,N\}$.
For each $t\in F$ and $r\in\{1,\dots,N\},$ we put $g_{t,r}=\lambda_{t}\,T_{t,r}(a_{t})p_{\alpha}.$
Then $\|g_{t,r}-p_{\alpha}^{2}\|\;<\;\varepsilon\quad\text{and}\quad\frac{1}{|\lambda_{t}|}\,\bigl\|\Phi_{t,r}^{-1}\!\bigl(b_{t,r,\alpha}^{-1}\bigr)\,\Phi_{t,r}^{-1}(g_{t,r})\bigr)\bigr\|\;=\;\frac{1}{|\lambda_{t}|}\,\|b_{t,r,\alpha}^{-1}g_{t,r}\|\;<\;\varepsilon.$
$\text{Similarly, for each }t\in F\text{ and }l\in\{1,\dots,N\},\text{ we have}$
$\|\lambda_{t}\,b_{t,l}\,\Phi_{t,l}(a_{t})\,\Phi_{t,l}(p_{\alpha})\|\;=\;\|(\lambda_{t}\,b_{t,l}\,\Phi_{t,l}(a_{t})-p_{\alpha})\,\Phi_{t,l}(p_{\alpha})\|\;\leq\;\|\lambda_{t}\,T_{t,l}(a_{t})-p_{\alpha}\|\;<\;\varepsilon,$
[…]
$\text{ set }d_{t}=a_{t}p_{\alpha}.$
$\text{Then }\|d_{t}-p_{\alpha}^{2}\|\;<\;\varepsilon\quad\text{and}\quad|\lambda_{t}|\,\bigl\|\Phi_{t,l}^{-1}(b_{t,l})\,d_{t}\bigr\|\;=\;\|\lambda_{t}\,b_{t,l}\,\Phi_{t,l}(a_{t}p_{\alpha})\|\;<\;\varepsilon.$
Hence, for all $t\in F$ and $r,l\in\{1,\dots,N\},$ we have
$\|b_{t,r,\alpha}^{-1}(g_{t,r})\|\|\Phi_{t,l}^{-1}(b_{t,l})d_{t}\|<\varepsilon^{2}.$
Further, for all $t\in F$ and each distinct $r,l\in\{1,\dots,N\}$ we get
$\Bigl\|\,\Phi_{t,r}^{-1}\!\bigl(b_{t,r}\bigr)\,\Phi_{t,l}^{-1}\!\bigl(b_{t,l,\alpha}^{-1}\bigr)\Phi_{t,l}^{-1}\!\bigl(g_{t,l}\bigr)\Bigr\|\;=\;\Bigl\|\,\!b_{t,r}\,\Phi_{t,r}\Phi_{t,l}^{-1}\!\bigl(b_{t,l,\alpha}^{-1}g_{t,l}\bigr)\Bigr\|$
$\bigl\|\,\lambda_{t}\,b_{t,r}\,\Phi_{t,r}(a_{t})\,\Phi_{t,r}\!\bigl(\Phi_{t,l}^{-1}(p_{\alpha})\bigr)\bigr\|$
$\bigl\|\,(\lambda_{t}\,T_{t,r}(a_{t})-p_{\alpha})\,\Phi_{t,r}\!\bigl(\Phi_{t,l}^{-1}(p_{\alpha})\bigr)\bigr\|\;<\;\varepsilon,$
where we have used that $p_{\alpha}\,\Phi_{t,r}\!\bigl(\Phi_{t,l}^{-1}(p_{\alpha})\bigr)=0\quad\text{for all }t\in F\text{ and each distinct }r,l\in\{1,\dots,N\}.$
Now we prove the implication $(2)\Rightarrow(1)$. Let $O,V_{1},\dots,V_{N}$ be non-empty open subsets of $\mathcal{A}$. Then we can find some non-zero $x\in O$ and non-zero $y_{l}\in V_{l}$ for each $l\in\{1,\dots,N\}$. There exists some $\alpha$ such that $p_{\alpha}^{3}x\in O$ and $p_{\alpha}^{3}y_{l}\in V_{l}\setminus\{0\}$ for each $l=1,\dots,N$. Clearly, this gives that $p_{\alpha}x\neq 0$ and $p_{\alpha}y_{l}\neq 0$ for each $l=1,\dots,N$. Also, since $O\setminus\{0\},\,V_{1}\setminus\{0\},\dots,V_{N}\setminus\{0\}$ are open and non-empty, there is some $\delta>0$ such that the $\delta$-neighbourhood of $p_{\alpha}^{3}x$ is contained in $O\setminus\{0\}$, and the $\delta$-neighbourhood of $p_{\alpha}^{3}y_{l}$ is contained in $V_{l}\setminus\{0\}$ for each $l=1,\dots,N$.
Let $C_{y}=\max_{1\leq l\leq N}\bigl\{\|y_{l}\|,\bigr\}$ and set
$\varepsilon:=\min\!\left\{\frac{\delta}{4\|x\|\|p_{\alpha}\|},\;\frac{\delta}{4N\sqrt{\|x\|C_{y}}\|p_{\alpha}\|}\frac{\delta}{4NC_{y}\|p_{\alpha}\|},\frac{\|p_{\alpha}^{3}x\|}{2\|p_{\alpha}x\|},\frac{\|p_{\alpha}^{3}y_{1}\|}{2\|p_{\alpha}y_{1}\|},\cdots,\frac{\|p_{\alpha}^{3}y_{N}\|}{2\|p_{\alpha}y_{N}\|}\right\}.$
Choose $F\in\mathcal{F}$ and the families $\{g_{t,1}\}_{t\in F},\dots,\{g_{t,N}\}_{t\in F},\{d_{t}\}_{t\in F}$ satisfying the assumptions in $(2)$ with respect to $\epsilon$ and $p_{\alpha}$.For each $\alpha$ and every $t\in F,l\in\{1,\dots,N\}$ we define the operator $S_{t,l,\alpha}$ on $\mathcal{A}$ by $S_{t,l,\alpha}(a)=\Phi_{t,l}^{-1}(b_{t,l,\alpha}^{-1})\Phi_{t,l}^{-1}(a)$ for all $a\in\mathcal{A}.$ We observe now that for each $t\in F$ and $l\in\{1,\dots,N\}$ it holds that
$\parallel T_{t,l}\!\bigl(d_{t}p_{\alpha}x\bigr)\parallel=\parallel\,b_{t,l}\,\Phi_{t,l}\bigl(d_{t}\bigr)\,\Phi_{t,l}(p_{\alpha}x)\,\parallel\leq\parallel\,b_{t,l}\,\Phi_{t,l}\bigl(d_{t}\bigr)\,\parallel\,\parallel\,\Phi_{t,l}(p_{\alpha}x)\,\parallel$
$\parallel\,\Phi_{t,l}^{-1}\,(b_{t,l})\,d_{t}\,\parallel\;\parallel\,x\,\parallel\|p_{\alpha}\|$
and
$\parallel S_{t,l,\alpha}\!\bigl(g_{t,l}p_{\alpha}\,y_{\ell}\bigr)\parallel=\parallel\,\Phi_{t,l}^{-1}\,(b_{t,l,\alpha}^{-1})\,\Phi_{t,l}^{-1}\!\bigl(g_{t,l}\bigr)\,\Phi_{t,l}^{-1}(p_{\alpha}\,y_{\ell})\,\parallel$
$\leq\parallel\,\Phi_{t,l}^{-1}\,(b_{t,l,\alpha}^{-1}\,g_{t,l})\,\parallel\,\parallel\,\Phi_{t,l}^{-1}(p_{\alpha}\,y_{\ell})\,\parallel=\parallel\,b_{t,l,\alpha}^{-1}\,g_{t,l}\,\parallel\;\parallel\,y_{\ell}\,\parallel\|p_{\alpha}\|.$
Moreover, for all $t\in F$ and each distinct $r,l\in\{1,\dots,N\}$ we have
$\parallel T_{t,r}\!\bigl(S_{t,l,\alpha}(g_{t,l}p_{\alpha}\,y_{\ell})\bigr)\parallel=\parallel\,b_{t,r}\,\Phi_{t,r}\,(\Phi_{t,l}^{-1}\,(b_{t,l,\alpha}^{-1})\bigr)\,\Phi_{t,r}\,\bigl(\Phi_{t,l}^{-1}\,(g_{t,l}p_{\alpha}\,y_{\ell}\bigr)\bigr)\,\parallel$
$\parallel\Phi_{t,r}^{-1}\,(b_{t,r}\,)\Phi_{t,l}^{-1}\,(b_{t,l,\alpha}^{-1})\,\Phi_{t,l}^{-1}\,(g_{t,l})\,\Phi_{t,l}^{-1}\,(p_{\alpha}\,y_{\ell})\,\parallel$
$\leq\parallel\Phi_{t,r}^{-1}\,(b_{t,r}\,)\Phi_{t,l}^{-1}\,(b_{t,l,\alpha}^{-1}g_{t,l})\,\,\parallel\,\parallel y_{\ell}\parallel\|p_{\alpha}\|.$
Finally, since $b_{t,l,\alpha}^{-1}$ is the $\alpha$-inverse of $b_{t,\ell},$ it follows that
$T_{t,l}\!\bigl(S_{t,l,\alpha}(g_{t,l}p_{\alpha}\,y_{\ell}))=g_{t,l}p_{\alpha}\,y_{\ell}$
for all $t\in F$ and each $\ell\in\{1,\dots,N\}.$
Since $\|d_{t}p_{\alpha}x-p_{\alpha}^{3}\,x\|<\|\epsilon\|\|p_{\alpha}x\|<\|p_{\alpha}^{3}\,x\|$, it follows that $d_{t}p_{\alpha}x\neq 0.$ Similarly, $g_{t,l}p_{\alpha}y_{\ell}\neq 0$ for all $t\in F$ and $\ell\in\{1,\dots,N\}$. Hence, since by the assumption we have that for each $t\in\mathcal{S}$ and $\ell\in\{1,\dots,N\}$ it holds that $b_{t,\ell}a=0$ if and only if $a=0,$ it follows that $T_{t,\ell}(d_{t}p_{\alpha}x)\neq 0$ for each $t\in F$ and $\ell\in\{1,\dots,N\}.$ Similarly, since $b_{t,l,\alpha}^{-1}$ is the $\alpha$-inverse of $b_{t,\ell},$ it folllows that $S_{t,\ell,\alpha}(g_{t,\ell}p_{\alpha}y_{\ell})\neq 0$ for all $t\in F$ and $\ell\in{1,\dots,N}.$
For each $t\in F$, we put
$x_{t}:=d_{t}p_{\alpha}x+\frac{\sqrt{\sum_{l=1}^{N}\,\parallel T_{t,l}(d_{t}p_{\alpha}x)\parallel}}{\sqrt{\sum_{l=1}^{N}\,\parallel S_{t,l,\alpha}(g_{t,l}p_{\alpha}y_{\ell})\parallel}\qquad}\,{\sum_{l=1}^{N}\,S_{t,r,\alpha}(g_{t,r}p_{\alpha}y_{r})}.$
By triangle inequality and choice of $\varepsilon$ one can deduce that for each $t\in F$, we have
$x_{t}\in O,\ \text{and}\ \frac{\sqrt{\sum_{\ell=1}^{N}\|S_{t,\ell,\alpha}(g_{t,\ell}p_{\alpha}y_{\ell})\|}}{\sqrt{\sum_{\ell=1}^{N}\|T_{t,\ell}(d_{t}p_{\alpha}x)\|}}T_{t,r}(x_{t})\in V_{r}.$
for all $r\in\{1,\dots,N\}.$
We notice that the assumption that $\{\Phi_{t,1}\}_{t\in\mathcal{S}},\cdots,\{\Phi_{t,N}\}_{t\in\mathcal{S}}$ is a disjoint aperiodic system is only needed for the proof of the implication $(1)\Rightarrow(2)$. Also, we notice that we have used the condition (R) only in the proof of the implication $(1)\Rightarrow(2)$ to obtain the desired inequalities.
The next corollary can easily be deduced from Theorem 5.1.
Corollary 5.2.
Let now $\mathcal{F}$ be the family of all infinite subsets of $\mathbb{N}.$ If there exist dense subsets $\mathcal{A}_{0},\mathcal{B}_{1},\cdots,\mathcal{B}_{N}$ of $\mathcal{A}$ and a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ such that for each $s,\ell\in\{1,\dots,N\},$ every fixed $\alpha$ and all $x\in\mathcal{A}_{0},y_{s}\in\mathcal{B}_{s}$ we have
$\lim_{k\to\infty}\|b_{n_{k},s,\alpha}^{-1}y_{s}\|\|\Phi_{n_{k},l}^{-1}(b_{n_{k},l})x\|=0,$
and, in addition, for each distinct $s,\ell\in\{1,\ldots,N\}$ it holds that
$\lim_{k\to\infty}\|\Phi_{n_{k},\ell}^{-1}(b_{n_{k},\ell})\Phi_{n_{k},s}^{-1}(b_{n_{k},s,\alpha}^{-1}y_{s})\|=0,$
then the families $\{T_{n,1}\}_{n\in\mathbb{N}},\cdots,\{T_{n,N}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
Remark 5.3.
The above results and the proofs can be extended and generalized to the case when $\mathcal{A}$ is a Banach bimodule over a unital normed algebra $\mathcal{A}_{1}.$ The condition (E) remains the same in this case, however, the condition $(\mathcal{P})$ should be then reformulated as follows:
Let $\{p_{\alpha}\}_{\alpha}$ be a set in $\mathcal{A}\cap\mathcal{A}_{1}$ satisfying that given any open subset $O$ of $\mathcal{A}$ there exists some $x\in\mathcal{A}\cup\mathcal{A}_{1}$ and some $p_{\alpha_{0}}\in\{p_{\alpha}\}_{\alpha}$ such that $p_{\alpha_{0}}x\in\mathcal{A}_{1}$ and $p_{\alpha_{0}}^{3}x\in O.$
Also, in this case, we let $\{\Phi_{t,1}\}_{t\in\mathcal{S}},\ \{\Phi_{t,2}\}_{t\in\mathcal{S}},\ \ldots,\ \{\Phi_{t,N}\}_{t\in\mathcal{S}}$ be families of isometric algebra isomorphisms of $\mathcal{A}_{1}$ and
$\{\Psi_{t,1}\}_{t\in\mathcal{S}},\ \{\Psi_{t,2}\}_{t\in\mathcal{S}},\ \ldots,\ \{\Psi_{t,N}\}_{t\in\mathcal{S}}$
be families of linear isometric isomorphisms of $\mathcal{A}$ such that for all $x,y\in\mathcal{A}_{1}$ and $a\in\mathcal{A}$ it holds that $\Psi_{t,j}(xay)=\Phi_{t,j}(x)\Psi_{t,j}(a)\Phi_{t,j}(y)$ for all $t\in\mathcal{S}$ and $j\in\{1,\dots,N\}.$ We express aperiodicity still in terms of the families
$\{\Phi_{t,1}\}_{t\in\mathcal{S}},\ \{\Phi_{t,2}\}_{t\in\mathcal{S}},\ \ldots,\ \{\Phi_{t,N}\}_{t\in\mathcal{S}},$ moreover, the conditions (R) and (C) remain the same in this case. Further, we define the operator
$T_{t,\ell}:\mathcal{A}\to\mathcal{A}\quad\text{by}\quad T_{t,\ell}(a)=b_{t,\ell}\,\Psi_{t,\ell}(a)$
for all $a\in\mathcal{A}.$ Under this set up and these assumptions, it is not hard to extend and generalize the proof of Theorem 5.1 in this setting.
5.2. Applications
We will now apply our results to the case of generalized weighted bilateral shift on $\ell_{2}(\mathcal{C}).$ More precisely, put $\mathcal{A}=\ell_{2}(\mathcal{C}),$ and
$\mathcal{A}_{1}:=\{\{y_{j}\}_{j\in\mathbb{Z}}\mid y_{j}\in B(H)\ \forall j,\,\exists M_{y}>0\text{ such that }\|y_{j}\|\leq M_{y}\ \forall j\}.$
The multiplication on $\mathcal{A}_{1}$ is defined similarly component-wise. We define the norm $\|\cdot\|_{1}$ on $\mathcal{A}_{1}$ as $\|\{y_{j}\}_{j\in\mathbb{Z}}\|_{1}=sup_{j\in\mathbb{Z}}\|y_{j}\|$ for all $\{y_{j}\}_{j\in\mathbb{Z}}\in\mathcal{A}_{1}.$ It is straightforward to check that the condition (E) is satisfied in this case.
For $N\in\mathbb{N},$ let $r_{1},\dots,r_{N}\in\mathbb{N}$ with $r_{1}<r_{2}<\dots<r_{N}$ and let $\mathcal{U}_{1},\dots,\mathcal{U}_{N}$ are unitary operators on $H$. For each $n\in\mathbb{N}$ and $s\in\{1,\ldots,N\},$ we define the map $\Phi_{n,s}:\mathcal{A}_{1}\to\mathcal{A}_{1}$ by
$\Phi_{n,s}\left(\{y_{j}\}_{j\in\mathbb{Z}}\right)=\{\mathcal{U}_{s}^{-r_{s}n}y_{j-r_{s}n}\mathcal{U}_{s}^{r_{s}n}\}_{j\in\mathbb{Z}},$
for all $\{y_{j}\}_{j\in\mathbb{Z}}\in\mathcal{A}_{1},$ (that is the $j$-th coordinate of $\Phi_{n,s}(\{y_{j}\}_{j\in\mathbb{Z}})=\mathcal{U}_{s}^{-r_{s}n}y_{j-r_{s}n}\mathcal{U}_{s}^{r_{s}n}$ for all $j\in\mathbb{Z}$).
Now, the set $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ forms a left approximate unit for $\mathcal{A}=\ell_{2}(\mathcal{C})$ and $\tilde{p}_{J,m}=\tilde{p}_{J,m}^{3}$ for all $(J,m)\in\mathbb{N}^{2}.$ If $\mathcal{F}$ is a family of subsets of $\mathbb{N}$ which is finitely invariant, that is, for each $n\in\mathbb{N}$ and $F\in\mathcal{F}$ it holds that $F\setminus\{1,\dots,n\}\in\mathcal{F}$ also, then, by some calculations it can be checked that $\{\Phi_{n,1}\}_{n\in\mathbb{N}},\cdots,\{\Phi_{n,N}\}_{t\in\mathbb{N}}$ is a disjoint aperiodic system with respect to the left approximate unit $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ and that the condition (R) is satisfied in this case.
For each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\},$ let $\{W_{n,j}^{(\ell)}\}_{j\in\mathbb{Z}}$ be a family of operators in $B(H)$ which is uniformly bounded in norm and such that each $W_{n,j}^{(\ell)}$ has a bounded inverse. Then, for each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\}$, we let $b_{n,\ell}\in\mathcal{A}_{1}$ be given by
$(b_{n,\ell})_{i}=W_{n,i}^{(\ell)}\mathcal{U}_{\ell}^{r_{\ell}n},$
for all $i\in\mathbb{Z}.$ Since for each $n\in\mathbb{N},i\in\mathbb{Z}$ and $\ell\in\{1,\dots,N\},$ the operator $W_{n,i}^{(\ell)}\mathcal{U}_{\ell}^{r_{\ell}n}$ is invertible, it follows that the condition (C) is satisfied in this case.
Then we let $T_{n,\ell}$ be the operator on $\mathcal{A}$ given by
$T_{n,\ell}(a)=b_{n,\ell}\,\Phi_{n,\ell}(a),\qquad\text{for all }a\in\mathcal{A}.$
Once again, since $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}=\{\tilde{p}_{J,m}^{3}\}_{(J,m)\in\mathbb{N}^{2}}$ is a left approximate unit for $\mathcal{A}$ (so that this left approximate unit is indexed over $\mathbb{N}^{2}$), by some calculations one can check that for all $n,J,m\in\mathbb{N},\,s\in\{1,\dots,N\}$ it holds that
$(b_{n,s,(J,m)}^{-1})_{i}=(b_{n,s,J}^{-1})_{i}=\mathcal{U}_{s}^{-r_{s}n}\,W^{(s)\,-1}_{n,i},$
for all $i\in\mathbb{Z}$ with $J\leq i\leq J,$ and $(b_{n,s,(J,m)}^{-1})_{i}=(b_{n,s,J}^{-1})_{i}=0$ for all $i\in\mathbb{Z}$ with $i\leq-J-1$ and all $i\in\mathbb{Z}$ with $i\geq J+1.$ In fact, we write just $b_{n,s,J}^{-1}$ instead of $b_{n,s,(J,m)}^{-1}$ because it does only depend on $J$ and not on $m.$
Finally, note also that for each $a\in\mathcal{A}$ and $j\in\mathbb{Z}$ it holds that $\|a_{j}\|\leq\|a\|.$
Taking all this into account, it is not hard to deduce Theorem 4.2 as a special case of Theorem 5.1.
From now on and in the rest of this section we will assume that $\mathcal{F}$ the family of all infinite subsets of $\mathbb{N}.$ In view or Corollary 5.2, by arguments similar to those in the proofs of [27, Theorem 3.2] and [28, Proposition 2.7], it is neither hard to deduce the following proposition in this special case.
Proposition 5.4.
Under the above notation and assumptions, if for each $j\in\mathbb{Z}$ there exist dense subsets $H_{0}^{(j)},H_{1}^{(j)},\dots,H_{N}^{(j)}$ of $H$ and a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ such that for each $s,\ell\in\{1,\dots,N\}$ and every $x^{(j)}\in H_{0}^{(j)}$, $y_{s}^{(j)}\in H_{s}^{(j)}$ we have
$\lim_{k\to\infty}\big\|W_{n_{k},j}^{(s)\,-1}\,y_{s}^{(j)}\big\|\big\|W_{n_{k},j+r_{\ell}n_{k}}^{(\ell)}\,x^{(j)}\big\|=0,$
and for each distinct $s,\ell\in\{1,\dots,N\}$ and every $y_{s}^{(j)}\in H_{s}^{(j)}$,we have
$\lim_{k\to\infty}\|W_{n_{k},j-r_{s}n_{k}+r_{\ell}n_{k}}^{(\ell)}\,W_{n_{k},j}^{(s)-1}y_{s}^{(j)}\|=0,$
then the sequences of operators $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
Next we let $\Omega$ be a locally compact Hausdorff space, $\alpha_{1},\dots,\alpha_{N}$ be homeomorphisms of $\Omega$, $\{w_{j}^{(1)}\}_{j\in\mathbb{Z}},\dots,\{w_{j}^{(N)}\}_{j\in\mathbb{Z}}\subseteq C_{b}(\Omega)$, where $C_{b}(\Omega)$ denotes the space of all bounded continuous functions on $\Omega$. We will assume that $w_{j}^{(\ell)}>0$ for all $j\in\mathbb{Z}$ and $\ell\in\{1,\dots,N\}$ and that $\|w_{j}^{(\ell)}\|_{\infty}<M$ for all $j\in\mathbb{Z}$, $\ell\in\{1,\dots,N\}$ and some $M>0.$ Now we consider $\mathcal{A}=\ell_{2}(C_{0}(\Omega))$, and
$\mathcal{A}_{1}:=\{\{y_{j}\}_{j\in\mathbb{Z}}\mid y_{j}\in C_{b}(\Omega)\ \forall j,\,\exists M_{y}>0\text{ such that }\|y_{j}\|_{\infty}\leq M_{y}\ \forall j\}.$
The multiplication and the norm on $\mathcal{A}_{1}$ are defined in the same way as in the previous case. Once again, it is straightforward to check that the condition (E) is satisfied in this case.
For each pair of compact subset $K_{1}$ and $K_{2}$ of $\Omega$ with $K_{1}\subseteq K_{2}$ , we let $u_{(K_{1},K_{2})}\in C_{0}(\Omega)$ such that $u_{{(K_{1},K_{2})}|_{K_{1}}}=1$, $\operatorname{supp}u_{(K_{1},K_{2})}=K_{2}$ and $0\leq u_{(K_{1},K_{2})}\leq 1$. For $J\in\mathbb{N}$, we let $\tilde{p}_{(K_{1},K_{2},J)}\in\mathcal{A}$ be given by
[…]
Then it is easily checked that $\{\tilde{p}_{(K_{1},K_{2},J)}^{3}\}_{K_{1}\subseteq K_{2}\subseteq\Omega,\,K_{1},K_{2}\text{ compact ,}J\in\mathbb{N}}$ satisfies the condition $(\mathcal{P}.)$
For each $\ell\in\{1,\dots,N\}$ and $k\in\mathbb{N}$, we let $\tilde{\Phi}_{k,\ell}:\mathcal{A}\to\mathcal{A}$ be given by
$\tilde{\Phi}_{k,\ell}\big(\{f_{j}\}_{j\in\mathbb{Z}}\big)=\{f_{j-k}\circ\alpha_{\ell}^{k}\,\}_{j\in\mathbb{Z}},$
for all $\{f_{j}\}_{j\in\mathbb{Z}}\subset\mathcal{A}$. Let now $r_{1},\dots,r_{N}\in\mathbb{N}$ with $r_{1}<r_{2}<\dots<r_{N}$. For each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\}$, set
$\Phi_{n,\ell}=\tilde{\Phi}_{r_{\ell}n,\ell}.$
Clearly, the system of isometric isomorphisms $\{\Phi_{k,1}\}_{k\in\mathbb{N}},\dots,\{\Phi_{k,N}\}_{k\in\mathbb{N}}$ is disjoint aperiodic with respect to
$\{\tilde{p}_{(K_{1},K_{2},J)}\}_{K_{1}\subseteq K_{2}\subseteq\Omega,\,K_{1},K_{2}\text{ compact ,}J\in\mathbb{N}}.$
Also, the condition (R) is satisfied in this case.
For each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\}$, let $b_{n,\ell}\in\mathcal{A}_{1}$ be given by
$(b_{n,\ell})_{j+r_{\ell}n}=\prod_{k=1}^{r_{\ell}n}\big(w_{j+k}^{(\ell)}\circ\alpha_{\ell}^{n-k}\big)\quad\text{for all }j\in\mathbb{Z}.$
Since for each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\},$ the function $w_{j}^{(\ell)}>0,$ it follows that the condition (C) is satisfied in this case.
Now, for each $K_{2}\subseteq\Omega$ compact, we can find some $v_{K_{2}}\in C_{0}(\Omega)$ with $0\leq v_{K_{2}}\leq 1$ such that ${v_{K_{2}}}_{|_{K_{2}}}=1.$ Recalling that the approximate unit for $\mathcal{A}$ in this case is indexed over all triples $(K_{1},K_{2},J)$ with $K_{1}\subseteq K_{2}\subseteq\Omega,\,K_{1},K_{2}\text{ compact, }J\in\mathbb{N},$ for each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\}$, and every fixed triple $(K_{1},K_{2},J)$ we let $b_{n,\ell,(K_{1},K_{2},J)}^{-1}\in\mathcal{A}_{1}$ be given by
$(b_{n,\ell,(K_{1},K_{2},J)}^{-1})_{j+r_{\ell}n}=\frac{v_{K_{2}}}{\prod_{k=1}^{r_{\ell}n}\big(w_{j+k}^{(\ell)}\circ\alpha_{\ell}^{n-k}\big)}$
for all $j\in\mathbb{Z}$ with $J-r_{\ell}n\leq j\leq J-r_{\ell}n$ and $(b_{n,\ell,(K_{1},K_{2},J)}^{-1})_{j+r_{\ell}n}=0$ else. Then, for all $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\},$ we let $T_{n,\ell}:\mathcal{A}\to\mathcal{A}$ be given by
$T_{n,\ell}(a)=b_{n,\ell}\,\Phi_{n,\ell}(a)\quad\text{for all }a\in\mathcal{A}.$
The next corollary can be deduced by some calculations from Theorem 5.1.
Corollary 5.5.
Under the above notation and assumptions, the following statements are equivalent:
(1) The sequences of operators
$\{T_{n,1}\}_{n\in\mathbb{N}},\ \dots,\ \{T_{n,N}\}_{n\in\mathbb{N}}$
are $d\mathcal{F}$-semi-transitive on $\mathcal{A}$.
(2) For every $J\in\mathbb{N}$ and each compact subset $K$ of $\Omega$, there exists a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ such that
$\lim_{k\to\infty}\Bigg(\sup_{t\in K}\frac{1}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j+i-r_{s}n_{k}}^{(s)}\circ\alpha_{s}^{r_{s}n_{k}-i}\big)(t)}\Bigg)\Bigg(\sup_{t\in K}\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j+i}^{(\ell)}\circ\alpha_{\ell}^{-i}\big)(t)\Bigg)=0$
for all $j\in\mathbb{Z}$ with $J\leq j\leq J$ and every $l,s\in\{1,\dots,N\}$. Moreover, for each distinct $s,\ell\in\{1,\dots,N\}$ and all $j\in\mathbb{Z}$ with $J\leq j\leq J$, it holds that
$\lim_{k\to\infty}\Bigg(\sup_{t\in K}\frac{\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j-r_{\ell}n_{k}-i}^{(\ell)}\circ\alpha_{\ell}^{-i}\circ\alpha_{s}^{r_{s}n_{k}}\big)(t)}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j-r_{s}n_{k}+i}^{(s)}\circ\alpha_{s}^{n_{k}-i}\big)(t)}\Bigg)=0.$
In sequel, we assume that $\Omega=\mathbb{R}.$ and $\mathcal{C}=C_{0}(\mathbb{R}).$ As in [26], for $\tau\in C_{b}(\mathbb{R}),$ we put
$\mathcal{C}_{\tau}:=\{f\in\mathcal{C}:\,\sum_{k=0}^{\infty}\|f\tau^{k}\|_{\infty}<\infty\}.$
For each $f\in\mathcal{C}_{\tau}$ we define
$\|f\|_{\tau}:=\sum_{k=0}^{\infty}\|f\tau^{k}\|_{\infty}.$
Then, $\mathcal{C}_{\tau}$ is a Banach algebra [23]. We will call this algebra Segal algebra corresponding to $\tau$.
As observed in [26], for a homeomorphism $\alpha$ of $\mathbb{R},$ we consider corresponding Segal algebra $\mathcal{C}_{\tau\circ\alpha^{k}}$ for all $k\in\mathbb{Z}$. Clearly, $f\cdot(\tau\circ\alpha^{k})\in\mathcal{C}$ for all $k\in\mathbb{Z}$ whenever $f\in C_{0}(\mathbb{R})$ since $\tau\circ\alpha^{k}\in C_{b}(\mathbb{R})$ for all $k\in\mathbb{Z}.$ By definition, $\|f\|_{\infty}\leq\|f\|_{\tau\circ\alpha^{k}}$ whenever $f\in\mathcal{C}_{\tau\circ\alpha^{k}}.$ We recall now the following lemmas from [26].
Lemma 5.6.
[26, Lemma 3.2] Let $\tau\in C_{b}(\mathbb{R})$ and $\epsilon_{1},\epsilon_{2}\in(0,1)$ with $\epsilon_{2}<\epsilon_{1}.$ Then for every compact subset $K_{\epsilon_{2}}^{(\tau)}\subseteq|\tau|^{-1}(\left[0,\epsilon_{2}\right]),$ there exists a function $\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}\in\mathcal{C}_{\tau}$ such that the followings hold:
  1. (1)
    $\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}=1$ on $K_{\epsilon_{2}}^{(\tau)}$,
  2. (2)
    $\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}\in C_{c}(\mathbb{R}),$
  3. (3)
    ${\rm supp}\,\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}\subseteq|\tau|^{-1}(\left[0,\epsilon_{1}\right])).$
Lemma 5.7.
[26, Lemma 3.3] For $\tau\in C_{b}(\mathbb{R})$ and each $f\in\mathcal{C}_{\tau}$, there is a sequence $(g_{n})$ in the set $\{\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}:\,\,0<\epsilon_{2}<\epsilon_{1}<1,\,\,K_{\epsilon_{2}}^{(\tau)}\text{ is compact}\}$ such that $fg_{n}\rightarrow f$ in $\mathcal{C}_{\tau}$.
Lemma 5.8.
[26, Lemma 3.5] For each $k\in\mathbb{Z}$, $f\circ\alpha\in\mathcal{C}_{\tau\circ\alpha^{k+1}}$ if and only if $f\in\mathcal{C}_{\tau\circ\alpha^{k}}$. Moreover, in this case, $\|f\circ\alpha\|_{\tau\circ\alpha^{k+1}}=\|f\|_{\tau\circ\alpha^{k}}$.
As in [26], let
${\rm c}_{0}^{\mathcal{C},\tau}:=\Big\{s:=(s_{k})\in\prod_{k\in\mathbb{Z}}\mathcal{C}_{\tau\circ\alpha^{k}}:\,\lim_{k\rightarrow\infty}\|s_{k}\|_{\tau\circ\alpha^{k}}=\lim_{k\rightarrow\infty}\|s_{-k}\|_{\tau\circ\alpha^{-k}}=0\Big\}$
and we equip ${\rm c_{0}}^{\mathcal{C},\tau}$ with the norm
$\|s\|_{0}:=\sup_{k\in\mathbb{Z}}\|s_{k}\|_{\tau\circ\alpha^{k}}.$
Since by [23] we have that $\mathcal{C}_{\tau\circ\alpha^{k}}$ is Banach algebra for all $k\in\mathbb{Z},$ it is not hard to check that ${\rm c}_{0}^{\mathcal{C},\tau}$ is a non-unital Banach algebra. We will from now on denote ${\rm c_{0}}^{\mathcal{C},\tau}$ by $\mathcal{A}.$ Moreover, we let $\mathcal{A}_{1}$ be the same as above in the previous case when we considered $\mathcal{A}=\ell_{2}(C_{0}(\Omega)).$ It is straightforward to check that the condition (E) is satisfied also in this case when ${\rm c_{0}}^{\mathcal{C},\tau}=\mathcal{A}$ (and $\mathcal{A}_{1}$ is as above in the previous case).
For $\epsilon_{1},\epsilon_{2}\in(0,1)$ with $\epsilon_{2}<\epsilon_{1},J\in\mathbb{N},$ and any collection of compact subsets $K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J+1})},\cdots,K_{\epsilon_{2}}^{(\tau\circ\alpha^{J})}$ of $\mathbb{R},$ we let $\tilde{p}_{(K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J+1})},\cdots,K_{\epsilon_{2}}^{(\tau\circ\alpha^{J})}),\epsilon_{1}}\in\mathcal{A}$ be given by
[…]
(where $\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}$ is the function from Lemma 5.6 with compact ${\rm supp}\,\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\subseteq\\ |\tau\circ\alpha^{j}|^{-1}(\left[0,\epsilon_{1}\right])$). By the proof of Lemma 5.7 (see [26]) one can deduce that the condition ($\mathcal{P}$) is satisfied in this case.
Let now $\alpha=\alpha_{1}=\cdots=\alpha_{N}.$ For each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\},$ let $\Phi_{n,\ell}$ be the same as above in the previous case when we considered $\mathcal{A}=\ell_{2}(C_{0}(\Omega)).$ It is not hard to check that the system $\{\Phi_{k,1}\}_{k\in\mathbb{N}},\dots,\{\Phi_{k,N}\}_{k\in\mathbb{N}}$ is again a system of isometric isomorphisms also in this case when ${\rm c_{0}}^{\mathcal{C},\tau}=\mathcal{A}.$ Moreover, it is disjoint apreiodic and satisfies the condition (R) with respect to
$\{\tilde{p}_{((K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J+1})},\cdots,K_{\epsilon_{2}}^{(\tau\circ\alpha^{J})}),\epsilon_{1})}\}.$
For each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\}$, let $\{w_{j}^{(1)}\}_{j\in\mathbb{Z}},\dots,\{w_{j}^{(N)}\}_{j\in\mathbb{Z}}\subseteq C_{b}(\mathbb{R}),b_{n,\ell}\in\mathcal{A}_{1}$ and $T_{n,\ell}:\mathcal{A}\to\mathcal{A}$ be as above in the previous case (when we considered $\mathcal{A}=\ell_{2}(C_{0}(\Omega))$).
Now, given $\epsilon_{1}\in(0,1),$ we choose some $\epsilon_{3}\in(\epsilon_{1},1).$ Since ${\rm supp}\,\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\subseteq\\ |\tau\circ\alpha^{j}|^{-1}(\left[0,\epsilon_{1}\right])$ and ${\rm supp}\,\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}$ is compact, by Lemma 5.6, we can find some function $v_{\bigl(supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\bigr),\epsilon_{3}}\in\mathcal{C}_{\tau\circ\alpha^{j}}$ such that $v_{\bigl(supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\bigr),\epsilon_{3}}(t)=1$ for all $t\in supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}},\text{ }supp\text{ }v_{\bigl(supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\bigr),\epsilon_{3}}$ is compact and $supp\text{ }v_{\bigl(supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\bigr),\epsilon_{3}}\subseteq|\tau\circ\alpha^{j}|^{-1}(\left[0,\epsilon_{3}\right]).$ Therefore, for each $n\in\mathbb{N},\ell\in\{1,\dots,N\},J\in\mathbb{N},\epsilon_{1},\epsilon_{2}\in(0,1)$ with $\epsilon_{2}<\epsilon_{1}$ and any collection of compact subsets
$K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J+1})},\cdots,K_{\epsilon_{2}}^{(\tau\circ\alpha^{J})},$
we let $b_{n,\ell,((K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J+1})},\cdots,K_{\epsilon_{2}}^{(\tau\circ\alpha^{J})}),\epsilon_{1})}^{-1}\in\mathcal{A}_{1}$ be given by
$(b_{n,\ell,((K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J+1})},\cdots,K_{\epsilon_{2}}^{(\tau\circ\alpha^{J})}),\epsilon_{1})}^{-1})_{j+r_{\ell}n}=\frac{v_{\bigl(supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j+r_{\ell}n})},\epsilon_{1}}\bigr),\epsilon_{3}}}{\prod_{k=1}^{r_{\ell}n}\big(w_{j+k}^{(\ell)}\circ\alpha^{n-k}\big)}$
for all $j\in\mathbb{Z}$ with $J-r_{\ell}n\leq j\leq J-r_{\ell}n$ (where we choose $\epsilon_{3}\in(\epsilon_{1},1)$) and $(b_{n,\ell,(\epsilon_{2},\epsilon_{1}J)}^{-1})_{j+r_{\ell}n}=0$ else.
By some calculations, it is not hard to deduce the next corollary as a corollary of Theorem 5.1.
Corollary 5.9.
Under the above notation and assumptions, the following statements are equivalent:
(1) The sequences of operators
$\{T_{n,1}\}_{n\in\mathbb{N}},\ \dots,\ \{T_{n,N}\}_{n\in\mathbb{N}}$
are $d\mathcal{F}$-semi-transitive on $\mathcal{A}$.
(2) For every $J\in\mathbb{N}$ and finite collection
$\{K_{\epsilon_{J}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{-J+1}}^{(\tau\circ\alpha^{-J+1})},\ldots,K_{\epsilon_{0}}^{\tau},\ldots,K_{\epsilon_{J}}^{(\tau\circ\alpha^{J})}\}$
of compact subsets of $\mathbb{R}$, there exists a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ such that
$\lim_{k\to\infty}\Bigg(\sup_{t\in K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})}}\frac{1}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j+i-r_{s}n_{k}}^{(s)}\circ\alpha^{r_{s}n_{k}-i}\big)(t)}\Bigg)\Bigg(\sup_{t\in K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})}}\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j+i}^{(\ell)}\circ\alpha^{-i}\big)(t)\Bigg)=0$
for all $j\in\mathbb{Z}$ with $J\leq j\leq J$ and every $l,s\in\{1,\dots,N\}$. Moreover, for each distinct $s,\ell\in\{1,\dots,N\}$ and all $j\in\mathbb{Z}$ with $J\leq j\leq J$, it holds that
$\lim_{k\to\infty}\Bigg(\sup_{t\in K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})}}\frac{\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j-r_{\ell}n_{k}-i}^{(\ell)}\circ\alpha^{-i}\circ\alpha^{r_{s}n_{k}}\big)(t)}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j-r_{s}n_{k}+i}^{(s)}\circ\alpha^{n_{k}-i}\big)(t)}\Bigg)=0.$
Example 5.10.
Let $N=1,r_{1}=1$ and choose some $\{w_{j}\}_{j\in\mathbb{Z}}\subseteq L^{\infty}(\mathbb{R})$ such that $w_{j}>0$, $w_{j}^{-1}\in L^{\infty}(\mathbb{R})$ and $\parallel w_{j}\parallel\leq M$ for all $j\in\mathbb{Z}$ and some $M>1$. Let $\alpha(t)=t-1$ for all $t\in\mathbb{R}.$ If there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}\leq 1-\varepsilon,w_{j}\chi_{(-\infty,0)}=1$ for all $j\geq 0$ and $w_{j}=1$ for all $j<0$; or if there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}=1,w_{j}\chi_{(-\infty,0)}\geq 1+\varepsilon$ for all $j<0$ and $w_{j}=1$ for all $j\geq 0,$ then the conditions of Corollary 5.5 and Corollary 5.9 are satisfied in this case.