### v7

# Disjoint supercyclicity in Banach algebras

Abstract.
In this paper, we consider operators that are a composition of an isometric isomorphism and a left multiplier on a normed algebra and we characterize disjoint $\mathcal{F}$-semi-transitive and disjoint supercyclic such operators on a large class of non-unital normed algebras. It turns out that generalized weighted bilateral shifts on the standard Hilbert C*-module is just a special case of our theory. Generalized weighted composition operators on the normed algebra of operator-valued continuous functions vanishing at infinity on a locally compact, non-compact Hausdorff space are another special case of our theory. Also, we illustrate our results with concrete examples.
Keywords:
Mathematics Subject Classification (2010) Primary MSC 47A16, Secondary MSC 54H20.
1. Introduction and preliminaries
We recall first the following definition.
Definition 1.1.
[18, Definition 4.1] Let $\mathcal{S}$ be a set, $\mathcal{F}$ be a family of subsets of $\mathcal{S}$ and $\{T_{t,1}\}_{t\in\mathcal{S}},\dots,\{T_{t,N}\}_{t\in\mathcal{S}}$ be families of bounded linear operators on a Banach space $X$. We say that $\{T_{t,1}\}_{t\in\mathcal{S}},\dots,\{T_{t,N}\}_{t\in\mathcal{S}}$ are disjoint $\mathcal{F}$-semi-transitive, or shortly $d\mathcal{F}$-semi-transitive, if for every collection of non-empty open subsets $\mathcal{O},V_{1},\dots,V_{N}$ of $X$, there exists some $F\in\mathcal{F}$ such that for all $t\in F$ there exists some $\lambda_{t}\in\mathbb{R}^{+}$ satisfying that
$\mathcal{O}\cap\lambda_{t}T_{t,1}^{-1}(V_{1})\cap\dots\cap\lambda_{t}T_{t,N}^{-1}(V_{N})\neq\varnothing.$
We recall also that the families $\{T_{t,1}\}_{t\in S},\dots,\{T_{t,N}\}_{t\in S}$ are said to be disjoint $\mathcal{F}$-transitive, or shortly $d\mathcal{F}$-transitive, if for every collection of non-empty open subsets $\mathcal{O},V_{1},\dots,V_{N}$ of $X$, there exists some $F\in\mathcal{F}$ such that for all $t\in F$ we have
$\mathcal{O}\cap T_{t,1}^{-1}(V_{1})\cap\dots\cap T_{t,N}^{-1}(V_{N})\neq\varnothing.$
Let now $\mathcal{A}$ be a non-unital Banach algebra such that $\mathcal{A}$ is a left ideal in a unital normed algebra $\mathcal{A}_{1}$. We do not assume that the norm on $\mathcal{A}$ extends to $\mathcal{A}_{1}$. We will say that $\mathcal{A}$ satisfies the condition (E) with respect to $\mathcal{A}_{1}$ if
$\|ba\|\leq\|b\|_{1}\|a\|$
for all $a\in\mathcal{A}$ and $b\in\mathcal{A}_{1},$ where  $\|\text{ }\|_{1}$ denotes the norm on $\mathcal{A}_{1}.$
Let $\{p_{\alpha}\}_{\alpha}$ be a set in $\mathcal{A}$ satisfying that given any open subset $O$ of $\mathcal{A}$ and $x\in O$ there exists some $p_{\alpha_{0}}\in\{p_{\alpha}\}_{\alpha}$ such that $p_{\alpha_{0}}^{3}x\in O$ as well. This condition will be called the condition $(\mathcal{P}.)$
For a set $\mathcal{S},$ let $\mathcal{F}$ be a family of subsets of $\mathcal{S}$, and let
$\{\Phi_{t,1}\}_{t\in\mathcal{S}},\ \{\Phi_{t,2}\}_{t\in\mathcal{S}},\ \ldots,\ \{\Phi_{t,N}\}_{t\in\mathcal{S}}$
be families of isometric algebra isomorphism of $\mathcal{A}_{1}$ such that
$\Phi_{t,j}(\mathcal{A})=\mathcal{A}\qquad(t\in\mathcal{S},\ 1\leq j\leq N).$
and such that $\Phi_{t,j|\mathcal{A}}$ is an isometry for all $j\in\{1,\dots,N\}.$ We will assume that the system
$\{\Phi_{t,1}\}_{t\in\mathcal{S}},\ \{\Phi_{t,2}\}_{t\in\mathcal{S}},\ \ldots,\ \{\Phi_{t,N}\}_{t\in\mathcal{S}}$
is disjoint aperiodic with respect to $\{p_{\alpha}\}_{\alpha}$, that is for each fixed $p_{\alpha}$ and every $H\in\mathcal{F}$ there exist some $F\subseteq H$ with $F\in\mathcal{F}$ such that
$p_{\alpha}\,\Phi_{t,\ell}(p_{\alpha})=0,\quad p_{\alpha}\,\Phi_{t,\ell}^{-1}(p_{\alpha})=0,\quad\text{and}\quad p_{\alpha}\,\Phi_{t,\ell}\!\big(\Phi_{t,r}^{-1}(p_{\alpha})\big)=0\quad\text{for all }t\in F$
$\text{ and all }r,\ell\in\{1,\dots,N\}\text{ with }r\neq\ell.$ Further, we will assume that for each $\alpha$ and all $t\in\mathcal{S}$, $r,\ell\in\{1,\dots,N\}$, and $a\in\mathcal{A}$, it holds that
$\|a\,p_{\alpha}\|\leq\|a\|,\quad\|a\,\Phi_{t,\ell}(p_{\alpha})\|\leq\|a\|,$
$\|a\,\Phi_{t,\ell}^{-1}(p_{\alpha})\|\leq\|a\|,\quad\quad\|a\,\Phi_{t,\ell}\big(\Phi_{t,r}^{-1}(p_{\alpha})\big)\|\leq\|a\|.$
This condition will be called the condition (R) throughout this section.
We now set up families $\{b_{t,1}\}_{t\in\mathcal{S}},\cdots,\{b_{t,N}\}_{t\in\mathcal{S}}$ of elements in $\mathcal{A}_{1}$, and for each $t\in\mathcal{S}$ and $\ell\in\{1,\dots,N\}$, we define the operator
$T_{t,\ell}:\mathcal{A}\to\mathcal{A}\quad\text{by}\quad T_{t,\ell}(a)=b_{t,\ell}\,\Phi_{t,\ell}(a).$
Due to the condition $(E)$, it is not hard to check that $T_{t,\ell}$ is a bounded linear operator on $\mathcal{A}$. Further, we will assume that for each $\alpha$ there exists so called $\alpha$-inverse” for $b_{t,\ell}$ denoted by $b_{t,\ell,\alpha}^{-1},$ that is for each $\alpha$ there exists an element $b_{t,\ell,\alpha}^{-1}\in\mathcal{A}_{1}$ satisfying that $b_{t,\ell}b_{t,\ell,\alpha}^{-1}ap_{\alpha}=b_{t,\ell,\alpha}^{-1}b_{t,\ell}ap_{\alpha}=ap_{\alpha}$ for all $a\in\mathcal{A}.$
Also, we will assume that for each $t\in\mathcal{S}$ and $\ell\in\{1,\dots,N\}$ it holds that $b_{t,\ell}a=0$ if and only if $a=0.$ Such condition will be called the condition (C) throughout the manuscript.
2. Main results
We have the following theorem.
Theorem 2.1.
Under the above notation and assumptions, the following statements are equivalent.
  1. (1)
    The families $\{T_{t,1}\}_{t\in\mathcal{S}},\cdots,\{T_{t,N}\}_{t\in\mathcal{S}}$ are $d\mathcal{F}$-semi-transitive.
  2. (2)
    For each fixed $p_{\alpha}$ and $\varepsilon>0$ there exists some $F\in\mathcal{F}$ and families $\{d_{t}\}_{t\in F}$, $\{g_{t,1}\}_{t\in F}\dots,\{g_{t,N}\}_{t\in F}$ in $\mathcal{A}$ such that $\|d_{t}-p_{\alpha}^{2}\|<\varepsilon,\quad\|g_{t,l}-p_{\alpha}^{2}\|<\varepsilon,$ and
    $\|b_{t,r,\alpha}^{-1}g_{t,r}\|\|\Phi_{t,l}^{-1}(b_{t,l})d_{t}\|<\varepsilon^{2}\quad\text{for all }t\in F\text{ and }r,l\in\{1,\ldots,N\}.$
    Moreover, for each distinct $r,l\in\{1,\ldots,N\}$ and $t\in F$ it holds that
    $\|\Phi_{t,r}^{-1}(b_{t,r})\Phi_{t,l}^{-1}(b_{t,l,\alpha}^{-1}g_{t,l})\|<\varepsilon.$
Proof.
We prove first (1) $\Rightarrow$ (2). Since the families $\{T_{\Phi_{t,1},b_{t,1}}\}_{t\in S},\dots,\{T_{\Phi_{t,N},b_{t,N}}\}_{t\in S}$ are $d\mathcal{F}$-semi-transitive, given $\varepsilon>0$ and some fixed $p_{\alpha}$ we can find $a_{t}\in\mathcal{A}$ and $\lambda_{t}\in\mathbb{R}^{+}$ such that $\|a_{t}-p_{\alpha}\|<\varepsilon$ and $\|\lambda_{t}T_{\Phi_{t,l}},b_{t,l}(a_{t})-p_{\alpha}\|<\varepsilon\text{ for all }t\in H\text{ and some }H\in\mathcal{F}.$ This gives $\bigl\|\,(a_{t}-p_{\alpha})\,p_{\alpha}\,\bigr\|\;\leq\;\bigl\|a_{t}-p_{\alpha}\bigr\|\;<\;\varepsilon.$
Moreover, since $\bigl\{\ {\Phi_{t,l}}\bigr\}_{t\in S,\;1\leq l\leq N}$ is a disjoint aperiodic system, we can find some $F\in\mathcal{F}$ with $F\subseteq H$ such that $p_{\alpha}\!\Phi_{t,l}^{-1}(p_{\alpha})=0,\;p_{\alpha}\!\Phi_{t,l}(p_{\alpha})=0,\;$ and $\Phi_{t,r}(\Phi_{t,l}^{-1}(p_{\alpha}))=0,\;\text{for all }\;t\in F\;\text{ and }r,l\in\{1,\dots N\}\text{ with }r\neq l.$ Hence, we get
$\qquad\frac{1}{|\lambda_{t}|}\,\left\|\Phi_{t,r}^{-1}\!\bigl(b_{t,r,\alpha}^{-1}\bigr)\,\Phi_{t,r}^{-1}\!\bigl(b_{t,r}\bigr)\,a_{t}\,\Phi_{t,r}^{-1}(p_{\alpha})\,\lambda_{t}\right\|=\left\|\!b_{t,r,\alpha}^{-1}\,\!\bigl(b_{t,r}\,\Phi_{t,r}(a_{t})\,(p_{\alpha})\,\bigr)\lambda_{t}\right\|$
$\left\|a_{t}\,\Phi_{t,r}^{-1}(p_{\alpha})\right\|=\left\|(a_{t}-p_{\alpha})\,\Phi_{t,r}^{-1}(p_{\alpha})\right\|\leq\left\|a_{t}-p_{\alpha}\right\|<\varepsilon.$
and
$\left\|\lambda_{t}\,b_{t,r}\,\Phi_{t,r}(a_{t})\,p_{\alpha}\;-\;p_{\alpha}^{2}\right\|\leq\left\|\lambda_{t}\,b_{t,r}\,\Phi_{t,r}(a_{t})-p_{\alpha}\right\|=\left\|\lambda_{t}\,T_{t,r}(a_{t})-p_{\alpha}\right\|<\epsilon,\qquad$
for all $t\in F$ and $r\in\{1,\dots,N\}$.
For each $t\in F$ and $r\in\{1,\dots,N\},$ we put $g_{t,r}=\lambda_{t}\,T_{t,r}(a_{t})p_{\alpha}.$
Then $\|g_{t,r}-p_{\alpha}^{2}\|\;<\;\varepsilon\quad\text{and}\quad\frac{1}{|\lambda_{t}|}\,\bigl\|\Phi_{t,r}^{-1}\!\bigl(b_{t,r,\alpha}^{-1}\bigr)\,\Phi_{t,r}^{-1}(g_{t,r})\bigr)\bigr\|\;=\;\frac{1}{|\lambda_{t}|}\,\|b_{t,r,\alpha}^{-1}g_{t,r}\|\;<\;\varepsilon.$
$\text{Similarly, for each }t\in F\text{ and }l\in\{1,\dots,N\},\text{ we have}$
$\|\lambda_{t}\,b_{t,l}\,\Phi_{t,l}(a_{t})\,\Phi_{t,l}(p_{\alpha})\|\;=\;\|(\lambda_{t}\,b_{t,l}\,\Phi_{t,l}(a_{t})-p_{\alpha})\,\Phi_{t,l}(p_{\alpha})\|\;\leq\;\|\lambda_{t}\,T_{t,l}(a_{t})-p_{\alpha}\|\;<\;\varepsilon,$
[…]
$\text{ set }d_{t}=a_{t}p_{\alpha}.$
$\text{Then }\|d_{t}-p_{\alpha}^{2}\|\;<\;\varepsilon\quad\text{and}\quad|\lambda_{t}|\,\bigl\|\Phi_{t,l}^{-1}(b_{t,l})\,d_{t}\bigr\|\;=\;\|\lambda_{t}\,b_{t,l}\,\Phi_{t,l}(a_{t}p_{\alpha})\|\;<\;\varepsilon.$
Hence, for all $t\in F$ and $r,l\in\{1,\dots,N\},$ we have
$\|b_{t,r,\alpha}^{-1}(g_{t,r})\|\|\Phi_{t,l}^{-1}(b_{t,l})d_{t}\|<\varepsilon^{2}.$
Further, for all $t\in F$ and each distinct $r,l\in\{1,\dots,N\}$ we get
$\Bigl\|\,\Phi_{t,r}^{-1}\!\bigl(b_{t,r}\bigr)\,\Phi_{t,l}^{-1}\!\bigl(b_{t,l,\alpha}^{-1}\bigr)\Phi_{t,l}^{-1}\!\bigl(g_{t,l}\bigr)\Bigr\|\;=\;\Bigl\|\,\!b_{t,r}\,\Phi_{t,r}\Phi_{t,l}^{-1}\!\bigl(b_{t,l,\alpha}^{-1}g_{t,l}\bigr)\Bigr\|$
$\bigl\|\,\lambda_{t}\,b_{t,r}\,\Phi_{t,r}(a_{t})\,\Phi_{t,r}\!\bigl(\Phi_{t,l}^{-1}(p_{\alpha})\bigr)\bigr\|$
$\bigl\|\,(\lambda_{t}\,T_{t,r}(a_{t})-p_{\alpha})\,\Phi_{t,r}\!\bigl(\Phi_{t,l}^{-1}(p_{\alpha})\bigr)\bigr\|\;<\;\varepsilon,$
where we have used that $p_{\alpha}\,\Phi_{t,r}\!\bigl(\Phi_{t,l}^{-1}(p_{\alpha})\bigr)=0\quad\text{for all }t\in F\text{ and each distinct }r,l\in\{1,\dots,N\}.$
Now we prove the implication $(2)\Rightarrow(1)$. Let $O,V_{1},\dots,V_{N}$ be non-empty open subsets of $\mathcal{A}$. Then we can find some non-zero $x\in O$ and non-zero $y_{l}\in V_{l}$ for each $l\in\{1,\dots,N\}$. There exists some $\alpha$ such that $p_{\alpha}^{3}x\in O$ and $p_{\alpha}^{3}y_{l}\in V_{l}\setminus\{0\}$ for each $l=1,\dots,N$. Clearly, this gives that $p_{\alpha}x\neq 0$ and $p_{\alpha}y_{l}\neq 0$ for each $l=1,\dots,N$. Also, since $O\setminus\{0\},\,V_{1}\setminus\{0\},\dots,V_{N}\setminus\{0\}$ are open and non-empty, there is some $\delta>0$ such that the $\delta$-neighbourhood of $p_{\alpha}^{3}x$ is contained in $O\setminus\{0\}$, and the $\delta$-neighbourhood of $p_{\alpha}^{3}y_{l}$ is contained in $V_{l}\setminus\{0\}$ for each $l=1,\dots,N$.
Let $C_{y}=\max_{1\leq l\leq N}\bigl\{\|y_{l}\|,\bigr\}$ and set
$\varepsilon:=\min\!\left\{\frac{\delta}{4\|x\|\|p_{\alpha}\|},\;\frac{\delta}{4N\sqrt{\|x\|C_{y}}\|p_{\alpha}\|}\frac{\delta}{4NC_{y}\|p_{\alpha}\|},\frac{\|p_{\alpha}^{3}x\|}{2\|p_{\alpha}x\|},\frac{\|p_{\alpha}^{3}y_{1}\|}{2\|p_{\alpha}y_{1}\|},\cdots,\frac{\|p_{\alpha}^{3}y_{N}\|}{2\|p_{\alpha}y_{N}\|}\right\}.$
Choose $F\in\mathcal{F}$ and the families $\{g_{t,1}\}_{t\in F},\dots,\{g_{t,N}\}_{t\in F},\{d_{t}\}_{t\in F}$ satisfying the assumptions in $(2)$ with respect to $\epsilon$ and $p_{\alpha}$.For each $\alpha$ and every $t\in F,l\in\{1,\dots,N\}$ we define the operator $S_{t,l,\alpha}$ on $\mathcal{A}$ by $S_{t,l,\alpha}(a)=\Phi_{t,l}^{-1}(b_{t,l,\alpha}^{-1})\Phi_{t,l}^{-1}(a)$ for all $a\in\mathcal{A}.$ We observe now that for each $t\in F$ and $l\in\{1,\dots,N\}$ it holds that
$\parallel T_{t,l}\!\bigl(d_{t}p_{\alpha}x\bigr)\parallel=\parallel\,b_{t,l}\,\Phi_{t,l}\bigl(d_{t}\bigr)\,\Phi_{t,l}(p_{\alpha}x)\,\parallel\leq\parallel\,b_{t,l}\,\Phi_{t,l}\bigl(d_{t}\bigr)\,\parallel\,\parallel\,\Phi_{t,l}(p_{\alpha}x)\,\parallel$
$\parallel\,\Phi_{t,l}^{-1}\,(b_{t,l})\,d_{t}\,\parallel\;\parallel\,x\,\parallel\|p_{\alpha}\|$
and
$\parallel S_{t,l,\alpha}\!\bigl(g_{t,l}p_{\alpha}\,y_{\ell}\bigr)\parallel=\parallel\,\Phi_{t,l}^{-1}\,(b_{t,l,\alpha}^{-1})\,\Phi_{t,l}^{-1}\!\bigl(g_{t,l}\bigr)\,\Phi_{t,l}^{-1}(p_{\alpha}\,y_{\ell})\,\parallel$
$\leq\parallel\,\Phi_{t,l}^{-1}\,(b_{t,l,\alpha}^{-1}\,g_{t,l})\,\parallel\,\parallel\,\Phi_{t,l}^{-1}(p_{\alpha}\,y_{\ell})\,\parallel=\parallel\,b_{t,l,\alpha}^{-1}\,g_{t,l}\,\parallel\;\parallel\,y_{\ell}\,\parallel\|p_{\alpha}\|.$
Moreover, for all $t\in F$ and each distinct $r,l\in\{1,\dots,N\}$ we have
$\parallel T_{t,r}\!\bigl(S_{t,l,\alpha}(g_{t,l}p_{\alpha}\,y_{\ell})\bigr)\parallel=\parallel\,b_{t,r}\,\Phi_{t,r}\,(\Phi_{t,l}^{-1}\,(b_{t,l,\alpha}^{-1})\bigr)\,\Phi_{t,r}\,\bigl(\Phi_{t,l}^{-1}\,(g_{t,l}p_{\alpha}\,y_{\ell}\bigr)\bigr)\,\parallel$
$\parallel\Phi_{t,r}^{-1}\,(b_{t,r}\,)\Phi_{t,l}^{-1}\,(b_{t,l,\alpha}^{-1})\,\Phi_{t,l}^{-1}\,(g_{t,l})\,\Phi_{t,l}^{-1}\,(p_{\alpha}\,y_{\ell})\,\parallel$
$\leq\parallel\Phi_{t,r}^{-1}\,(b_{t,r}\,)\Phi_{t,l}^{-1}\,(b_{t,l,\alpha}^{-1}g_{t,l})\,\,\parallel\,\parallel y_{\ell}\parallel\|p_{\alpha}\|.$
Finally, since $b_{t,l,\alpha}^{-1}$ is the $\alpha$-inverse of $b_{t,\ell},$ it follows that
$T_{t,l}\!\bigl(S_{t,l,\alpha}(g_{t,l}p_{\alpha}\,y_{\ell}))=g_{t,l}p_{\alpha}\,y_{\ell}$
for all $t\in F$ and each $\ell\in\{1,\dots,N\}.$
Since $\|d_{t}p_{\alpha}x-p_{\alpha}^{3}\,x\|<\|\epsilon\|\|p_{\alpha}x\|<\|p_{\alpha}^{3}\,x\|$, it follows that $d_{t}p_{\alpha}x\neq 0.$ Similarly, $g_{t,l}p_{\alpha}y_{\ell}\neq 0$ for all $t\in F$ and $\ell\in\{1,\dots,N\}$. Hence, since by the assumption we have that for each $t\in\mathcal{S}$ and $\ell\in\{1,\dots,N\}$ it holds that $b_{t,\ell}a=0$ if and only if $a=0,$ it follows that $T_{t,\ell}(d_{t}p_{\alpha}x)\neq 0$ for each $t\in F$ and $\ell\in\{1,\dots,N\}.$ Similarly, since $b_{t,l,\alpha}^{-1}$ is the $\alpha$-inverse of $b_{t,\ell},$ it folllows that $S_{t,\ell,\alpha}(g_{t,\ell}p_{\alpha}y_{\ell})\neq 0$ for all $t\in F$ and $\ell\in{1,\dots,N}.$
For each $t\in F$, we put
$x_{t}:=d_{t}p_{\alpha}x+\frac{\sqrt{\sum_{l=1}^{N}\,\parallel T_{t,l}(d_{t}p_{\alpha}x)\parallel}}{\sqrt{\sum_{l=1}^{N}\,\parallel S_{t,l,\alpha}(g_{t,l}p_{\alpha}y_{\ell})\parallel}\qquad}\,{\sum_{l=1}^{N}\,S_{t,r,\alpha}(g_{t,r}p_{\alpha}y_{r})}.$
By triangle inequality and choice of $\varepsilon$ one can deduce that for each $t\in F$, we have
$x_{t}\in O,\ \text{and}\ \frac{\sqrt{\sum_{\ell=1}^{N}\|S_{t,\ell,\alpha}(g_{t,\ell}p_{\alpha}y_{\ell})\|}}{\sqrt{\sum_{\ell=1}^{N}\|T_{t,\ell}(d_{t}p_{\alpha}x)\|}}T_{t,r}(x_{t})\in V_{r}.$
for all $r\in\{1,\dots,N\}.$
We notice that the assumption that $\{\Phi_{t,1}\}_{t\in\mathcal{S}},\cdots,\{\Phi_{t,N}\}_{t\in\mathcal{S}}$ is a disjoint aperiodic system is only needed for the proof of the implication $(1)\Rightarrow(2)$. Also, we notice that we have used the condition (R) only in the proof of the implication $(1)\Rightarrow(2)$ to obtain the desired inequalities.
The next corollary can easily be deduced from Theorem 2.1.
Corollary 2.2.
Let now $\mathcal{F}$ be the family of all infinite subsets of $\mathbb{N}.$ If there exist dense subsets $\mathcal{A}_{0},\mathcal{B}_{1},\cdots,\mathcal{B}_{N}$ of $\mathcal{A}$ and a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ such that for each $s,\ell\in\{1,\dots,N\},$ every fixed $\alpha$ and all $x\in\mathcal{A}_{0},y_{s}\in\mathcal{B}_{s}$ we have
$\lim_{k\to\infty}\|b_{n_{k},s,\alpha}^{-1}y_{s}\|\|\Phi_{n_{k},l}^{-1}(b_{n_{k},l})x\|=0,$
and, in addition, for each distinct $s,\ell\in\{1,\ldots,N\}$ it holds that
$\lim_{k\to\infty}\|\Phi_{n_{k},\ell}^{-1}(b_{n_{k},\ell})\Phi_{n_{k},s}^{-1}(b_{n_{k},s,\alpha}^{-1}y_{s})\|=0,$
then the families $\{T_{n,1}\}_{n\in\mathbb{N}},\cdots,\{T_{n,N}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
Remark 2.3.
The above results and the proofs can be extended and generalized to the case when $\mathcal{A}$ is a Banach bimodule over a unital normed algebra $\mathcal{A}_{1}.$ The condition (E) remains the same in this case, however, the condition $(\mathcal{P})$ should be then reformulated as follows:
Let $\{p_{\alpha}\}_{\alpha}$ be a set in $\mathcal{A}\cap\mathcal{A}_{1}$ satisfying that given any open subset $O$ of $\mathcal{A}$ there exists some $x\in\mathcal{A}\cup\mathcal{A}_{1}$ and some $p_{\alpha_{0}}\in\{p_{\alpha}\}_{\alpha}$ such that $p_{\alpha_{0}}x\in\mathcal{A}_{1}$ and $p_{\alpha_{0}}^{3}x\in O.$
Also, in this case, we let $\{\Phi_{t,1}\}_{t\in\mathcal{S}},\ \{\Phi_{t,2}\}_{t\in\mathcal{S}},\ \ldots,\ \{\Phi_{t,N}\}_{t\in\mathcal{S}}$ be families of isometric algebra isomorphisms of $\mathcal{A}_{1}$ and
$\{\Psi_{t,1}\}_{t\in\mathcal{S}},\ \{\Psi_{t,2}\}_{t\in\mathcal{S}},\ \ldots,\ \{\Psi_{t,N}\}_{t\in\mathcal{S}}$
be families of linear isometric isomorphisms of $\mathcal{A}$ such that for all $x,y\in\mathcal{A}_{1}$ and $a\in\mathcal{A}$ it holds that $\Psi_{t,j}(xay)=\Phi_{t,j}(x)\Psi_{t,j}(a)\Phi_{t,j}(y)$ for all $t\in\mathcal{S}$ and $j\in\{1,\dots,N\}.$ We express aperiodicity still in terms of the families
$\{\Phi_{t,1}\}_{t\in\mathcal{S}},\ \{\Phi_{t,2}\}_{t\in\mathcal{S}},\ \ldots,\ \{\Phi_{t,N}\}_{t\in\mathcal{S}},$ moreover, the conditions (R) and (C) remain the same in this case. Further, we define the operator
$T_{t,\ell}:\mathcal{A}\to\mathcal{A}\quad\text{by}\quad T_{t,\ell}(a)=b_{t,\ell}\,\Psi_{t,\ell}(a)$
for all $a\in\mathcal{A}.$ Under this set up and these assumptions, it is not hard to extend and generalize the proof of Theorem 2.1 in this setting.
Now we set $N=1$ and we will therefore write $b_{t}$, $\Phi_{t}$ and $T_{t}$ instead of $b_{t,1}$, $\Phi_{t,1}$ and $T_{t,1},$ respectively. We will now assume that $b_{t}$ is invertible for each $t\in\mathcal{S}$ and we will consider the map $S_{t}:\mathcal{A}\rightarrow\mathcal{A}$ given by $S_{t}(a)=\Phi_{t}^{-1}(b_{t}^{-1})\,\Phi_{t}^{-1}(a)$ for all $a\in\mathcal{A}$. Obviously, $S_{t}=T_{t}^{-1}$ for each $t\in\mathcal{S}.$
We set $C_{t}:=\frac{1}{2}(T_{t}+S_{t})$ for all $t\in\mathcal{S}$.
Proposition 2.4.
Under the above notation and assumptions, we have $(2)\Rightarrow(1)$.
(1) The family $\{C_{t}\}_{t\in\mathcal{S}}$ is $\mathcal{F}$–semi–transitive.
(2) For each fixed $p_{\alpha}$ and $\varepsilon>0$ there exists some $F\in\mathcal{F}$ and families $\{d_{t}\}_{t\in F}$, $\{g_{t}\}_{t\in F}$, $\{e_{t}\}_{t\in F}$, $\{f_{t}\}_{t\in F}$ with $e_{t}+f_{t}=g_{t}$ for all $t\in F$ such that
$\|d_{t}-p_{\alpha}^{2}\|<\varepsilon,\qquad\|g_{t}-p_{\alpha}^{2}\|<\varepsilon,$
$\|\Phi_{t}^{-2}(b_{t})\Phi_{t}^{-1}(b_{t})e_{t}\|<\varepsilon,$
$\|\Phi_{t}(b_{t}^{-1})\,b_{t}^{-1}\,f_{t}\|<\varepsilon,$
$\|\Phi_{t}^{-1}(b_{t})\,d_{t}\|\,\|\Phi_{t}^{-1}(b_{t})\,e_{t}\|<\varepsilon^{2},$
$\|\Phi_{t}^{-1}(b_{t})\,d_{t}\|\,\|b_{t}^{-1}f_{t}\|<\varepsilon^{2},$
$\|b_{t}^{-1}\,d_{t}\|\,\|\Phi_{t}^{-1}(b_{t})\,e_{t}\|<\varepsilon^{2},$
$\|b_{t}^{-1}\,d_{t}\|\,\|b_{t}^{-1}\,f_{t}\|<\varepsilon^{2},\text{ for all }t\in F.$
Proof.
Given non–empty open subsets $\mathcal{O}$ and $\mathcal{V}$ of $\mathcal{A}$, as in the proof of Theorem 2.1 part $(2)\Rightarrow(1)$ we can find some $\delta>0$ and some $p_{\alpha}^{3}x\in\mathcal{O}\setminus\{0\}$, $p_{\alpha}^{3}y\in\mathcal{V}\setminus\{0\}$ such that $\delta$–neighbourhoods of $p_{\alpha}^{3}x$ and $p_{\alpha}^{3}y$ are contained in $\mathcal{O}\setminus\{0\}$ and $V\setminus\{0\}$, respectively.
Set
$\varepsilon:=\min\left\{\frac{\delta}{4\|p_{\alpha}\|(\|x\|+\|y\|)},\;\frac{\delta}{16\|p_{\alpha}\|\sqrt{(\|x\|\|y\|)}},\;\frac{\|p_{\alpha}^{3}x\|}{2\|p_{\alpha}x\|},\;\frac{\|p_{\alpha}^{3}y\|}{2\|p_{\alpha}y\|}\right\}.$
Choose $F\in\mathcal{F}$ and families $\{d_{t}\}_{t\in F}$, $\{g_{t}\}_{t\in F}$ , $\{e_{t}\}_{t\in F}$, $\{f_{t}\}_{t\in F}$ satisfying the assumptions in (2) with respect to $p_{\alpha}$ and $\varepsilon$. Since $e_{t}+f_{t}=g_{t}$ and $g_{t}\,p_{\alpha}y\neq 0$ (as shown in the proof of Theorem 2.1 part $(2)\Rightarrow(1)$), it follows that either $e_{t}\,p_{\alpha}y\neq 0$ or $f_{t}\,p_{\alpha}y\neq 0$. Then, by the same arguments as in the proof of Theorem 2.1 part $(2)\Rightarrow(1)$ we can show that either $T_{t}(e_{t},p_{\alpha}y)\neq 0$ or $S_{t}(f_{t},p_{\alpha}y)\neq 0$.
Further, we notice that for each $t\in F$ it holds that
$\|T_{t}^{2}(e_{t},p_{\alpha}y)\|=\|b_{t}\,\Phi_{t}\,(b_{t})\,\Phi_{t}^{2}\,(e_{t}\,p_{\alpha}y)\|$
$\leq\|\Phi_{t}^{-2}\,(b_{t})\,\Phi_{t}^{-1}\,(b_{t})\,e_{t}\|\,\|p_{\alpha}\|\,\|y\|$
and,
$\|S_{t}^{2}(f_{t},p_{\alpha}y)\|=\|\Phi_{t}^{-1}(b_{t}^{-1})\cdot\Phi_{t}^{-2}(b_{t}^{-1}f_{t}\,p_{\alpha}y)\|\leq\|\Phi_{t}(b_{t}^{-1})b_{t}^{-1}f_{t}\|\,\|p_{\alpha}\|\,\|\|y\|.$
For each $t\in F,$ put
$v_{t}:=d_{t}p_{\alpha}x+\frac{2\sqrt{\|T_{t}(d_{t}p_{\alpha}x)\|+\|S_{t}(d_{t}p_{\alpha}x)\|}}{\sqrt{\|T_{t}(e_{t}+p_{\alpha}y)\|+\|S_{t}(f_{t}p_{\alpha}y)\|}}(T_{t}(e_{t}+p_{\alpha}x)+S_{t}(e_{t}p_{\alpha}x)).$
By the above inequalities and inequalities in the proof of Theorem 2.1 part $(2)\Rightarrow(1)$, the choice of $\varepsilon$ and the triangle inequality, we can deduce that $v_{t}\in\mathcal{O}$ and
$\frac{\sqrt{\|T_{t}(e_{t}p_{\alpha}y)\|+\|S_{t}(f_{t}p_{\alpha}y)\|}}{\sqrt{\|T_{t}(d_{t}+p_{\alpha}x)\|+\|S_{t}(d_{t}p_{\alpha}x)\|}}C_{t}(v_{t})\in\mathcal{V}\text{ for all }t\in F.$
By putting either $e_{t}=0$ (so that $f_{t}=g_{t}$), or by putting $f_{t}=0$ (so that $e_{t}=g_{t}$) in Proposition 2.4, then we can get the following corollary.
Corollary 2.5.
Let now $\mathcal{F}$ be the family of all infinite subsets of $\mathbb{N}$. If there exist dense subsets $\mathcal{B}_{1},\mathcal{B}_{2}$ of $\mathcal{A}$ and a strictly increasing sequence $\{n_{k}\}\subseteq\mathbb{N}$ such that for every fixed $p_{\alpha}$ it holds that either and
$\lim_{k\to\infty}\|\Phi_{n_{k}}^{-2}(b_{n_{k}})\,\Phi_{n_{k}}^{-1}(b_{n_{k}})\,y\|=\lim_{k\to\infty}\|b^{-1}_{n_{k}}x\|\,\|\Phi_{n_{k}}^{-1}(b_{n_{k}})y\|=$
$\lim_{k\to\infty}\|\Phi_{n_{k}}^{-1}(b_{n_{k}})x\|\,\|\Phi_{n_{k}}^{-1}(b_{n_{k}})y\|=0$
or
$\lim_{k\to\infty}\|\Phi_{n_{k}}(b_{n_{k}}^{-1})b_{n_{k}}^{-1}y\|=\lim_{k\to\infty}\|b_{n_{k}}^{-1}x)\|\cdot\|b_{n_{k}}^{-1}y\|=\lim_{k\to\infty}\|\Phi_{n_{k}}^{-1}(b_{n_{k}})\|\|b_{n_{k}}^{-1}y\|=0$
for all $x\in\mathcal{B}_{1}$ and $y\in\mathcal{B}_{2}$, then the family $\{C_{n}\}_{n\in\mathbb{N}}$ is $\mathcal{F}$–semi-transitive.
3. Applications
From now on and in the rest of this section we will assume that $\mathcal{F}$ the family of all infinite subsets of $\mathbb{N}.$ Further, in this section, for a separable Hilbert space $H$ we let $B(H)$ be the space of all bounded linear operators on $H$, and $\mathcal{C}$ be the $C^{*}$-algebra of compact operators on $H$. For an orthonormal basis $\{e_{j}\}_{j\in\mathbb{Z}}$ for $H$, we let for each $m\in\mathbb{N}$, $P_{m}$ be the orthogonal projection onto $\text{Span}\{e_{-m},\dots,e_{m}\}$.
Moreover, we will denote by $\Omega$ be a locally compact Hausdorff, non-compact space. As usual, $C_{b}(\Omega)$ will denote the space of all bounded continuous functions on $\Omega$ and $C_{0}(\Omega)$ will denote the space of all continuous functions on $\Omega$ vanishing at infinity. Both these spaces will be equipped with supremum norm. In addition, we let $C_{c}(\Omega)$ is the space of all continuous compactly supported functions on $\Omega.$ For each pair of compact subset $K_{1}$ and $K_{2}$ of $\Omega$ with $K_{1}\subseteq K_{2}$ , we let $u_{(K_{1},K_{2})}\in C_{0}(\Omega)$ such that $u_{{(K_{1},K_{2})}|_{K_{1}}}=1$, $\operatorname{supp}u_{(K_{1},K_{2})}=K_{2}$ and $0\leq u_{(K_{1},K_{2})}\leq 1$.
3.1. Applications to generalized weighted shifts on Hilbert C*-modules
In this subsection, we will denote by $\ell_{2}(\mathcal{C})$ the standard (right) Hilbert module over $\mathcal{C}$, see [21, Example 1.3.5]. Notice that $\ell_{2}(\mathcal{C})$ is a Banach algebra. Indeed, we can define multiplication on $\ell_{2}(\mathcal{C})$ as pointwise multiplication, i.e., if $\{x_{j}\}_{j\in\mathbb{Z}},\{y_{j}\}_{j\in\mathbb{Z}}\in\ell_{2}(\mathcal{C})$, then
$\{x_{j}\}_{j\in\mathbb{Z}}\cdot\{y_{j}\}_{j\in\mathbb{Z}}=\{x_{j}y_{j}\}_{j\in\mathbb{Z}}.$
To see that $\{x_{j}y_{j}\}_{j\in\mathbb{Z}}$ belongs to $\ell_{2}(\mathcal{C})$, it suffices to observe that
$\sum_{j}y_{j}^{*}x_{j}^{*}x_{j}y_{j}\leq\sum_{j}y_{j}^{*}\|x_{j}\|^{2}y_{j}\leq\parallel\{x_{j}\}_{j\in\mathbb{Z}}\parallel^{2}\sum y_{j}^{*}y_{j}.$
For each $m,J\in\mathbb{N},$ we let $\tilde{p}_{J,m}\in\ell_{2}(\mathcal{C})$ be given by
[…]
As shown in the proof of [16, Proposition 3.1], the set $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ forms a left approximate unit for $\ell_{2}(\mathcal{C})$. We put then $\mathcal{A}=\ell_{2}(\mathcal{C}),$ and
$\mathcal{A}_{1}:=\{\{y_{j}\}_{j\in\mathbb{Z}}\mid y_{j}\in B(H)\ \forall j,\,\exists M_{y}>0\text{ such that }\|y_{j}\|\leq M_{y}\ \forall j\}.$
The multiplication on $\mathcal{A}_{1}$ is defined similarly component-wise. We define the norm $\|\cdot\|_{1}$ on $\mathcal{A}_{1}$ as $\|\{y_{j}\}_{j\in\mathbb{Z}}\|_{1}=sup_{j\in\mathbb{Z}}\|y_{j}\|$ for all $\{y_{j}\}_{j\in\mathbb{Z}}\in\mathcal{A}_{1}.$ It is straightforward to check that the condition (E) is satisfied in this case.
Let now $\mathcal{U}$ be a unitary operator on $H$. For each $n\in\mathbb{N}$ we define the map $\Phi_{n}:\mathcal{A}_{1}\to\mathcal{A}_{1}$ by
$\Phi_{n}\left(\{y_{j}\}_{j\in\mathbb{Z}}\right)=\{\mathcal{U}^{-n}y_{j-n}\mathcal{U}^{n}\}_{j\in\mathbb{Z}},$
for all $\{y_{j}\}_{j\in\mathbb{Z}}\in\mathcal{A}_{1},$ (that is the $j$-th coordinate of $\Phi_{n}(\{y_{j}\}_{j\in\mathbb{Z}})=\mathcal{U}^{-n}y_{j-n}\mathcal{U}^{n}$ for all $j\in\mathbb{Z}$).
Now, the set $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ forms a left approximate unit for $\mathcal{A}=\ell_{2}(\mathcal{C})$ and $\tilde{p}_{J,m}=\tilde{p}_{J,m}^{3}$ for all $(J,m)\in\mathbb{N}^{2}.$ Moreover, by some calculations it can be checked that $\{\Phi_{n}\}_{n\in\mathbb{N}}$ is aperiodic with respect to the left approximate unit $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ and that the condition (R) is satisfied in this case.
For each $n\in\mathbb{N},$ let $\{W_{n,j}\}_{j\in\mathbb{Z}}$ be a family of operators in $B(H)$ which is uniformly bounded in norm such that each $W_{n,j}$ has a bounded inverse and such that $\{W_{n,j}^{-1}\}_{j\in\mathbb{Z}}$ is also uniformly bounded in the norm. Then, for each $n\in\mathbb{N}$ we let $b_{n}\in\mathcal{A}_{1}$ be given by
$(b_{n})_{i}=W_{n,i}\mathcal{U}^{n},$
for all $i\in\mathbb{Z}.$ Obviously, $b_{n}$ is invertible in $\mathcal{A}_{1}$ for each $n\in\mathbb{N}$ and
$(b_{n}^{-1})_{i}=\mathcal{U}^{-n}\,W^{\,-1}_{n,i},$
for all $i\in\mathbb{Z},$ hence, the condition $(C)$ is satisfied in this case.
Finally, for each $n\in\mathbb{N}$ we let $T_{n}$ be the operator on $\mathcal{A}$ given by
$T_{n}(a)=b_{n}\,\Phi_{n}(a),\qquad\text{for all }a\in\mathcal{A}.$
Noticing that for each $a\in\mathcal{A}$ and $j\in\mathbb{Z}$ it holds that $\|a_{j}\|\leq\|a\|$, in view of Corollary 2.5 and by arguments similar to those in the proofs of [16, Theorem 3.2] and [17, Proposition 2.7] it is not hard to deduce the following proposition in this special case.
Proposition 3.1.
Under the above notation and assumptions, if there exist dense subsets $H_{0}$ and $H_{1}$ of $H$ and a strictly increasing sequence $\{n_{k}\}_{k}$ of natural numbers such that either
$\lim_{k\to\infty}\|W^{-1}_{n_{k},j}x\|\,\|W_{n_{k,j+n_{k}}}y\|=\lim_{k\to\infty}\|W_{n_{k},j+n_{k}}x\|\,\|W_{n_{k},j+n_{k}}y\|$
$\lim_{k\to\infty}\|W_{n_{k},j+2n_{k}}W_{n_{k},j+n_{k}}y\|=0,$
or
$\lim_{k\to\infty}\|W_{n_{k},j+n_{k}}x\|\,\|W^{-1}_{n_{k},j}y\|=\lim_{k\to\infty}\|W^{-1}_{n_{k},j}x\|\,\|W^{-1}_{n_{k},j}y\|=$
$\lim_{k\to\infty}\|W^{-1}_{n_{k},j-n_{k}}W^{-1}_{n_{k},j}y\|=0,$
for all $x\in H_{0}$ and $y\in H_{1}$, then $\{C_{n}\}_{n\in\mathbb{N}}$ is $\mathcal{F}$–semi-transitive on $\mathcal{A}$.
Example 3.2.
Let $H=L^{2}(\mathbb{R})$. As in [16, Example 3.3], choose some $\{w_{j}\}_{j\in\mathbb{Z}}\subseteq L^{\infty}(\mathbb{R})$ such that $w_{j}>0$, $w_{j}^{-1}\in L^{\infty}(\mathbb{R})$ and $\parallel w_{j}\parallel,\parallel w_{j}^{-1}\parallel\leq M$ for all $j\in\mathbb{Z}$ and some $M>1$. Let $\{c_{j}\}_{j\in\mathbb{Z}}\subseteq\mathbb{R}^{+}$ be such that $c_{j}\geq K$ for all $j\in\mathbb{Z}$ and some $K>0$. For each $j\in\mathbb{Z},$ let $\alpha_{j}$ be the translation on $\mathbb{R}$ given by $\alpha_{j}(t)=t-c_{j}$ for all $t\in\mathbb{R},$ and let then $V_{j}\in B(H)$ be given by
$V_{j}(f)=w_{j}\cdot(f\circ\alpha_{j})$
for all $f\in H.$ For each $n\in\mathbb{N},i\in\mathbb{Z}$, let $W_{n,i}\in B(H)$ be given by
$W_{n,i}:=V_{i}V_{i-1}\cdots V_{i-n+1}.$
If there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}\leq 1-\varepsilon,w_{j}\chi_{(-\infty,0)}=1$ for all $j\geq 0$ and $w_{j}=1$ for all $j<0$; then by the same arguments as in [16, Example 3.3] one can show that the conditions of Proposition 3.1 are satisfied with $H_{0}=H_{1}=C_{c}(\mathbb{R})$ in this case. On the other hand, if there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}=1,w_{j}\chi_{(-\infty,0)}\geq 1+\varepsilon$ for all $j<0$ and $w_{j}=1$ for all $j\geq 0,$ then again the conditions of Proposition 3.1 will be satisfied with $H_{0}=H_{1}=C_{c}(\mathbb{R}).$
Next, for $N\in\mathbb{N},$ let $r_{1},\dots,r_{N}\in\mathbb{N}$ with $r_{1}<r_{2}<\dots<r_{N}.$
Let $\alpha_{1},\dots,\alpha_{N}$ be homeomorphisms of $\Omega$, $\{w_{j}^{(1)}\}_{j\in\mathbb{Z}},\dots,\{w_{j}^{(N)}\}_{j\in\mathbb{Z}}\subseteq C_{b}(\Omega)$. We will assume that $w_{j}^{(\ell)}>0$ for all $j\in\mathbb{Z}$ and $\ell\in\{1,\dots,N\}$ and that $\|w_{j}^{(\ell)}\|_{\infty}<M$ for all $j\in\mathbb{Z}$, $\ell\in\{1,\dots,N\}$ and some $M>0.$ Now we consider $\mathcal{A}=\ell_{2}(C_{0}(\Omega))$, and
$\mathcal{A}_{1}:=\{\{y_{j}\}_{j\in\mathbb{Z}}\mid y_{j}\in C_{b}(\Omega)\ \forall j,\,\exists M_{y}>0\text{ such that }\|y_{j}\|_{\infty}\leq M_{y}\ \forall j\}.$
The multiplication and the norm on $\mathcal{A}_{1}$ are defined in the same way as in the previous case. Once again, it is straightforward to check that the condition (E) is satisfied in this case.
For $J\in\mathbb{N}$, we let $\tilde{p}_{(K_{1},K_{2},J)}\in\mathcal{A}$ be given by
[…]
Then it is easily checked that $\{\tilde{p}_{(K_{1},K_{2},J)}^{3}\}_{K_{1}\subseteq K_{2}\subseteq\Omega,\,K_{1},K_{2}\text{ compact ,}J\in\mathbb{N}}$ satisfies the condition $(\mathcal{P}.)$
For each $\ell\in\{1,\dots,N\}$ and $k\in\mathbb{N}$, we let $\tilde{\Phi}_{k,\ell}:\mathcal{A}\to\mathcal{A}$ be given by
$\tilde{\Phi}_{k,\ell}\big(\{f_{j}\}_{j\in\mathbb{Z}}\big)=\{f_{j-k}\circ\alpha_{\ell}^{k}\,\}_{j\in\mathbb{Z}},$
for all $\{f_{j}\}_{j\in\mathbb{Z}}\subset\mathcal{A}$. Let now $r_{1},\dots,r_{N}\in\mathbb{N}$ with $r_{1}<r_{2}<\dots<r_{N}$. For each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\}$, set
$\Phi_{n,\ell}=\tilde{\Phi}_{r_{\ell}n,\ell}.$
Clearly, the system of isometric isomorphisms $\{\Phi_{k,1}\}_{k\in\mathbb{N}},\dots,\{\Phi_{k,N}\}_{k\in\mathbb{N}}$ is disjoint aperiodic with respect to
$\{\tilde{p}_{(K_{1},K_{2},J)}\}_{K_{1}\subseteq K_{2}\subseteq\Omega,\,K_{1},K_{2}\text{ compact ,}J\in\mathbb{N}}.$
Also, the condition (R) is satisfied in this case.
For each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\}$, let $b_{n,\ell}\in\mathcal{A}_{1}$ be given by
$(b_{n,\ell})_{j+r_{\ell}n}=\prod_{k=1}^{r_{\ell}n}\big(w_{j+k}^{(\ell)}\circ\alpha_{\ell}^{n-k}\big)\quad\text{for all }j\in\mathbb{Z}.$
Since for each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\},$ the function $w_{j}^{(\ell)}>0,$ it follows that the condition $(C)$ is satisfied in this case.
Now, for each $K_{2}\subseteq\Omega$ compact, we can find some $v_{K_{2}}\in C_{0}(\Omega)$ with $0\leq v_{K_{2}}\leq 1$ such that ${v_{K_{2}}}_{|_{K_{2}}}=1.$ Recalling that the approximate unit for $\mathcal{A}$ in this case is indexed over all triples $(K_{1},K_{2},J)$ with $K_{1}\subseteq K_{2}\subseteq\Omega,\,K_{1},K_{2}\text{ compact, }J\in\mathbb{N},$ for each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\}$, and every fixed triple $(K_{1},K_{2},J)$ we let $b_{n,\ell,(K_{1},K_{2},J)}^{-1}\in\mathcal{A}_{1}$ be given by
$(b_{n,\ell,(K_{1},K_{2},J)}^{-1})_{j+r_{\ell}n}=\frac{v_{K_{2}}}{\prod_{k=1}^{r_{\ell}n}\big(w_{j+k}^{(\ell)}\circ\alpha_{\ell}^{n-k}\big)}$
for all $j\in\mathbb{Z}$ with $J-r_{\ell}n\leq j\leq J-r_{\ell}n$ and $(b_{n,\ell,(K_{1},K_{2},J)}^{-1})_{j+r_{\ell}n}=0$ else. Then, for all $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\},$ we let $T_{n,\ell}:\mathcal{A}\to\mathcal{A}$ be given by
$T_{n,\ell}(a)=b_{n,\ell}\,\Phi_{n,\ell}(a)\quad\text{for all }a\in\mathcal{A}.$
The next corollary can be deduced by some calculations from Theorem 2.1.
Corollary 3.3.
Under the above notation and assumptions, the following statements are equivalent:
(1) The sequences of operators
$\{T_{n,1}\}_{n\in\mathbb{N}},\ \dots,\ \{T_{n,N}\}_{n\in\mathbb{N}}$
are $d\mathcal{F}$-semi-transitive on $\mathcal{A}$.
(2) For every $J\in\mathbb{N}$ and each compact subset $K$ of $\Omega$, there exists a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ such that
$\lim_{k\to\infty}\Bigg(\sup_{t\in K}\frac{1}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j+i-r_{s}n_{k}}^{(s)}\circ\alpha_{s}^{r_{s}n_{k}-i}\big)(t)}\Bigg)\cdot\Bigg(\sup_{t\in K}\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j+i}^{(\ell)}\circ\alpha_{\ell}^{-i}\big)(t)\Bigg)=0$
for all $j\in\mathbb{Z}$ with $J\leq j\leq J$ and every $l,s\in\{1,\dots,N\}$. Moreover, for each distinct $s,\ell\in\{1,\dots,N\}$ and all $j\in\mathbb{Z}$ with $J\leq j\leq J$, it holds that
$\lim_{k\to\infty}\Bigg(\sup_{t\in K}\frac{\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j-r_{\ell}n_{k}-i}^{(\ell)}\circ\alpha_{\ell}^{-i}\circ\alpha_{s}^{r_{s}n_{k}}\big)(t)}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j-r_{s}n_{k}+i}^{(s)}\circ\alpha_{s}^{n_{k}-i}\big)(t)}\Bigg)=0.$
Example 3.4.
Let $N=1,r_{1}=1$ and choose some $\{w_{j}\}_{j\in\mathbb{Z}}\subseteq L^{\infty}(\mathbb{R})$ such that $w_{j}>0$, $w_{j}^{-1}\in L^{\infty}(\mathbb{R})$ and $\parallel w_{j}\parallel\leq M$ for all $j\in\mathbb{Z}$ and some $M>1$. Let $\alpha(t)=t-1$ for all $t\in\mathbb{R}.$ If there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}\leq 1-\varepsilon,w_{j}\chi_{(-\infty,0)}=1$ for all $j\geq 0$ and $w_{j}=1$ for all $j<0$; or if there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}=1,w_{j}\chi_{(-\infty,0)}\geq 1+\varepsilon$ for all $j<0$ and $w_{j}=1$ for all $j\geq 0,$ then the conditions of Corollary 3.3 are satisfied in this case.
3.2. Applications to generalized weighted composition operators on spaces of operator-valued continuous functions
In this subsection, we let $C_{0}(\Omega,\mathcal{C})$ denote the non-unital normed algebra of all continuous $\mathcal{C}$-valued functions on $\Omega$ vanishing at infinity and equipped with the norm
$\|f\|_{\infty}:=\sup_{x\in\Omega}\|f(x)\|$
and $C_{b}(\Omega,B(H))$ be the unital normed algebra consisting of all uniformly bounded $B(H)$-valued continuous functions on $\Omega$ equipped with the same norm $\|\cdot\|_{\infty}$.
Obviously, $C_{0}(\Omega,\mathcal{C})$ is (left) ideal in $C_{b}(\Omega,B(H))$ and the condition $(E)$ is satisfied. We will therefore denote $C_{0}(\Omega,\mathcal{C})$ and $C_{b}(\Omega,B(H))$ by $\mathcal{A}$ and $A_{1},$ respectively. For every triple $(K_{1},K_{2},m)$ with $K_{1}\subseteq K_{2}\subseteq\Omega,\,K_{1},K_{2}\text{ compact, }m\in\mathbb{N},$ let $\tilde{P}_{(K_{1},K_{2},m)}\in C_{0}(\Omega,\mathcal{C})$ be given by
$\tilde{P}_{(K_{1},K_{2},m)}(x)=u_{(K_{1},K_{2})}(x)P_{m}\quad\text{for all }x\in\Omega.$
We have the following lemma.
Lemma 3.5.
The set $\{\tilde{P}_{(K_{1},K_{2},m)}\}_{K_{1}\subseteq K_{2}\in\Omega,K_{1},K_{2}\text{ compact, }m\in\mathbb{N}}$ satisfies the condition $(\mathcal{P})$.
Proof.
Let $O$ be a non-empty open subset of $\mathcal{A}$ and $f\in O$. Then there exists some $\varepsilon>0$ such that $\varepsilon$– neighbourhood of $f$ is contained in $O$ as well. Since $f\in\mathcal{A},$ there exists some compact subset $K_{1}$ of $\Omega$ such that $\|f(x)\|<\frac{\varepsilon}{2}$ for all $x\in\Omega\setminus K_{1}$. Now, by continuity of $f,$ for each $x\in K_{1}$ we can find some open set $O_{x}$ of $\Omega$ containing $x$ such that $\|f(u)-f(x)\|<\frac{\varepsilon}{2}$ whenever $y\in O_{x}$. Then for all $m\in\mathbb{N}$ and each $y\in O_{x}$ we obtain
$\|(I-P_{m})f(y)\|\leq\|(I-P_{m})(f(y)-f(x))\|+\|(I-P_{m})f(x)\|$
$\leq\|f(y)-f(x)\|+\|(I-P_{m})f(x)\|<\frac{\varepsilon}{2}+\|(I-P_{m})f(x)\|$
By [21, Proposition 2.2.1], since $f(x)\in\mathcal{C},$ we can find sufficiently large $m_{x}\in\mathbb{N}$ such that $\|(I-P_{m})f(x)\|<\frac{\varepsilon}{2}$ for all $m\geq m_{x}.$ Hence, $\|(I-P_{m})f(y)\|<\varepsilon$ each $y\in O_{x}$ and all $m\geq m_{x}.$
The collection $\{O_{x}\}_{x\in K_{1}}$ is an open cover of $K_{1}$, hence there exist a finite subcover $\{O_{x_{1}},\ldots,O_{x_{M}}\}.$ Let $\tilde{m}:=\max\{m_{x_{1}},\ldots,m_{x_{M}}\}.$ Then, by the above we get
$\|\tilde{P}^{3}_{(K_{1},K_{2},m)}(x)f(x)-f(x)\|=\|(I-P_{m})f(x)\|<\varepsilon$ for all $x\in K_{1}$ and $m\geq\tilde{m}$. If $x\in\Omega\setminus K_{1},$ then for all $m\in\mathbb{N}$ we get
$\|\tilde{P}^{3}_{(K_{1},K_{2},m)}(x)f(x)\|+\|f(x)\|=\|u^{3}_{(K_{1},K_{2})}(x)P_{m}f(x)\|+\|f(x)\|\leq 2\|f(x)\|<2\frac{\varepsilon}{2}=\varepsilon.$
Thus,
$\|\tilde{P}^{3}_{(K_{1},K_{2},\tilde{m})}f-f\|_{\infty}<\varepsilon,\text{ so }\tilde{P}^{3}_{(K_{1},K_{2},m)}f\in O.$
Now we consider families $\{\alpha_{n,1}\}_{n\in\mathbb{N}}\dots,\{\alpha_{n,N}\}_{n\in\mathbb{N}}$ of homeomorphisms of $\Omega.$ We will assume that for every compact subset $K$ of $X$ and every $S\in\mathcal{F}$ there exists some $F\in\mathcal{F}$ with $F\subseteq S$ such that
$K\cap\alpha_{n,\ell}(K)=\varnothing\quad\text{and}\quad\alpha_{n,\ell}^{-1}(\alpha_{n,s}(K))\cap K=\varnothing$
for all $n\in F$ and each distinct $s,\ell\in\{1,\dots,N\}.$ Also, we consider families of unitary operators $\{U_{n,1}\}_{n\in\mathbb{N}}\dots,\{U_{n,N}\}_{n\in\mathbb{N}}$ on $H.$ For every $n\in\mathbb{N}$ and $l\in\{1,\ldots,N\}$ we let then
$\Phi_{n,l}:\mathcal{A}_{1}\rightarrow\mathcal{A}_{1}\text{ be given by }\Phi_{t,l}(f):=U_{n,l}^{*}(f\circ\alpha_{n,l})U_{n,l}$
for all $f\in\mathcal{A}.$ Obviously the system $\{\Phi_{n,1}\}_{n\in\mathbb{N}},\dots,\{\Phi_{n,N}\}_{n\in\mathbb{N}}$ is a system of isometric isomorphisms on $\mathcal{A}$ which is disjoint aperiodic with respect to $\{\tilde{P}_{(K_{1},K_{2},m)}\}$ and, clearly, the condition ($R$) is satisfied. Let $\{w_{n,1}\}_{n\in\mathbb{N}},\dots,\{w_{n,N}\}_{n\in\mathbb{N}}$ be families in $\mathcal{A}_{1}$ such that $w_{n,l}(x)$ is invertible for all $x\in\Omega$, $n\in\mathbb{N}$, $l\in\{1,\ldots,N\}$, and such that $\{w^{-1}_{n,1}\}_{n\in\mathbb{N}},\dots,\{w^{-1}_{n,N}\}_{n\in\mathbb{N}}$ also belong to $\mathcal{A}_{1}$ i.e.
$\sup_{x\in\Omega}\|w^{-1}_{n,l}(x)\|<\infty\quad\text{for all }n\in\mathbb{N}\text{ and }l\in\{1,\ldots,N\}.$
For each $n\in\mathbb{N}$, and $l\in\{1,\ldots,N\},$ let then $b_{n,l}:=w_{n,l}U_{n,l}$. Since each $b_{n,l}$ is invertible in $\mathcal{A}_{1}$, it follows that the condition $(C)$ is satisfied in this case.
Corollary 3.6.
Under the above notation and assumptions, if there exist dense subsets $H_{0},H_{1},\ldots,H_{N}$ of $H$ such that for each $s,l\in\{1,\ldots,N\}$ and every compact subset $K$ of $\Omega$ we have that
$\lim_{n\to\infty}\left[\left(\sup_{x\in K}\|w_{n,s}^{-1}(x)h_{s}\|\right)\cdot\left(\sup_{y\in K}\|(w_{n,\ell}\circ\alpha_{n,\ell}^{-1})(y)h_{0}\|\right)\right]=0$
for every $h_{s}\in H_{s}$ and $h_{0}\in H_{0},$ and, in addition, for each distinct $s,l\in\{1,\ldots,N\}$ and every $h_{s}\in H_{s}$ we have
$\lim_{n\to\infty}\left(\sup_{x\in K}\|(w_{n,l}\circ\alpha_{n,l}^{-1}\circ\alpha_{n,s})(x)\,w_{n,s}^{-1}(x)\,h_{s}\|\right)=0,$
then the families $\{T_{n,1}\}_{n\in\mathbb{N}},\ldots,\{T_{n,N}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
Proof.
For each $s\in\{1,\ldots,N\}$, the set consisting of elements of the form
$u_{(K_{1},K_{2})}G_{s,m},$
where $G_{s,m}\in\mathcal{C}$ is such that
$G_{s,m}e_{i}\in H_{s}\text{ for all }j\in\{-m,\ldots,m\}\text{ and }G_{s,m}e_{j}=0\text{ for }j\in\mathbb{Z}\setminus\{-m,\ldots,m\}$
is obviously dense in $\mathcal{A}$ by Lemma 3.5 and the density of $H_{s}$ in $H$. (Here of course $m$ runs through the whole $\mathbb{N}$ and $(K_{1},K_{2})$ runs through all pairs of compact subsets $K_{1},K_{2}$ of $\Omega$ with $K_{1}\subseteq K_{2}.$) Let us denote this set by $\mathcal{A}_{s}$ for each $s\in\{1,\ldots,N\}$. Similarly, we let $\mathcal{A}_{0}$ be the set consisting of all elements of the form $u_{(K_{1},K_{2})}D_{m}$ where $D_{m}\in\mathcal{C}$ is such that
$D_{m}e_{j}\in H_{0}\text{ for all }j\in\{-m,\dots m\}\quad\text{and}\quad D_{m}e_{j}=0\text{ for }j\in\mathbb{Z}\setminus\{-m,\dots m\}.$
By some calculations similar to those in the proof of [16, Theorem 3.2] and [17, Proposition 2.7], it is then not hard to deduce that the conditions of Corollary 2.2 are satisfied. ∎
Example 3.7.
[…]
3.3. Applications to generalized weighted shifts on function sequence spaces
In this subsection, we assume that $\Omega=\mathbb{R}.$ and we let now $\mathcal{C}=C_{0}(\mathbb{R}).$ As in [15], for $\tau\in C_{b}(\mathbb{R}),$ we put
$\mathcal{C}_{\tau}:=\{f\in\mathcal{C}:\,\sum_{k=0}^{\infty}\|f\tau^{k}\|_{\infty}<\infty\}.$
For each $f\in\mathcal{C}_{\tau}$ we define
$\|f\|_{\tau}:=\sum_{k=0}^{\infty}\|f\tau^{k}\|_{\infty}.$
Then, $\mathcal{C}_{\tau}$ is a Banach algebra [14]. We will call this algebra Segal algebra corresponding to $\tau$.
As observed in [15], for a homeomorphism $\alpha$ of $\mathbb{R},$ we consider corresponding Segal algebra $\mathcal{C}_{\tau\circ\alpha^{k}}$ for all $k\in\mathbb{Z}$. Clearly, $f\cdot(\tau\circ\alpha^{k})\in\mathcal{C}$ for all $k\in\mathbb{Z}$ whenever $f\in C_{0}(\mathbb{R})$ since $\tau\circ\alpha^{k}\in C_{b}(\mathbb{R})$ for all $k\in\mathbb{Z}.$ By definition, $\|f\|_{\infty}\leq\|f\|_{\tau\circ\alpha^{k}}$ whenever $f\in\mathcal{C}_{\tau\circ\alpha^{k}}.$ We recall now the following lemmas from [15].
Lemma 3.8.
[15, Lemma 3.2] Let $\tau\in C_{b}(\mathbb{R})$ and $\epsilon_{1},\epsilon_{2}\in(0,1)$ with $\epsilon_{2}<\epsilon_{1}.$ Then for every compact subset $K_{\epsilon_{2}}^{(\tau)}\subseteq|\tau|^{-1}(\left[0,\epsilon_{2}\right]),$ there exists a function $\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}\in\mathcal{C}_{\tau}$ such that the followings hold:
  1. (1)
    $\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}=1$ on $K_{\epsilon_{2}}^{(\tau)}$,
  2. (2)
    $\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}\in C_{c}(\mathbb{R}),$
  3. (3)
    ${\rm supp}\,\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}\subseteq|\tau|^{-1}(\left[0,\epsilon_{1}\right])).$
Lemma 3.9.
[15, Lemma 3.3] For $\tau\in C_{b}(\mathbb{R})$ and each $f\in\mathcal{C}_{\tau}$, there is a sequence $(g_{n})$ in the set $\{\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}:\,\,0<\epsilon_{2}<\epsilon_{1}<1,\,\,K_{\epsilon_{2}}^{(\tau)}\text{ is compact}\}$ such that $fg_{n}\rightarrow f$ in $\mathcal{C}_{\tau}$.
Lemma 3.10.
[15, Lemma 3.5] For each $k\in\mathbb{Z}$, $f\circ\alpha\in\mathcal{C}_{\tau\circ\alpha^{k+1}}$ if and only if $f\in\mathcal{C}_{\tau\circ\alpha^{k}}$. Moreover, in this case, $\|f\circ\alpha\|_{\tau\circ\alpha^{k+1}}=\|f\|_{\tau\circ\alpha^{k}}$.
As in [15], let
${\rm c}_{0}^{\mathcal{C},\tau}:=\Big\{s:=(s_{k})\in\prod_{k\in\mathbb{Z}}\mathcal{C}_{\tau\circ\alpha^{k}}:\,\lim_{k\rightarrow\infty}\|s_{k}\|_{\tau\circ\alpha^{k}}=\lim_{k\rightarrow\infty}\|s_{-k}\|_{\tau\circ\alpha^{-k}}=0\Big\}$
and we equip ${\rm c_{0}}^{\mathcal{C},\tau}$ with the norm
$\|s\|_{0}:=\sup_{k\in\mathbb{Z}}\|s_{k}\|_{\tau\circ\alpha^{k}}.$
Since by [14] we have that $\mathcal{C}_{\tau\circ\alpha^{k}}$ is Banach algebra for all $k\in\mathbb{Z},$ it is not hard to check that ${\rm c}_{0}^{\mathcal{C},\tau}$ is a non-unital Banach algebra. We will from now on denote ${\rm c_{0}}^{\mathcal{C},\tau}$ by $\mathcal{A}.$ Moreover, we let $\mathcal{A}_{1}$ be the same as above in the previous case when we considered $\mathcal{A}=\ell_{2}(C_{0}(\Omega)).$ It is straightforward to check that the condition (E) is satisfied also in this case when ${\rm c_{0}}^{\mathcal{C},\tau}=\mathcal{A}$ (and $\mathcal{A}_{1}$ is as above in the previous case).
For $\epsilon_{1},\epsilon_{2}\in(0,1)$ with $\epsilon_{2}<\epsilon_{1},J\in\mathbb{N},$ and any collection of compact subsets $K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J+1})},\cdots,K_{\epsilon_{2}}^{(\tau\circ\alpha^{J})}$ of $\mathbb{R},$ we let $\tilde{p}_{(K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J+1})},\cdots,K_{\epsilon_{2}}^{(\tau\circ\alpha^{J})}),\epsilon_{1}}\in\mathcal{A}$ be given by
[…]
(where $\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}$ is the function from Lemma 3.8 with compact ${\rm supp}\,\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\subseteq\\ |\tau\circ\alpha^{j}|^{-1}(\left[0,\epsilon_{1}\right])$). By the proof of Lemma 3.9 (see [15]) one can deduce that the condition ($\mathcal{P}$) is satisfied in this case.
Let now $\alpha=\alpha_{1}=\cdots=\alpha_{N}.$ For each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\},$ let $\Phi_{n,\ell}$ be the same as above in the previous case when we considered $\mathcal{A}=\ell_{2}(C_{0}(\Omega)).$ It is not hard to check that the system $\{\Phi_{k,1}\}_{k\in\mathbb{N}},\dots,\{\Phi_{k,N}\}_{k\in\mathbb{N}}$ is again a system of isometric isomorphisms also in this case when ${\rm c_{0}}^{\mathcal{C},\tau}=\mathcal{A}.$ Moreover, it is disjoint apreiodic and satisfies the condition (R) with respect to
$\{\tilde{p}_{((K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J+1})},\cdots,K_{\epsilon_{2}}^{(\tau\circ\alpha^{J})}),\epsilon_{1})}\}.$
For each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\}$, let $\{w_{j}^{(1)}\}_{j\in\mathbb{Z}},\dots,\{w_{j}^{(N)}\}_{j\in\mathbb{Z}}\subseteq C_{b}(\mathbb{R}),b_{n,\ell}\in\mathcal{A}_{1}$ and $T_{n,\ell}:\mathcal{A}\to\mathcal{A}$ be as above in the previous case (when we considered $\mathcal{A}=\ell_{2}(C_{0}(\Omega))$).
Now, given $\epsilon_{1}\in(0,1),$ we choose some $\epsilon_{3}\in(\epsilon_{1},1).$ Since ${\rm supp}\,\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\subseteq\\ |\tau\circ\alpha^{j}|^{-1}(\left[0,\epsilon_{1}\right])$ and ${\rm supp}\,\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}$ is compact, by Lemma 3.8, we can find some function $v_{\bigl(supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\bigr),\epsilon_{3}}\in\mathcal{C}_{\tau\circ\alpha^{j}}$ such that $v_{\bigl(supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\bigr),\epsilon_{3}}(t)=1$ for all $t\in supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}},\text{ }supp\text{ }v_{\bigl(supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\bigr),\epsilon_{3}}$ is compact and $supp\text{ }v_{\bigl(supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})},\epsilon_{1}}\bigr),\epsilon_{3}}\subseteq|\tau\circ\alpha^{j}|^{-1}(\left[0,\epsilon_{3}\right]).$ Therefore, for each $n\in\mathbb{N},\ell\in\{1,\dots,N\},J\in\mathbb{N},\epsilon_{1},\epsilon_{2}\in(0,1)$ with $\epsilon_{2}<\epsilon_{1}$ and any collection of compact subsets
$K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J+1})},\cdots,K_{\epsilon_{2}}^{(\tau\circ\alpha^{J})},$
we let $b_{n,\ell,((K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J+1})},\cdots,K_{\epsilon_{2}}^{(\tau\circ\alpha^{J})}),\epsilon_{1})}^{-1}\in\mathcal{A}_{1}$ be given by
$(b_{n,\ell,((K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{2}}^{(\tau\circ\alpha^{-J+1})},\cdots,K_{\epsilon_{2}}^{(\tau\circ\alpha^{J})}),\epsilon_{1})}^{-1})_{j+r_{\ell}n}=\frac{v_{\bigl(supp\text{ }\mu_{K_{\epsilon_{2}}^{(\tau\circ\alpha^{j+r_{\ell}n})},\epsilon_{1}}\bigr),\epsilon_{3}}}{\prod_{k=1}^{r_{\ell}n}\big(w_{j+k}^{(\ell)}\circ\alpha^{n-k}\big)}$
for all $j\in\mathbb{Z}$ with $J-r_{\ell}n\leq j\leq J-r_{\ell}n$ (where we choose $\epsilon_{3}\in(\epsilon_{1},1)$) and $(b_{n,\ell,(\epsilon_{2},\epsilon_{1}J)}^{-1})_{j+r_{\ell}n}=0$ else.
By some calculations, it is not hard to deduce the next corollary as a corollary of Theorem 2.1.
Corollary 3.11.
Under the above notation and assumptions, the following statements are equivalent:
(1) The sequences of operators
$\{T_{n,1}\}_{n\in\mathbb{N}},\ \dots,\ \{T_{n,N}\}_{n\in\mathbb{N}}$
are $d\mathcal{F}$-semi-transitive on $\mathcal{A}$.
(2) For every $J\in\mathbb{N}$ and finite collection
$\{K_{\epsilon_{J}}^{(\tau\circ\alpha^{-J})},K_{\epsilon_{-J+1}}^{(\tau\circ\alpha^{-J+1})},\ldots,K_{\epsilon_{0}}^{\tau},\ldots,K_{\epsilon_{J}}^{(\tau\circ\alpha^{J})}\}$
of compact subsets of $\mathbb{R}$, there exists a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ such that
$\lim_{k\to\infty}\Bigg(\sup_{t\in K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})}}\frac{1}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j+i-r_{s}n_{k}}^{(s)}\circ\alpha^{r_{s}n_{k}-i}\big)(t)}\Bigg)\cdot\Bigg(\sup_{t\in K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})}}\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j+i}^{(\ell)}\circ\alpha^{-i}\big)(t)\Bigg)=0$
for all $j\in\mathbb{Z}$ with $J\leq j\leq J$ and every $l,s\in\{1,\dots,N\}$. Moreover, for each distinct $s,\ell\in\{1,\dots,N\}$ and all $j\in\mathbb{Z}$ with $J\leq j\leq J$, it holds that
$\lim_{k\to\infty}\Bigg(\sup_{t\in K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})}}\frac{\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j-r_{\ell}n_{k}-i}^{(\ell)}\circ\alpha^{-i}\circ\alpha^{r_{s}n_{k}}\big)(t)}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j-r_{s}n_{k}+i}^{(s)}\circ\alpha^{n_{k}-i}\big)(t)}\Bigg)=0.$