### v8

# Disjoint $\mathcal{F}$-semi-transitivity in Banach modules

Abstract.
Motivated by the concept of disjoint topological transitivity for supercyclicity given in [28], in this paper, we consider the concept of disjoint Furstenberg-semi-transitivity for operators that are a composition of an isometric isomorphism and a left multiplier on a normed algebra. Thus, we characterize disjoint $\mathcal{F}$-semi-transitive and disjoint supercyclic such operators on a large class of non-unital normed algebras. Generalized weighted composition operators on the normed algebra of operator-valued continuous functions vanishing at infinity on a locally compact, non-compact Hausdorff space are a special case of our theory. At the end of the paper, we characterize disjoint $\mathcal{F}$-semi-transitive adjoints of weighted composition operators acting on the weighted space of Radon measures.
Keywords:
Mathematics Subject Classification (2010) Primary MSC 47A16, Secondary MSC 54H20.
1. Introduction
Supercyclicity is a fundamental concept in the dynamics of operators and numerous papers have been written on the topics related to supercyclicity of operators. For example, Hilden and Wallen in [17] proved that any unilateral backward weighted shift is supercyclic. Afterwards, Salas in [33] characterized supercyclic bilateral weighted shift operators on $l^{p}(\mathbb{Z})$ in terms of a supercyclicity criterion. Supercyclicity of several kinds of operators has also been studied in for instance [12, 28, 30, 36]. Moreover, quite recently, a characterization of supercyclic and topologically semi-transitive weighted composition operators on solid Banach function spaces, in particular Morrey spaces, and Segal algebras have been given in [24]. Further, there is a close connection between supercyclicity and semi-Fredholm theory, see [1, 6, 13]. On the other hand, Furstenberg topological transitivity as a generalization of topological transitivity and hypercyclicity has been studied in several papers, see [2, 5, 15, 16, 27]. Motivated by these facts, in this paper we consider the concept of $\mathcal{F}$-semi-transitivity (Definition 2.1), as a generalization of the concept of topological transitivity for supercyclicity given in [28] and we characterize disjoint $\mathcal{F}$-semi-transitive operators on a large class of non-unital normed algebra (Theorem 3.1 and Corollary 3.2). The operators we study are in fact a composition of an isometric isomorphism and a left multiplier on a non-unital normed algebra. The reason for considering these operators is because many known classes of operators that have earlier been studied in connection with the dynamics, such as bilateral weighted shifts on $l^{2}(\mathbb{Z}),$ weighted composition operators on algebras of functions, generalized weighted shifts on Hilbert C*-modules and so on, turn out to be of this kind.
Next, the dynamics of cosine operator functions has been studied in several papers, see [7, 8, 9, 10, 11, 35, 20, 23, 25, 26, 38]. Although there are several published papers on the dynamics of cosine operator functions, supercyclicity of cosine operator functions has not been studied in any of the published papers, as far as we know. All these facts naturally lead us to study in this paper supercyclicity and $\mathcal{F}$-semi-transitivity (as a generalization of supercyclicity) of cosine operator functions generated by the above mentioned operators on normed non-unital algebras, i.e. the operators that are a composition of an isometric isomorphism and a left multiplier. In Proposition 3.4 and Corollary 3.5 we provide some sufficient conditions for the cosine operators functions generated by these operators to be $\mathcal{F}$-semi-transitive on the respective normed algebra.
Now, an attempt to axiomatically study hypercyclicity of operators on C*-algebras has been done in [20, Section 4]. This approach is limited to C*-algebras and does not cover many important and relevant cases in linear dynamics. The approach in Section 3 in this paper is more general because we consider a large class of non-unital normed algebras that are not necessarily C*-algebras. Our approach covers also the case of generalized bilateral weighted shift operators on standard Hilbert C*-modules defined in [19, 20].
In Section 4 we provide a new application in the context of generalized weighted composition operators on the space of operator valued continuous functions. More precisely, we denote by $\mathcal{C}$ the C*-algebra of compact operators on a separable Hilbert space $H,$ and in Corollary 4.2 we characterize disjoint $\mathcal{F}$-semi-transitive generalized weighted composition operators on the algebra $C_{0}(\Omega,\mathcal{C})$ equipped with the supremum-norm, i.e. the normed algebra of all $\mathcal{C}$-valued continuous functions vanishing at infinity on a locally compact, non-compact Hausdorff space $\Omega.$ The generalized weighted composition operators that we study in this context are in fact a composition of a weighted composition operator on $C_{0}(\Omega,\mathcal{C})$ and a right multiplication by a unitary operator on $H.$ Thus, these operators are at a same time a generalization of (ordinary) weighted composition operators on the classical space $C_{0}(\Omega),$ and elementary (wedge) operators on $\mathcal{C}$ considered in [21]. We give also a concrete example where the conditions of Corollary 4.2 are satisfied (Example 4.3).
In Section 5, we study the dynamics of adjoints of weighted composition operators. These adjoints act on the weighted space of Radon measures on a locally compact non-compact Hausdorff space. The motivation for considering the dynamics of operators on weighted spaces of Radon measures comes from the recent paper [22] where the dynamics of weighted composition operators on weighted solid Banach function spaces have been studied. In Theorem 5.5 we characterize the adjoints of weighted composition operators that have a dense set of periodic elements, obtaining in this way an improvement of [23, Proposition 10.3] by deducing equivalent conditions that are weaker than sufficient conditions given in [23, Proposition 10.3] for chaos of the adjoints of weighted composition operators. Next, in Theorem 5.6, we characterize disjoint $\mathcal{F}$-semi-transitive adjoints of weighted composition operators on weighted spaces of Radon measures.
2. Preliminaries
We need first the following definition.
Definition 2.1.
Let $\mathcal{S}$ be a set, $\mathcal{F}$ be a family of subsets of $\mathcal{S}$ and
$\{T_{t,1}\}_{t\in\mathcal{S}},\dots,\{T_{t,N}\}_{t\in\mathcal{S}}$
be families of bounded linear operators on a Banach space $X$. We say that $\{T_{t,1}\}_{t\in\mathcal{S}},\dots,\{T_{t,N}\}_{t\in\mathcal{S}}$ are disjoint $\mathcal{F}$-semi-transitive, or shortly $d\mathcal{F}$-semi-transitive, if for every collection of non-empty open subsets $\mathcal{O},V_{1},\dots,V_{N}$ of $X$, there exists some $F\in\mathcal{F}$ such that for all $t\in F$ there exists some $\lambda_{t}\in\mathbb{R}^{+}$ satisfying that
$\mathcal{O}\cap\lambda_{t}T_{t,1}^{-1}(V_{1})\cap\dots\cap\lambda_{t}T_{t,N}^{-1}(V_{N})\neq\varnothing.$
We recall also that the families $\{T_{t,1}\}_{t\in S},\dots,\{T_{t,N}\}_{t\in S}$ are said to be disjoint $\mathcal{F}$-transitive, or shortly $d\mathcal{F}$-transitive, if for every collection of non-empty open subsets $\mathcal{O},V_{1},\dots,V_{N}$ of $X$, there exists some $F\in\mathcal{F}$ such that for all $t\in F$ we have
$\mathcal{O}\cap T_{t,1}^{-1}(V_{1})\cap\dots\cap T_{t,N}^{-1}(V_{N})\neq\varnothing.$
Definition 2.2.
Let $\mathcal{X}$ be a Banach space, and $(T_{n})_{n\in\mathbb{N}_{0}}$ be a sequence of operators in $B(\mathcal{X})$. A vector $x\in\mathcal{X}$ is called a periodic element of $(T_{n})_{n\in\mathbb{N}_{0}}$ if there exists a constant $N\in\mathbb{N}$ such that for each $k\in\mathbb{N}$, $T_{kN}x=x$. The set of all periodic elements of $(T_{n})_{n\in\mathbb{N}_{0}}$ is denoted by ${\mathcal{P}}((T_{n})_{n\in\mathbb{N}_{0}})$. The sequence $(T_{n})_{n\in\mathbb{N}_{0}}$ is called chaotic if $(T_{n})_{n\in\mathbb{N}_{0}}$ is topologically transitive and ${\mathcal{P}}((T_{n})_{n\in\mathbb{N}_{0}})$ is dense in $\mathcal{X}$}. An operator $T\in B(\mathcal{X})$ is called chaotic if the sequence $\{T^{n}\}_{n\in\mathbb{N}_{0}}$ is chaotic.
Let now $\mathcal{A}$ be a non-unital normed algebra such that $\mathcal{A}$ is a left ideal in a unital normed algebra $\mathcal{A}_{1}$. We do not assume that the norm on $\mathcal{A}$ extends to $\mathcal{A}_{1}$. We will say that $\mathcal{A}$ satisfies the condition (E) with respect to $\mathcal{A}_{1}$ if
$\|ba\|\leq\|b\|_{1}\|a\|$
for all $a\in\mathcal{A}$ and $b\in\mathcal{A}_{1},$ where  $\|\text{ }\|_{1}$ denotes the norm on $\mathcal{A}_{1}.$
Let $\{p_{\alpha}\}_{\alpha}$ be a set in $\mathcal{A}$ satisfying that given any open subset $O$ of $\mathcal{A}$ and $x\in O$ there exists some $p_{\alpha_{0}}\in\{p_{\alpha}\}_{\alpha}$ such that $p_{\alpha_{0}}^{3}x\in O$ as well. This condition will be called the condition $(\mathcal{P}.)$
For a set $\mathcal{S},$ let $\mathcal{F}$ be a family of subsets of $\mathcal{S}$, and let
$\{\Phi_{t,1}\}_{t\in\mathcal{S}},\ \{\Phi_{t,2}\}_{t\in\mathcal{S}},\ \ldots,\ \{\Phi_{t,N}\}_{t\in\mathcal{S}}$
be families of isometric algebra isomorphism of $\mathcal{A}_{1}$ such that
$\Phi_{t,j}(\mathcal{A})=\mathcal{A}\qquad(t\in\mathcal{S},\ 1\leq j\leq N).$
and such that $\Phi_{t,j|\mathcal{A}}$ is an isometry for all $j\in\{1,\dots,N\}.$ We will assume that the system
$\{\Phi_{t,1}\}_{t\in\mathcal{S}},\ \{\Phi_{t,2}\}_{t\in\mathcal{S}},\ \ldots,\ \{\Phi_{t,N}\}_{t\in\mathcal{S}}$
is disjoint aperiodic with respect to $\{p_{\alpha}\}_{\alpha}$, that is for each fixed $p_{\alpha}$ and every $H\in\mathcal{F}$ there exist some $F\subseteq H$ with $F\in\mathcal{F}$ such that
$p_{\alpha}\,\Phi_{t,\ell}(p_{\alpha})=0,\quad p_{\alpha}\,\Phi_{t,\ell}^{-1}(p_{\alpha})=0,\quad\text{and}\quad p_{\alpha}\,\Phi_{t,\ell}\!\big(\Phi_{t,r}^{-1}(p_{\alpha})\big)=0\quad\text{for all }t\in F$
$\text{ and all }r,\ell\in\{1,\dots,N\}\text{ with }r\neq\ell.$ Further, we will assume that for each $\alpha$ and all $t\in\mathcal{S}$, $r,\ell\in\{1,\dots,N\}$, and $a\in\mathcal{A}$, it holds that
$\|a\,p_{\alpha}\|\leq\|a\|,\quad\|a\,\Phi_{t,\ell}(p_{\alpha})\|\leq\|a\|,$
$\|a\,\Phi_{t,\ell}^{-1}(p_{\alpha})\|\leq\|a\|,\quad\quad\|a\,\Phi_{t,\ell}\big(\Phi_{t,r}^{-1}(p_{\alpha})\big)\|\leq\|a\|.$
This condition will be called the condition (R) throughout this section.
We now set up families $\{b_{t,1}\}_{t\in\mathcal{S}},\cdots,\{b_{t,N}\}_{t\in\mathcal{S}}$ of elements in $\mathcal{A}_{1}$, and for each $t\in\mathcal{S}$ and $\ell\in\{1,\dots,N\}$, we define the operator
$T_{t,\ell}:\mathcal{A}\to\mathcal{A}\quad\text{by}\quad T_{t,\ell}(a)=b_{t,\ell}\,\Phi_{t,\ell}(a).$
Due to the condition $(E)$, it is not hard to check that $T_{t,\ell}$ is a bounded linear operator on $\mathcal{A}$. Further, we will assume that for each $\alpha$ there exists so called $\alpha$-inverse” for $b_{t,\ell}$ denoted by $b_{t,\ell,\alpha}^{-1},$ that is for each $\alpha$ there exists an element $b_{t,\ell,\alpha}^{-1}\in\mathcal{A}_{1}$ satisfying that $b_{t,\ell}b_{t,\ell,\alpha}^{-1}ap_{\alpha}=b_{t,\ell,\alpha}^{-1}b_{t,\ell}ap_{\alpha}=ap_{\alpha}$ for all $a\in\mathcal{A}.$
Also, we will assume that for each $t\in\mathcal{S}$ and $\ell\in\{1,\dots,N\}$ it holds that $b_{t,\ell}a=0$ if and only if $a=0.$ Such condition will be called the condition (C) throughout the manuscript.
3. Main results
We have the following theorem.
Theorem 3.1.
Under the above notation and assumptions, the following statements are equivalent.
  1. (1)
    The families $\{T_{t,1}\}_{t\in\mathcal{S}},\cdots,\{T_{t,N}\}_{t\in\mathcal{S}}$ are $d\mathcal{F}$-semi-transitive.
  2. (2)
    For each fixed $p_{\alpha}$ and $\varepsilon>0$ there exists some $F\in\mathcal{F}$ and families $\{d_{t}\}_{t\in F}$, $\{g_{t,1}\}_{t\in F}\dots,\{g_{t,N}\}_{t\in F}$ in $\mathcal{A}$ such that $\|d_{t}-p_{\alpha}^{2}\|<\varepsilon,\quad\|g_{t,l}-p_{\alpha}^{2}\|<\varepsilon,$ and
    $\|b_{t,r,\alpha}^{-1}g_{t,r}\|\|\Phi_{t,l}^{-1}(b_{t,l})d_{t}\|<\varepsilon^{2}\quad\text{for all }t\in F\text{ and }r,l\in\{1,\ldots,N\}.$
    Moreover, for each distinct $r,l\in\{1,\ldots,N\}$ and $t\in F$ it holds that
    $\|\Phi_{t,r}^{-1}(b_{t,r})\Phi_{t,l}^{-1}(b_{t,l,\alpha}^{-1}g_{t,l})\|<\varepsilon.$
Proof.
We prove first (1) $\Rightarrow$ (2). Since the families $\{T_{\Phi_{t,1},b_{t,1}}\}_{t\in S},\dots,\{T_{\Phi_{t,N},b_{t,N}}\}_{t\in S}$ are $d\mathcal{F}$-semi-transitive, given $\varepsilon>0$ and some fixed $p_{\alpha}$ we can find $a_{t}\in\mathcal{A}$ and $\lambda_{t}\in\mathbb{R}^{+}$ such that $\|a_{t}-p_{\alpha}\|<\varepsilon$ and $\|\lambda_{t}T_{\Phi_{t,l}},b_{t,l}(a_{t})-p_{\alpha}\|<\varepsilon\text{ for all }t\in H\text{ and some }H\in\mathcal{F}.$ This gives $\bigl\|\,(a_{t}-p_{\alpha})\,p_{\alpha}\,\bigr\|\;\leq\;\bigl\|a_{t}-p_{\alpha}\bigr\|\;<\;\varepsilon.$
Moreover, since $\bigl\{\ {\Phi_{t,l}}\bigr\}_{t\in S,\;1\leq l\leq N}$ is a disjoint aperiodic system, we can find some $F\in\mathcal{F}$ with $F\subseteq H$ such that $p_{\alpha}\!\Phi_{t,l}^{-1}(p_{\alpha})=0,\;p_{\alpha}\!\Phi_{t,l}(p_{\alpha})=0,\;$ and $\Phi_{t,r}(\Phi_{t,l}^{-1}(p_{\alpha}))=0,\;\text{for all }\;t\in F\;\text{ and }r,l\in\{1,\dots N\}\text{ with }r\neq l.$ Hence, we get
$\qquad\frac{1}{|\lambda_{t}|}\,\left\|\Phi_{t,r}^{-1}\!\bigl(b_{t,r,\alpha}^{-1}\bigr)\,\Phi_{t,r}^{-1}\!\bigl(b_{t,r}\bigr)\,a_{t}\,\Phi_{t,r}^{-1}(p_{\alpha})\,\lambda_{t}\right\|=\left\|\!b_{t,r,\alpha}^{-1}\,\!\bigl(b_{t,r}\,\Phi_{t,r}(a_{t})\,(p_{\alpha})\,\bigr)\lambda_{t}\right\|$
$\left\|a_{t}\,\Phi_{t,r}^{-1}(p_{\alpha})\right\|=\left\|(a_{t}-p_{\alpha})\,\Phi_{t,r}^{-1}(p_{\alpha})\right\|\leq\left\|a_{t}-p_{\alpha}\right\|<\varepsilon.$
and
$\left\|\lambda_{t}\,b_{t,r}\,\Phi_{t,r}(a_{t})\,p_{\alpha}\;-\;p_{\alpha}^{2}\right\|\leq\left\|\lambda_{t}\,b_{t,r}\,\Phi_{t,r}(a_{t})-p_{\alpha}\right\|=\left\|\lambda_{t}\,T_{t,r}(a_{t})-p_{\alpha}\right\|<\epsilon,\qquad$
for all $t\in F$ and $r\in\{1,\dots,N\}$.
For each $t\in F$ and $r\in\{1,\dots,N\},$ we put $g_{t,r}=\lambda_{t}\,T_{t,r}(a_{t})p_{\alpha}.$ Then
$\|g_{t,r}-p_{\alpha}^{2}\|\;<\;\varepsilon$
and
$\frac{1}{|\lambda_{t}|}\,\bigl\|\Phi_{t,r}^{-1}\!\bigl(b_{t,r,\alpha}^{-1}\bigr)\,\Phi_{t,r}^{-1}(g_{t,r})\bigr)\bigr\|\;=\;\frac{1}{|\lambda_{t}|}\,\|b_{t,r,\alpha}^{-1}g_{t,r}\|\;<\;\varepsilon.$
Similarly, for each $t\in F$ and $l\in\{1,\dots,N\},$ we have
$\|\lambda_{t}\,b_{t,l}\,\Phi_{t,l}(a_{t})\,\Phi_{t,l}(p_{\alpha})\|\;=\;\|(\lambda_{t}\,b_{t,l}\,\Phi_{t,l}(a_{t})-p_{\alpha})\,\Phi_{t,l}(p_{\alpha})\|\;\leq\;\|\lambda_{t}\,T_{t,l}(a_{t})-p_{\alpha}\|\;<\;\varepsilon,$
and, as shown above,
$\|a_{t}p_{\alpha}-p_{\alpha}^{2}\|\;<\;\varepsilon$
For each $t\in F$ and $l\in\{1,\dots,N\},$ set $d_{t}=a_{t}p_{\alpha}.$ Then
$\|d_{t}-p_{\alpha}^{2}\|\;<\;\varepsilon$
and
$|\lambda_{t}|\,\bigl\|\Phi_{t,l}^{-1}(b_{t,l})\,d_{t}\bigr\|\;=\;\|\lambda_{t}\,b_{t,l}\,\Phi_{t,l}(a_{t}p_{\alpha})\|\;<\;\varepsilon.$
Hence, for all $t\in F$ and $r,l\in\{1,\dots,N\},$ we have
$\|b_{t,r,\alpha}^{-1}(g_{t,r})\|\|\Phi_{t,l}^{-1}(b_{t,l})d_{t}\|<\varepsilon^{2}.$
Further, for all $t\in F$ and each distinct $r,l\in\{1,\dots,N\}$ we get
$\Bigl\|\,\Phi_{t,r}^{-1}\!\bigl(b_{t,r}\bigr)\,\Phi_{t,l}^{-1}\!\bigl(b_{t,l,\alpha}^{-1}\bigr)\Phi_{t,l}^{-1}\!\bigl(g_{t,l}\bigr)\Bigr\|\;=\;\Bigl\|\,\!b_{t,r}\,\Phi_{t,r}\Phi_{t,l}^{-1}\!\bigl(b_{t,l,\alpha}^{-1}g_{t,l}\bigr)\Bigr\|$
$\bigl\|\,\lambda_{t}\,b_{t,r}\,\Phi_{t,r}(a_{t})\,\Phi_{t,r}\!\bigl(\Phi_{t,l}^{-1}(p_{\alpha})\bigr)\bigr\|$
$\bigl\|\,(\lambda_{t}\,T_{t,r}(a_{t})-p_{\alpha})\,\Phi_{t,r}\!\bigl(\Phi_{t,l}^{-1}(p_{\alpha})\bigr)\bigr\|\;<\;\varepsilon,$
where we have used that $p_{\alpha}\,\Phi_{t,r}\!\bigl(\Phi_{t,l}^{-1}(p_{\alpha})\bigr)=0\quad\text{for all }t\in F\text{ and each distinct }r,l\in\{1,\dots,N\}.$
Now we prove the implication $(2)\Rightarrow(1)$. Let $O,V_{1},\dots,V_{N}$ be non-empty open subsets of $\mathcal{A}$. Then we can find some non-zero $x\in O$ and non-zero $y_{l}\in V_{l}$ for each $l\in\{1,\dots,N\}$. There exists some $\alpha$ such that $p_{\alpha}^{3}x\in O$ and $p_{\alpha}^{3}y_{l}\in V_{l}\setminus\{0\}$ for each $l=1,\dots,N$. Clearly, this gives that $p_{\alpha}x\neq 0$ and $p_{\alpha}y_{l}\neq 0$ for each $l=1,\dots,N$. Also, since $O\setminus\{0\},\,V_{1}\setminus\{0\},\dots,V_{N}\setminus\{0\}$ are open and non-empty, there is some $\delta>0$ such that the $\delta$-neighbourhood of $p_{\alpha}^{3}x$ is contained in $O\setminus\{0\}$, and the $\delta$-neighbourhood of $p_{\alpha}^{3}y_{l}$ is contained in $V_{l}\setminus\{0\}$ for each $l=1,\dots,N$.
Let $C_{y}=\max_{1\leq l\leq N}\bigl\{\|y_{l}\|,\bigr\}$ and set
$\varepsilon:=\min\!\left\{\frac{\delta}{4\|x\|\|p_{\alpha}\|},\;\frac{\delta}{4N\sqrt{\|x\|C_{y}}\|p_{\alpha}\|}\frac{\delta}{4NC_{y}\|p_{\alpha}\|},\frac{\|p_{\alpha}^{3}x\|}{2\|p_{\alpha}x\|},\frac{\|p_{\alpha}^{3}y_{1}\|}{2\|p_{\alpha}y_{1}\|},\cdots,\frac{\|p_{\alpha}^{3}y_{N}\|}{2\|p_{\alpha}y_{N}\|}\right\}.$
Choose $F\in\mathcal{F}$ and the families $\{g_{t,1}\}_{t\in F},\dots,\{g_{t,N}\}_{t\in F},\{d_{t}\}_{t\in F}$ satisfying the assumptions in $(2)$ with respect to $\epsilon$ and $p_{\alpha}$.For each $\alpha$ and every $t\in F,l\in\{1,\dots,N\}$ we define the operator $S_{t,l,\alpha}$ on $\mathcal{A}$ by $S_{t,l,\alpha}(a)=\Phi_{t,l}^{-1}(b_{t,l,\alpha}^{-1})\Phi_{t,l}^{-1}(a)$ for all $a\in\mathcal{A}.$ We observe now that for each $t\in F$ and $l\in\{1,\dots,N\}$ it holds that
$\parallel T_{t,l}\!\bigl(d_{t}p_{\alpha}x\bigr)\parallel=\parallel\,b_{t,l}\,\Phi_{t,l}\bigl(d_{t}\bigr)\,\Phi_{t,l}(p_{\alpha}x)\,\parallel\leq\parallel\,b_{t,l}\,\Phi_{t,l}\bigl(d_{t}\bigr)\,\parallel\,\parallel\,\Phi_{t,l}(p_{\alpha}x)\,\parallel$
$\parallel\,\Phi_{t,l}^{-1}\,(b_{t,l})\,d_{t}\,\parallel\;\parallel\,x\,\parallel\|p_{\alpha}\|$
and
$\parallel S_{t,l,\alpha}\!\bigl(g_{t,l}p_{\alpha}\,y_{\ell}\bigr)\parallel=\parallel\,\Phi_{t,l}^{-1}\,(b_{t,l,\alpha}^{-1})\,\Phi_{t,l}^{-1}\!\bigl(g_{t,l}\bigr)\,\Phi_{t,l}^{-1}(p_{\alpha}\,y_{\ell})\,\parallel$
$\leq\parallel\,\Phi_{t,l}^{-1}\,(b_{t,l,\alpha}^{-1}\,g_{t,l})\,\parallel\,\parallel\,\Phi_{t,l}^{-1}(p_{\alpha}\,y_{\ell})\,\parallel=\parallel\,b_{t,l,\alpha}^{-1}\,g_{t,l}\,\parallel\;\parallel\,y_{\ell}\,\parallel\|p_{\alpha}\|.$
Moreover, for all $t\in F$ and each distinct $r,l\in\{1,\dots,N\}$ we have
$\parallel T_{t,r}\!\bigl(S_{t,l,\alpha}(g_{t,l}p_{\alpha}\,y_{\ell})\bigr)\parallel=\parallel\,b_{t,r}\,\Phi_{t,r}\,(\Phi_{t,l}^{-1}\,(b_{t,l,\alpha}^{-1})\bigr)\,\Phi_{t,r}\,\bigl(\Phi_{t,l}^{-1}\,(g_{t,l}p_{\alpha}\,y_{\ell}\bigr)\bigr)\,\parallel$
$\parallel\Phi_{t,r}^{-1}\,(b_{t,r}\,)\Phi_{t,l}^{-1}\,(b_{t,l,\alpha}^{-1})\,\Phi_{t,l}^{-1}\,(g_{t,l})\,\Phi_{t,l}^{-1}\,(p_{\alpha}\,y_{\ell})\,\parallel$
$\leq\parallel\Phi_{t,r}^{-1}\,(b_{t,r}\,)\Phi_{t,l}^{-1}\,(b_{t,l,\alpha}^{-1}g_{t,l})\,\,\parallel\,\parallel y_{\ell}\parallel\|p_{\alpha}\|.$
Finally, since $b_{t,l,\alpha}^{-1}$ is the $\alpha$-inverse of $b_{t,\ell},$ it follows that
$T_{t,l}\!\bigl(S_{t,l,\alpha}(g_{t,l}p_{\alpha}\,y_{\ell}))=g_{t,l}p_{\alpha}\,y_{\ell}$
for all $t\in F$ and each $\ell\in\{1,\dots,N\}.$
Since $\|d_{t}p_{\alpha}x-p_{\alpha}^{3}\,x\|<\|\epsilon\|\|p_{\alpha}x\|<\|p_{\alpha}^{3}\,x\|$, it follows that $d_{t}p_{\alpha}x\neq 0.$ Similarly, $g_{t,l}p_{\alpha}y_{\ell}\neq 0$ for all $t\in F$ and $\ell\in\{1,\dots,N\}$. Hence, since by the assumption we have that for each $t\in\mathcal{S}$ and $\ell\in\{1,\dots,N\}$ it holds that $b_{t,\ell}a=0$ if and only if $a=0,$ it follows that $T_{t,\ell}(d_{t}p_{\alpha}x)\neq 0$ for each $t\in F$ and $\ell\in\{1,\dots,N\}.$ Similarly, since $b_{t,l,\alpha}^{-1}$ is the $\alpha$-inverse of $b_{t,\ell},$ it folllows that $S_{t,\ell,\alpha}(g_{t,\ell}p_{\alpha}y_{\ell})\neq 0$ for all $t\in F$ and $\ell\in{1,\dots,N}.$
For each $t\in F$, we put
$x_{t}:=d_{t}p_{\alpha}x+\frac{\sqrt{\sum_{l=1}^{N}\,\parallel T_{t,l}(d_{t}p_{\alpha}x)\parallel}}{\sqrt{\sum_{l=1}^{N}\,\parallel S_{t,l,\alpha}(g_{t,l}p_{\alpha}y_{\ell})\parallel}\qquad}\,{\sum_{l=1}^{N}\,S_{t,r,\alpha}(g_{t,r}p_{\alpha}y_{r})}.$
By triangle inequality and choice of $\varepsilon$ one can deduce that for each $t\in F$, we have
$x_{t}\in O,\ \text{and}\ \frac{\sqrt{\sum_{\ell=1}^{N}\|S_{t,\ell,\alpha}(g_{t,\ell}p_{\alpha}y_{\ell})\|}}{\sqrt{\sum_{\ell=1}^{N}\|T_{t,\ell}(d_{t}p_{\alpha}x)\|}}T_{t,r}(x_{t})\in V_{r}.$
for all $r\in\{1,\dots,N\}.$
The next corollary can easily be deduced from Theorem 3.1.
Corollary 3.2.
Let now $\mathcal{F}$ be the family of all infinite subsets of $\mathbb{N}.$ If there exist dense subsets $\mathcal{A}_{0},\mathcal{B}_{1},\cdots,\mathcal{B}_{N}$ of $\mathcal{A}$ and a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ such that for each $s,\ell\in\{1,\dots,N\},$ every fixed $\alpha$ and all $x\in\mathcal{A}_{0},y_{s}\in\mathcal{B}_{s}$ we have
$\lim_{k\to\infty}\|b_{n_{k},s,\alpha}^{-1}y_{s}\|\|\Phi_{n_{k},l}^{-1}(b_{n_{k},l})x\|=0,$
and, in addition, for each distinct $s,\ell\in\{1,\ldots,N\}$ it holds that
$\lim_{k\to\infty}\|\Phi_{n_{k},\ell}^{-1}(b_{n_{k},\ell})\Phi_{n_{k},s}^{-1}(b_{n_{k},s,\alpha}^{-1}y_{s})\|=0,$
then the families $\{T_{n,1}\}_{n\in\mathbb{N}},\cdots,\{T_{n,N}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
Remark 3.3.
We notice that the assumption that $\{\Phi_{t,1}\}_{t\in\mathcal{S}},\cdots,\{\Phi_{t,N}\}_{t\in\mathcal{S}}$ is a disjoint aperiodic system in Theorem 3.1 is only needed for the proof of the implication $(1)\Rightarrow(2)$. Also, we notice that we have used the condition (R) only in the proof of the implication $(1)\Rightarrow(2)$ to obtain the desired inequalities. Hence, the assumption that $\{\Phi_{t,1}\}_{t\in\mathcal{S}},\cdots,\{\Phi_{t,N}\}_{t\in\mathcal{S}}$ is disjoint aperiodic and the condition (R) are not needed in Corollary 3.2.
Now we set $N=1$ and we will therefore write $b_{t}$, $\Phi_{t}$ and $T_{t}$ instead of $b_{t,1}$, $\Phi_{t,1}$ and $T_{t,1},$ respectively. We will now assume that $b_{t}$ is invertible for each $t\in\mathcal{S}$ and we will consider the map $S_{t}:\mathcal{A}\rightarrow\mathcal{A}$ given by $S_{t}(a)=\Phi_{t}^{-1}(b_{t}^{-1})\,\Phi_{t}^{-1}(a)$ for all $a\in\mathcal{A}$. Obviously, $S_{t}=T_{t}^{-1}$ for each $t\in\mathcal{S}.$
We set $C_{t}:=\frac{1}{2}(T_{t}+S_{t})$ for all $t\in\mathcal{S}$.
Proposition 3.4.
Under the above notation and assumptions, we have $(2)\Rightarrow(1)$.
(1) The family $\{C_{t}\}_{t\in\mathcal{S}}$ is $\mathcal{F}$–semi–transitive.
(2) For each fixed $p_{\alpha}$ and $\varepsilon>0$ there exists some $F\in\mathcal{F}$ and families $\{d_{t}\}_{t\in F}$, $\{g_{t}\}_{t\in F}$, $\{e_{t}\}_{t\in F}$, $\{f_{t}\}_{t\in F}$ with $e_{t}+f_{t}=g_{t}$ for all $t\in F$ such that
$\|d_{t}-p_{\alpha}^{2}\|<\varepsilon,\qquad\|g_{t}-p_{\alpha}^{2}\|<\varepsilon,$
$\|\Phi_{t}^{-2}(b_{t})\Phi_{t}^{-1}(b_{t})e_{t}\|<\varepsilon,$
$\|\Phi_{t}(b_{t}^{-1})\,b_{t}^{-1}\,f_{t}\|<\varepsilon,$
$\|\Phi_{t}^{-1}(b_{t})\,d_{t}\|\,\|\Phi_{t}^{-1}(b_{t})\,e_{t}\|<\varepsilon^{2},$
$\|\Phi_{t}^{-1}(b_{t})\,d_{t}\|\,\|b_{t}^{-1}f_{t}\|<\varepsilon^{2},$
$\|b_{t}^{-1}\,d_{t}\|\,\|\Phi_{t}^{-1}(b_{t})\,e_{t}\|<\varepsilon^{2},$
$\|b_{t}^{-1}\,d_{t}\|\,\|b_{t}^{-1}\,f_{t}\|<\varepsilon^{2},\text{ for all }t\in F.$
Proof.
Given non–empty open subsets $\mathcal{O}$ and $\mathcal{V}$ of $\mathcal{A}$, as in the proof of Theorem 3.1 part $(2)\Rightarrow(1)$ we can find some $\delta>0$ and some $p_{\alpha}^{3}x\in\mathcal{O}\setminus\{0\}$, $p_{\alpha}^{3}y\in\mathcal{V}\setminus\{0\}$ such that $\delta$–neighbourhoods of $p_{\alpha}^{3}x$ and $p_{\alpha}^{3}y$ are contained in $\mathcal{O}\setminus\{0\}$ and $V\setminus\{0\}$, respectively.
Set
$\varepsilon:=\min\left\{\frac{\delta}{4\|p_{\alpha}\|(\|x\|+\|y\|)},\;\frac{\delta}{16\|p_{\alpha}\|\sqrt{(\|x\|\|y\|)}},\;\frac{\|p_{\alpha}^{3}x\|}{2\|p_{\alpha}x\|},\;\frac{\|p_{\alpha}^{3}y\|}{2\|p_{\alpha}y\|}\right\}.$
Choose $F\in\mathcal{F}$ and families $\{d_{t}\}_{t\in F}$, $\{g_{t}\}_{t\in F}$ , $\{e_{t}\}_{t\in F}$, $\{f_{t}\}_{t\in F}$ satisfying the assumptions in (2) with respect to $p_{\alpha}$ and $\varepsilon$. Since $e_{t}+f_{t}=g_{t}$ and $g_{t}\,p_{\alpha}y\neq 0$ (as shown in the proof of Theorem 3.1 part $(2)\Rightarrow(1)$), it follows that either $e_{t}\,p_{\alpha}y\neq 0$ or $f_{t}\,p_{\alpha}y\neq 0$. Then, by the same arguments as in the proof of Theorem 3.1 part $(2)\Rightarrow(1)$ we can show that either $T_{t}(e_{t},p_{\alpha}y)\neq 0$ or $S_{t}(f_{t},p_{\alpha}y)\neq 0$.
Further, we notice that for each $t\in F$ it holds that
$\|T_{t}^{2}(e_{t},p_{\alpha}y)\|=\|b_{t}\,\Phi_{t}\,(b_{t})\,\Phi_{t}^{2}\,(e_{t}\,p_{\alpha}y)\|$
$\leq\|\Phi_{t}^{-2}\,(b_{t})\,\Phi_{t}^{-1}\,(b_{t})\,e_{t}\|\,\|p_{\alpha}\|\,\|y\|$
and,
$\|S_{t}^{2}(f_{t},p_{\alpha}y)\|=\|\Phi_{t}^{-1}(b_{t}^{-1})\cdot\Phi_{t}^{-2}(b_{t}^{-1}f_{t}\,p_{\alpha}y)\|\leq\|\Phi_{t}(b_{t}^{-1})b_{t}^{-1}f_{t}\|\,\|p_{\alpha}\|\,\|\|y\|.$
For each $t\in F,$ put
$v_{t}:=d_{t}p_{\alpha}x+\frac{2\sqrt{\|T_{t}(d_{t}p_{\alpha}x)\|+\|S_{t}(d_{t}p_{\alpha}x)\|}}{\sqrt{\|T_{t}(e_{t}p_{\alpha}y)\|+\|S_{t}(f_{t}p_{\alpha}y)\|}}(T_{t}(e_{t}+p_{\alpha}x)+S_{t}(e_{t}p_{\alpha}x)).$
By the above inequalities and inequalities in the proof of Theorem 3.1 part $(2)\Rightarrow(1)$, the choice of $\varepsilon$ and the triangle inequality, we can deduce that $v_{t}\in\mathcal{O}$ and
$\frac{\sqrt{\|T_{t}(e_{t}p_{\alpha}y)\|+\|S_{t}(f_{t}p_{\alpha}y)\|}}{\sqrt{\|T_{t}(d_{t}p_{\alpha}x)\|+\|S_{t}(d_{t}p_{\alpha}x)\|}}C_{t}(v_{t})\in\mathcal{V}\text{ for all }t\in F.$
By putting either $e_{t}=0$ (so that $f_{t}=g_{t}$), or by putting $f_{t}=0$ (so that $e_{t}=g_{t}$) in Proposition 3.4, then we can get the following corollary.
Corollary 3.5.
Let now $\mathcal{F}$ be the family of all infinite subsets of $\mathbb{N}$. If there exist dense subsets $\mathcal{B}_{1},\mathcal{B}_{2}$ of $\mathcal{A}$ and a strictly increasing sequence $\{n_{k}\}\subseteq\mathbb{N}$ such that for every fixed $p_{\alpha}$ it holds that either and
$\lim_{k\to\infty}\|\Phi_{n_{k}}^{-2}(b_{n_{k}})\,\Phi_{n_{k}}^{-1}(b_{n_{k}})\,y\|=\lim_{k\to\infty}\|b^{-1}_{n_{k}}x\|\,\|\Phi_{n_{k}}^{-1}(b_{n_{k}})y\|=$
$\lim_{k\to\infty}\|\Phi_{n_{k}}^{-1}(b_{n_{k}})x\|\,\|\Phi_{n_{k}}^{-1}(b_{n_{k}})y\|=0$
or
$\lim_{k\to\infty}\|\Phi_{n_{k}}(b_{n_{k}}^{-1})b_{n_{k}}^{-1}y\|=\lim_{k\to\infty}\|b_{n_{k}}^{-1}x)\|\cdot\|b_{n_{k}}^{-1}y\|=\lim_{k\to\infty}\|\Phi_{n_{k}}^{-1}(b_{n_{k}})\|\|b_{n_{k}}^{-1}y\|=0$
for all $x\in\mathcal{B}_{1}$ and $y\in\mathcal{B}_{2}$, then the family $\{C_{n}\}_{n\in\mathbb{N}}$ is $\mathcal{F}$–semi-transitive.
Remark 3.6.
The above results and the proofs can be extended and generalized to the case when $\mathcal{A}$ is a Banach bimodule over a unital normed algebra $\mathcal{A}_{1}.$ The condition (E) remains the same in this case, however, the condition $(\mathcal{P})$ should be then reformulated as follows:
Let $\{p_{\alpha}\}_{\alpha}$ be a set in $\mathcal{A}\cap\mathcal{A}_{1}$ satisfying that given any open subset $O$ of $\mathcal{A}$ there exists some $x\in\mathcal{A}\cup\mathcal{A}_{1}$ and some $p_{\alpha_{0}}\in\{p_{\alpha}\}_{\alpha}$ such that $p_{\alpha_{0}}x\in\mathcal{A}_{1}$ and $p_{\alpha_{0}}^{3}x\in O.$
Also, in this case, we let $\{\Phi_{t,1}\}_{t\in\mathcal{S}},\ \{\Phi_{t,2}\}_{t\in\mathcal{S}},\ \ldots,\ \{\Phi_{t,N}\}_{t\in\mathcal{S}}$ be families of isometric algebra isomorphisms of $\mathcal{A}_{1}$ and
$\{\Psi_{t,1}\}_{t\in\mathcal{S}},\ \{\Psi_{t,2}\}_{t\in\mathcal{S}},\ \ldots,\ \{\Psi_{t,N}\}_{t\in\mathcal{S}}$
be families of linear isometric isomorphisms of $\mathcal{A}$ such that for all $x,y\in\mathcal{A}_{1}$ and $a\in\mathcal{A}$ it holds that $\Psi_{t,j}(xay)=\Phi_{t,j}(x)\Psi_{t,j}(a)\Phi_{t,j}(y)$ for all $t\in\mathcal{S}$ and $j\in\{1,\dots,N\}.$ We express aperiodicity still in terms of the families
$\{\Phi_{t,1}\}_{t\in\mathcal{S}},\ \{\Phi_{t,2}\}_{t\in\mathcal{S}},\ \ldots,\ \{\Phi_{t,N}\}_{t\in\mathcal{S}},$ moreover, the conditions (R) and (C) remain the same in this case. Further, we define the operator
$T_{t,\ell}:\mathcal{A}\to\mathcal{A}\quad\text{by}\quad T_{t,\ell}(a)=b_{t,\ell}\,\Psi_{t,\ell}(a)$
for all $a\in\mathcal{A}.$ Under this set up and these assumptions, it is not hard to extend and generalize the proof of Theorem 3.1 in this setting.
4. Applications
From now on and in the rest of this section we will assume that $\mathcal{F}$ is the family of all infinite subsets of $\mathbb{N}.$ Further, in this section, for a separable Hilbert space $H$ we let $B(H)$ be the space of all bounded linear operators on $H$, and $\mathcal{C}$ be the $C^{*}$-algebra of compact operators on $H$. For an orthonormal basis $\{e_{j}\}_{j\in\mathbb{Z}}$ for $H$, we let for each $m\in\mathbb{N}$, $P_{m}$ be the orthogonal projection onto $\text{Span}\{e_{-m},\dots,e_{m}\}$.
Moreover, we will denote by $\Omega$ be a locally compact, non-compact Hausdorff, space. As usual, $C_{b}(\Omega)$ will denote the space of all bounded continuous functions on $\Omega$ and $C_{0}(\Omega)$ will denote the space of all continuous functions on $\Omega$ vanishing at infinity. Both these spaces will be equipped with supremum norm. In addition, we let $C_{c}(\Omega)$ be the space of all continuous compactly supported functions on $\Omega.$ For each pair of compact subsets $K_{1}$ and $K_{2}$ of $\Omega$ with $K_{1}\subseteq K_{2}$ , we let $u_{(K_{1},K_{2})}\in C_{0}(\Omega)$ be such that $u_{{(K_{1},K_{2})}|_{K_{1}}}=1$, $\operatorname{supp}u_{(K_{1},K_{2})}=K_{2}$ and $0\leq u_{(K_{1},K_{2})}\leq 1$.
4.1. Applications to generalized weighted composition operators on spaces of operator-valued continuous functions
In this subsection, we let $C_{0}(\Omega,\mathcal{C})$ denote the non-unital normed algebra of all continuous $\mathcal{C}$-valued functions on $\Omega$ vanishing at infinity and equipped with the norm
$\|f\|_{\infty}:=\sup_{x\in\Omega}\|f(x)\|,$
and $C_{b}(\Omega,B(H))$ be the unital normed algebra consisting of all uniformly bounded $B(H)$-valued continuous functions on $\Omega$ equipped with the same norm $\|\cdot\|_{\infty}$.
Obviously, $C_{0}(\Omega,\mathcal{C})$ is (left) ideal in $C_{b}(\Omega,B(H))$ and the condition $(E)$ is satisfied. We will therefore denote $C_{0}(\Omega,\mathcal{C})$ and $C_{b}(\Omega,B(H))$ by $\mathcal{A}$ and $A_{1},$ respectively. For every triple $(K_{1},K_{2},m)$ with $K_{1}\subseteq K_{2}\subseteq\Omega,\,K_{1},K_{2}\text{ compact, }m\in\mathbb{N},$ let $\tilde{P}_{(K_{1},K_{2},m)}\in C_{0}(\Omega,\mathcal{C})$ be given by
$\tilde{P}_{(K_{1},K_{2},m)}(x)=u_{(K_{1},K_{2})}(x)P_{m}\quad\text{for all }x\in\Omega.$
We have the following lemma.
Lemma 4.1.
The set $\{\tilde{P}_{(K_{1},K_{2},m)}\}_{K_{1}\subseteq K_{2}\subseteq\Omega,K_{1},K_{2}\text{ compact, }m\in\mathbb{N}}$ satisfies the condition $(\mathcal{P})$.
Proof.
Let $O$ be a non-empty open subset of $\mathcal{A}$ and $f\in O$. Then there exists some $\varepsilon>0$ such that $\varepsilon$– neighbourhood of $f$ is contained in $O$ as well. Since $f\in\mathcal{A},$ there exists some compact subset $K_{1}$ of $\Omega$ such that $\|f(x)\|<\frac{\varepsilon}{2}$ for all $x\in\Omega\setminus K_{1}$. Now, by continuity of $f,$ for each $x\in K_{1}$ we can find some open set $O_{x}$ of $\Omega$ containing $x$ such that $\|f(u)-f(x)\|<\frac{\varepsilon}{2}$ whenever $y\in O_{x}$. Then for all $m\in\mathbb{N}$ and each $y\in O_{x}$ we obtain
$\|(I-P_{m})f(y)\|\leq\|(I-P_{m})(f(y)-f(x))\|+\|(I-P_{m})f(x)\|$
$\leq\|f(y)-f(x)\|+\|(I-P_{m})f(x)\|<\frac{\varepsilon}{2}+\|(I-P_{m})f(x)\|$
By [29, Proposition 2.2.1], since $f(x)\in\mathcal{C},$ we can find sufficiently large $m_{x}\in\mathbb{N}$ such that $\|(I-P_{m})f(x)\|<\frac{\varepsilon}{2}$ for all $m\geq m_{x}.$ Hence, $\|(I-P_{m})f(y)\|<\varepsilon$ each $y\in O_{x}$ and all $m\geq m_{x}.$
The collection $\{O_{x}\}_{x\in K_{1}}$ is an open cover of $K_{1}$, hence there exist a finite subcover $\{O_{x_{1}},\ldots,O_{x_{M}}\}.$ Let $\tilde{m}:=\max\{m_{x_{1}},\ldots,m_{x_{M}}\}.$ Then, by the above we get
$\|\tilde{P}^{3}_{(K_{1},K_{2},m)}(x)f(x)-f(x)\|=\|(I-P_{m})f(x)\|<\varepsilon$ for all $x\in K_{1}$ and $m\geq\tilde{m}$. If $x\in\Omega\setminus K_{1},$ then for all $m\in\mathbb{N}$ we get
$\|\tilde{P}^{3}_{(K_{1},K_{2},m)}(x)f(x)\|+\|f(x)\|=\|u^{3}_{(K_{1},K_{2})}(x)P_{m}f(x)\|+\|f(x)\|\leq 2\|f(x)\|<2\frac{\varepsilon}{2}=\varepsilon.$
Thus,
$\|\tilde{P}^{3}_{(K_{1},K_{2},\tilde{m})}f-f\|_{\infty}<\varepsilon,\text{ so }\tilde{P}^{3}_{(K_{1},K_{2},m)}f\in O.$
Now we consider families $\{\alpha_{n,1}\}_{n\in\mathbb{N}}\dots,\{\alpha_{n,N}\}_{n\in\mathbb{N}}$ of homeomorphisms of $\Omega.$ and families of unitary operators $\{U_{n,1}\}_{n\in\mathbb{N}}\dots,\{U_{n,N}\}_{n\in\mathbb{N}}$ on $H.$ For every $n\in\mathbb{N}$ and $l\in\{1,\ldots,N\}$ we let then
$\Phi_{n,l}:\mathcal{A}_{1}\rightarrow\mathcal{A}_{1}\text{ be given by }\Phi_{t,l}(f):=U_{n,l}^{*}(f\circ\alpha_{n,l})U_{n,l}$
for all $f\in\mathcal{A}.$ Obviously the system $\{\Phi_{n,1}\}_{n\in\mathbb{N}},\dots,\{\Phi_{n,N}\}_{n\in\mathbb{N}}$ is a system of isometric isomorphisms on $\mathcal{A}$ which is disjoint aperiodic with respect to $\{\tilde{P}_{(K_{1},K_{2},m)}\}$ and, clearly, the condition ($R$) is satisfied. Let $\{w_{n,1}\}_{n\in\mathbb{N}},\dots,\{w_{n,N}\}_{n\in\mathbb{N}}$ be families in $\mathcal{A}_{1}$ such that $w_{n,l}(x)$ is invertible for all $x\in\Omega$, $n\in\mathbb{N}$, $l\in\{1,\ldots,N\}$, and such that $\{w^{-1}_{n,1}\}_{n\in\mathbb{N}},\dots,\{w^{-1}_{n,N}\}_{n\in\mathbb{N}}$ also belong to $\mathcal{A}_{1}$ i.e.
$\sup_{x\in\Omega}\|w^{-1}_{n,l}(x)\|<\infty\quad\text{for all }n\in\mathbb{N}\text{ and }l\in\{1,\ldots,N\}.$
For each $n\in\mathbb{N}$, and $l\in\{1,\ldots,N\},$ let then $b_{n,l}:=w_{n,l}U_{n,l}$. Since each $b_{n,l}$ is invertible in $\mathcal{A}_{1}$, it follows that the condition $(C)$ is satisfied in this case.
Corollary 4.2.
Under the above notation and assumptions, if there exist dense subsets $H_{0},H_{1},\ldots,H_{N}$ of $H$ such that for each $s,l\in\{1,\ldots,N\}$ and every compact subset $K$ of $\Omega$ we have that
$\lim_{n\to\infty}\left[\left(\sup_{x\in K}\|w_{n,s}^{-1}(x)h_{s}\|\right)\cdot\left(\sup_{y\in K}\|(w_{n,\ell}\circ\alpha_{n,\ell}^{-1})(y)h_{0}\|\right)\right]=0$
for every $h_{s}\in H_{s}$ and $h_{0}\in H_{0},$ and, in addition, for each distinct $s,l\in\{1,\ldots,N\}$ and every $h_{s}\in H_{s}$ we have
$\lim_{n\to\infty}\left(\sup_{x\in K}\|(w_{n,l}\circ\alpha_{n,l}^{-1}\circ\alpha_{n,s})(x)\,w_{n,s}^{-1}(x)\,h_{s}\|\right)=0,$
then the families $\{T_{n,1}\}_{n\in\mathbb{N}},\ldots,\{T_{n,N}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
Proof.
For each $s\in\{1,\ldots,N\}$, the set consisting of elements of the form
$u_{(K_{1},K_{2})}G_{s,m},$
where $G_{s,m}\in\mathcal{C}$ is such that
$G_{s,m}e_{i}\in H_{s}\text{ for all }j\in\{-m,\ldots,m\}\text{ and }G_{s,m}e_{j}=0\text{ for }j\in\mathbb{Z}\setminus\{-m,\ldots,m\}$
is obviously dense in $\mathcal{A}$ by Lemma 4.1 and the density of $H_{s}$ in $H$. (Here of course $m$ runs through the whole $\mathbb{N}$ and $(K_{1},K_{2})$ runs through all pairs of compact subsets $K_{1},K_{2}$ of $\Omega$ with $K_{1}\subseteq K_{2}.$) Let us denote this set by $\mathcal{A}_{s}$ for each $s\in\{1,\ldots,N\}$. Similarly, we let $\mathcal{A}_{0}$ be the set consisting of all elements of the form $u_{(K_{1},K_{2})}D_{m}$ where $D_{m}\in\mathcal{C}$ is such that
$D_{m}e_{j}\in H_{0}\text{ for all }j\in\{-m,\dots m\}\quad\text{and}\quad D_{m}e_{j}=0\text{ for }j\in\mathbb{Z}\setminus\{-m,\dots m\}.$
By some calculations similar to those in the proof of [20, Theorem 3.2] and [21, Proposition 2.7], it is then not hard to deduce that the conditions of Corollary 3.2 are satisfied. ∎
Example 4.3.
[…]
5. Dynamics on weighted spaces of Radon measures
If $T_{\alpha,w}$ denotes the weighted composition operator on $C_{0}(\Omega)$ with respect to $\alpha$ and $w,$ that is $T_{\alpha,w}(f)=w\cdot(f\circ\alpha)$ for all $f\in C_{0}(\Omega),$ then the adjoint $T_{\alpha,w}^{*}$ is a bounded operator on $M(\Omega)$ where $M(\Omega)$ stands for the Banach space of all Radon measures on $\Omega$ equipped with the total variation norm. It is straightforward to check that
$T_{\alpha,w}^{*}(\mu)(E)=\int_{E}w\circ\alpha^{-1}\,d\mu\ \circ\alpha^{-1}$
for every $\mu\in M(\Omega)$ and every measurable subset $E$ od $\Omega.$ Here $\mu\ \circ\alpha^{-1}(E)=\mu(\alpha^{-1}(E))$ for every $\mu\in M(\Omega)$ and every measurable subset $E$ od $\Omega.$ By $(1)$ and $(2)$ it follows then that for every $n\in\mathbb{N},$ $\mu\in M(\Omega)$ and a Borel measurable subset $E\subseteq\Omega$ we have that
${T_{\alpha,w}^{*n}}(\mu)(E)=\int_{E}\prod_{j=0}^{n-1}w\circ\alpha^{j-n}\,d\mu\ \circ\alpha^{-n}$
and
${S_{\alpha,w}^{*n}}(\mu)(E)=\int_{E}\prod_{j=1}^{n}(w\circ\alpha^{n-j})^{-1}\,d\mu\ \circ\alpha^{n}.$
Throughout this paper, $M_{+}(\Omega)$ will denote the set of all non-negative Radon measures on $\Omega.$ For every Radon measure $\mu$ on $\Omega,$ we let as usual $|\mu|$ denote the total variation of $\mu.$ A weight on $\Omega$ with respect to $\alpha$ is a continuous positive function $\rho$ on $\Omega$ satisfying that
$(\rho\circ\alpha)(t)\leq C\rho(t)\quad\text{and}\quad(\rho\circ\alpha^{-1})(t)\leq C\rho(t)$
for all $t\in\Omega$ and some constant $C>0$.
If $\mu\in M(\Omega)$ and $\rho\in L^{1}(|\mu|)$, we let $\mu^{(\rho)}$ be the Borel measure on $\Omega$ given by
$\mu^{(\rho)}(E)=\int_{E}\rho\,d\mu$
for every Borel subset $E$ of $\Omega$ and we set
$M_{\rho}(\Omega):=\left\{\mu\in M(\Omega)\;\middle|\;\rho\in L^{1}(|\mu|)\text{ and }\mu^{(\rho)}\text{ is regular}\right\}.$
Let $\|\cdot\|_{\rho}:M_{\rho}(\Omega)\to\mathbb{R}^{+}$ be given by
$\|\mu\|_{\rho}=|\mu^{(\rho)}|(\Omega).$
Lemma 5.1.
$M_{\rho}(\Omega)$ is a subspace and $\|\cdot\|_{\rho}$ is a norm on $M_{\rho}(\Omega).$
Proof.
Let $\mu,\nu\in M_{\rho}(\Omega)$ and $\mu=\mu_{1}-\mu_{2}+i(\mu_{3}-\mu_{4})$ be the Jordan decomposition of $\mu.$ It is easy to check that $|\mu+\nu|\leq|\mu|+|\nu|.$ Now, since $\rho\in L^{1}(|\mu|)\cap L^{1}(|\nu|),$ we get that $\rho\in L^{1}(|\mu+\nu|).$ Next, we have that $\mu^{(\rho)}=\mu_{1}^{(\rho)}-\mu_{2}^{(\rho)}+i\big(\mu_{3}^{(\rho)}-\mu_{4}^{(\rho)}\big),$ is the Jordan decomposition of $\mu^{(\rho)}.$
Therefore, $\mu_{1}^{(\rho)}+\mu_{2}^{(\rho)},$ $\mu_{3}^{(\rho)}+\mu_{4}^{(\rho)}\leq|\mu^{(\rho)}|,$ hence, if $|\mu^{(\rho)}|=0,$ then $\mu_{j}^{(\rho)}=0$ for all $j\in\{1,\ldots,4\},$ but this means that $\mu_{j}=0$ for all $j\in\{1,\ldots,4\}$ since $\rho$ is a positive continuous function and $\mu_{j}$ is a Borell measure. Thus, $\mu=0$ in this case. ∎
Lemma 5.2.
Let $K$ be a compact subset of $\Omega,$ $\delta>0$ and $\mu\in M_{\rho}(\Omega).$ If $|\mu^{(\rho)}|(K)<\delta,$ then $|\mu|(K)\leq\dfrac{2\delta}{m_{K}},$ where $m_{K}:=\inf_{x\in K}\rho(x)>0.$
Proof.
Let again $\mu=\mu_{1}-\mu_{2}+i(\mu_{3}-\mu_{4})$ be the Jordan decomposition of $\mu,$ then $\mu^{(\rho)}=\mu_{1}^{(\rho)}-\mu_{2}^{(\rho)}+i(\mu_{3}^{(\rho)}-\mu_{4}^{(\rho)})$ is the Jordan decomposition of $\mu^{(\rho)}.$ If $|\mu^{(\rho)}|(K)<\delta,$ then $(\mu_{1}^{(\rho)}+\mu_{2}^{(\rho)})(K)<\delta$ and $(\mu_{3}^{(\rho)}+\mu_{4}^{(\rho)})(K)<\delta.$ Since $\rho$ is positive and continuous, we must have that $m_{K}>0,$ hence, the above inequalities give $(\mu_{1}+\mu_{2})(K)<\dfrac{\delta}{m_{K}}$ and $(\mu_{3}+\mu_{4})(K)<\dfrac{\delta}{m_{K}}.$ Now, since $|\mu|\leq\sum_{j=1}^{4}\mu_{j},$ the statement follows. ∎
Remark 5.3.
If $\mu\in M_{\rho}(\Omega)$ and $\delta>0,$ it is not hard to check that there exists a compact subset $K\subseteq\Omega$ and $\nu\in M(\Omega)$ such that
$|\nu|(K^{c})=0\quad\text{and}\quad|\mu^{(\rho)}-\nu^{(\rho)}|(\Omega)<\delta.$
Indeed, since $\mu^{(\rho)}$ is a regular Borel measure, there exists a compact subset $K\subseteq\Omega$ such that
$|\mu^{(\rho)}|(K^{c})<\frac{\delta}{2},$
hence we can let $\nu$ be the Borel measure on $\Omega$ given by
$\nu(E):=\mu(E\cap K),$
for every Borel subset $E\subseteq\Omega.$
Lemma 5.4.
The operators $T_{\alpha,w}^{*}$ and $S_{\alpha,w}^{*}$ are bounded linear self-mappings on $M_{\rho}(\Omega).$
Proof.
Let $\mu\in M_{\rho}(\Omega)$ and
$\mu=\mu_{1}-\mu_{2}+i(\mu_{3}-\mu_{4})$
be the Jordan decomposition of $\mu.$ Then $\rho\in\bigcap_{j=1}^{4}L^{1}(\mu_{j}).$ By the change of variables, we obtain
$\int\rho\,(w\circ\alpha^{-1})\,d(\mu_{j}\circ\alpha^{-1})=\int(\rho\circ\alpha)\,w\,d\mu_{j}\leq$
$C\Big(\sup_{t\in\Omega}w(t)\Big)\int\rho\,d\mu_{j}<\infty,$
for all $j\in\{1,2,3,4\},$ because
$(\rho\circ\alpha^{-1})(t)\leq C\rho(t),\quad\text{for all }t\in\Omega,$
by the assumption. Thus, $\rho\in L^{1}\big(T_{\alpha,w}^{*}(\mu_{j})\big)$ for all $j\in\{1,2,3,4\}.$ Since
$T^{*}_{\alpha,w}(\mu)=T^{*}_{\alpha,w}(\mu_{1})-T^{*}_{\alpha,w}(\mu_{2})+i\Big(T^{*}_{\alpha,w}(\mu_{3})-T^{*}_{\alpha,w}(\mu_{4})\Big)$
is the Jordan decomposition of $T^{*}_{\alpha,w}(\mu)$ by the arguments from the proof of [23, Proposition 10.1] part $(i)\Rightarrow(ii)$, it follows that $\rho\in L^{1}\big(|T_{\alpha,w}^{*}(\mu)|\big).$
Next, given $\epsilon>0$ and a Borel subset $E$ of $\Omega,$ choose some open subset $\mathcal{O}\subseteq E$ and a compact subset $K\subseteq\Omega$ such that $K$ and $\alpha^{-1}(E)$ are subsets of $\mathcal{O}$ and
$|\mu^{(\rho)}|(\mathcal{O}\setminus\alpha^{-1}(E)),|\mu^{(\rho)}|(\mathcal{O}\setminus K)<\varepsilon.$
Since $\mu^{(\rho)}=\mu_{1}^{(\rho)}-\mu_{2}^{(\rho)}+i\big(\mu_{3}^{(\rho)}-\mu_{4}^{(\rho)}\big)$ is the Jordan decomposition of $\mu^{(\rho)}$, we have that $|\mu_{j}^{(\rho)}|\leq|\mu^{(\rho)}|$ for all $j\in\{1,\dots,4\}$, hence
$|\mu_{j}^{(\rho)}|\big(\mathcal{O}\setminus\alpha^{-1}(E)\big)<\varepsilon$
and
$|\mu_{j}^{(\rho)}|\big(\mathcal{O}\setminus K\big)<\varepsilon$
for all $j\in\{1,\dots,4\}$. By the previous inequality, we get that
$\varepsilon\,C\Big(\sup_{t\in\Omega}w(t)\Big)>C\Big(\sup_{t\in\Omega}w(t)\Big)\int_{\mathcal{O}\setminus\alpha^{-1}(E)}\rho\,d\mu_{j}\geq\int_{\alpha(\mathcal{O})\setminus(E)}\rho\big(w\circ\alpha^{-1}\big)\,d\big(\mu_{j}\circ\alpha^{-1}\big),$
and similarly,
$\varepsilon\,C\Big(\sup_{t\in\Omega}w(t)\Big)\geq\int_{\alpha(\mathcal{O})\setminus\alpha(K)}\rho\big(w\circ\alpha^{-1}\big)\,d\big(\mu_{j}\circ\alpha^{-1}\big)$
for all $j\in\{1,\dots,4\}$.
Since
$\big(T^{*}_{\alpha,w}(\mu)\big)^{(\rho)}=\big(T^{*}_{\alpha,w}(\mu_{1})\big)^{(\rho)}-\big(T^{*}_{\alpha,w}(\mu_{2})\big)^{(\rho)}+i\Big(\big(T^{*}_{\alpha,w}(\mu_{3})\big)^{(\rho)}-\big(T^{*}_{\alpha,w}(\mu_{4})\big)^{(\rho)}\Big)$
is the Jordan decomposition of $\big(T^{*}_{\alpha,w}(\mu)\big)^{(\rho)}$, we deduce that
$\big|\big(T^{*}_{\alpha,w}(\mu)\big)^{(\rho)}\big|\big(\alpha(\mathcal{O})\setminus E)\big)<4\varepsilon\,C\Big(\sup_{t\in\Omega}w(t)\Big)$
and
$\big|\big(T^{*}_{\alpha,w}(\mu)\big)^{(\rho)}\big|\big(\alpha(\mathcal{O})\setminus\alpha(K)\big)<4\varepsilon\,C\Big(\sup_{t\in\Omega}w(t)\Big).$
Now, $\alpha(\mathcal{O})$ is open and $\alpha(K)$ is compact because $\alpha$ is a homeomorphism, and, moreover, $E\subseteq\alpha(\mathcal{O}),$ hence $\big(T^{*}_{\alpha,w}(\mu)\big)^{(\rho)}$ is regular.
Finally, by the above,
$\big|\big(T^{*}_{\alpha,w}(\mu)\big)^{(\rho)}\big|\leq 4C\Big(\sup_{t\in\Omega}w(t)\Big)|\mu^{(\rho)}|.$
Theorem 5.5.
Under the above notation and assumptions, the following statements are equivalent.
  1. (1)
    The set of periodic elements of $T_{\alpha,\omega}^{*}$ is dense in $M_{\rho}(\Omega)$.
  2. (2)
    For every compact subset $K$ of $\Omega$ and every measure $\mu\in M_{\rho}(\Omega)$ with $|\mu|(K^{c})=0$, there exist a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ and sequences $\{A_{k}\}_{k\in\mathbb{N}},\{B_{k}\}_{k\in\mathbb{N}}$ of Borel subsets of $K$ such that $\alpha^{n_{k}}(K)\cap K=\varnothing$ for all $k\in\mathbb{N}$ and
    $\lim_{k\to\infty}|\mu|(B_{k})=\lim_{k\to\infty}|\mu|(A_{k})=0,$
    and
    $\lim_{k\to\infty}\sup_{t\in A_{k}^{c}\cap K}\sum_{\ell=1}^{\infty}(\rho\circ\alpha^{\ell n_{k}})(t)\left(\prod_{r=0}^{\ell n_{k}-1}(w\circ{\alpha^{r}})(t)\right)=$
    $\lim_{k\to\infty}\sup_{t\in B_{k}^{c}\cap K}\sum_{\ell=1}^{\infty}(\rho\circ\alpha^{-\ell n_{k}})(t)\left(\prod_{r=1}^{\ell n_{k}}(w\circ{\alpha^{-r}})\right)^{-1}(t)=0.$
Proof.
First we will prove $1.\Rightarrow 2.$ If 1. holds, that is if the set of periodic elements of $T_{\alpha,\omega}^{*}$ is dense in $M_{\rho}(\Omega)$, then given a compact subset $K\subseteq\Omega$ and $\mu\in M_{\rho}(\Omega)$ with $|\mu|(K^{c})=0$, for each $k\in\mathbb{N}$ we can find some $\gamma_{k}\in M_{\rho}(\Omega)$ with
$T_{\alpha,\omega}^{*\,\ell n_{k}}(\gamma_{k})=\gamma_{k}\quad\text{for all }\ell\in\mathbb{N}\quad\text{ and some }n_{k}\in\mathbb{N},$
and such that
$|\gamma_{k}^{(\rho)}-\mu^{(\rho)}|(\Omega)<\frac{1}{4^{k}}.$
Then
$T_{\alpha,\omega}^{*\,\ell(mn_{k})}(\gamma_{k})=\gamma_{k}\quad\text{for all }\ell,m\in\mathbb{N},$
hence the sequence $\{n_{k}\}_{1}$ can be chosen to be strictly increasing, and such that
$\alpha^{r}(K)\cap K=\varnothing$
for all $r\geq n_{k}$ due to the fact that $\alpha$ is aperiodic. Notice that this gives that
$\alpha^{j_{1}n_{k}}(K)\cap\alpha^{j_{2}n_{k}}(K)=\varnothing$
for each distinct $j_{1},j_{2}\in\mathbb{Z}$ and all $k\in\mathbb{N}$. Moreover, since $|\mu|(K^{c})=0$, then
$|\mu^{(p)}|(K^{c})=0,$
hence
$|\gamma_{k}^{(p)}|(K^{c})<\frac{1}{4^{k}}\quad\text{for all }k\in\mathbb{N}.$
Taking all this into account, for each $k\in\mathbb{N}$ we then get
$|\gamma_{k}^{(\rho)}|(K^{c})\geq\sum_{\ell=1}^{\infty}|\gamma_{k}^{(\rho)}|\big(\alpha^{\ell n_{k}}(K)\big)+\sum_{\ell=1}^{\infty}|\gamma_{k}^{(\rho)}|\big(\alpha^{-\ell n_{k}}(K)\big)$
$\sum_{\ell=1}^{\infty}\left|\big(T_{\alpha,\omega}^{*\,\ell n_{k}}(\gamma_{k})\big)^{(\rho)}\right|\big(\alpha^{\ell n_{k}}(K)\big)+\sum_{\ell=1}^{\infty}\left|\big(S_{\alpha,\omega}^{*\,\ell n_{k}}(\gamma_{k})\big)^{(\rho)}\right|\big(\alpha^{-\ell n_{k}}(K)\big).$
For each $k,\ell\in\mathbb{N}$ and $j\in\{1,2\}$, let $\eta_{j,\pm}^{(k,\ell)}\in M(\Omega)$ be given by
$\eta_{j,\pm}^{(k,\ell)}(E)=\int_{E}\prod_{r=1}^{\ell n_{k}}\big(w\circ\alpha^{\ell n_{k}-r}\big)^{-1}\,d\big(\gamma_{j,\pm}^{(k)}\circ\alpha^{\ell n_{k}}\big)$
for every Borel subset $E$ of $\Omega$, where
$\gamma_{k}=\gamma_{1,+}^{(k)}-\gamma_{1,-}^{(k)}+i\big(\gamma_{2,+}^{(k)}-\gamma_{2,-}^{(k)}\big)$
is the Jordan decomposition of $\gamma_{k}$.
Then, by the same arguments as in the proof of [23, Proposition 10.1] part $(i)\Rightarrow(ii)$, we have that
$S_{\alpha,\omega}^{*\,\ell n_{k}}(\gamma_{k})=\eta_{1,+}^{(k,\ell)}-\eta_{1,-}^{(k,\ell)}+i\big(\eta_{2,+}^{(k,\ell)}-\eta_{2,-}^{(k,\ell)}\big)$
is the Jordan decomposition of $S_{\alpha,\omega}^{*\,\ell n_{k}}(\gamma_{k})$.
Hence,
$\big(\eta_{1,+}^{(k,\ell)}\big)^{(\rho)}-\big(\eta_{1,-}^{(k,\ell)}\big)^{(\rho)}+i\left[\big(\eta_{2,+}^{(k,\ell)}\big)^{(\rho)}-\big(\eta_{2,-}^{(k,\ell)}\big)^{(\rho)}\right]$
is the Jordan decomposition of $\big(S_{2,\omega}^{*\,\ell n_{k}}(\gamma_{k})\big)^{(\rho)}$. Therefore,
$\big(\eta_{j,+}^{(k,\ell)}\big)^{(\rho)}+\big(\eta_{j,-}^{(k,\ell)}\big)^{(\rho)}\leq\left|\big(S_{2,\omega}^{*\,\ell n_{k}}(\gamma_{k})\big)^{(\rho)}\right|\quad\text{for }j\in\{1,2\}.$
Hence,
$\sum_{\ell=1}^{\infty}\Big((\eta_{j,+}^{(k,\ell)})^{(\rho)}+(\eta_{j,-}^{(k,\ell)})^{(\rho)}\Big)\big(\alpha^{-\ell n_{k}}(K)\big)<\frac{1}{4^{k}}$
for all $k\in\mathbb{N}$ and $j\in\{1,2\}$. Thus,
$\frac{1}{4^{k}}>\sum_{\ell=1}^{\infty}\int_{\alpha^{-\ell n_{k}}(K)}\rho\,\Big(\prod_{r=1}^{\ell n_{k}}\big(w\circ\alpha^{\ell n_{k}-r}\big)\Big)^{-1}\,d\big((\gamma_{j,+}^{(k)}+\gamma_{j,-}^{(k)})\circ\alpha^{\ell n_{k}}\big)=$
$\sum_{\ell=1}^{\infty}\int_{K}(\rho\circ\alpha^{-\ell n_{k}})\Big(\prod_{r=1}^{\ell n_{k}}(w\circ\alpha^{-r})\Big)^{-1}\,d\big(\gamma_{j,+}^{(k)}+\gamma_{j,-}^{(k)}\big)=$
$\int_{K}\sum_{\ell=1}^{\infty}(\rho\circ\alpha^{-\ell n_{k}})\Big(\prod_{r=1}^{\ell n_{k}}(w\circ\alpha^{-r})\Big)^{-1}\,d\big(\gamma_{j,+}^{(k)}+\gamma_{j,-}^{(k)}\big)$
for each $k\in\mathbb{N}$ and $j\in\{1,2\}$, where in the first equality we have applied the change of variables and in the second equality we have applied the Monotone Convergence Theorem.
For each $k\in\mathbb{N}$, let
$B_{k}=\left\{t\in K\ \middle|\ \sum_{\ell=1}^{\infty}(\rho\circ\alpha^{-\ell n_{k}})(t)\Big(\prod_{r=1}^{\ell n_{k}}(w\circ\alpha^{-r})\Big)^{-1}(t)>\frac{1}{2^{k}}\right\}.$
Since
$|\gamma_{k}^{(\rho)}-\mu^{(\rho)}|(K)<\frac{1}{4^{k}},$
by Lemma 5.2 we have
$|\gamma_{k}-\mu|(K)<\frac{2}{m_{K}4^{k}},$
hence, by similar arguments as in the proof of [23, Proposition 10.1] part $(i)\Rightarrow(ii)$, we can deduce that
$|\mu|(B_{k})<\frac{1}{2^{k-1}}+\frac{2}{m_{K}4^{k}}\quad\text{for all }k\in\mathbb{N}.$
Next, by observing that
$T_{\alpha,\omega}^{*\,\ell n_{k}}(\gamma_{k})=\Theta_{1,+}^{(k,\ell)}-\Theta_{1,-}^{(k,\ell)}+i\big(\Theta_{2,+}^{(k,\ell)}-\Theta_{2,-}^{(k,\ell)}\big)$
is the Jordan decomposition of $T_{\alpha,\omega}^{*\,\ell n_{k}}(\gamma_{k})$, where
$\Theta_{j,\pm}^{(k,\ell)}(E)=\int_{E}\prod_{r=0}^{\ell n_{k}-1}\big(w\circ\alpha^{r-\ell n_{k}}\big)\,d\big(\gamma_{j,\pm}^{(k)}\circ\alpha^{-\ell n_{k}}\big)$
for each $\ell,k\in\mathbb{N}$, $j\in\{1,2\}$ and every Borel subset $E$ of $\Omega$, and using that
$\sum_{\ell=1}^{\infty}\left|\big(T_{\alpha,\omega}^{*\,\ell n_{k}}(\gamma_{k})\big)^{(p)}\right|\big(\alpha^{\ell n_{k}}(K)\big)<\frac{1}{4^{k}},$
in a similar way as above we obtain that
$|\mu|(A_{k})<\frac{1}{2^{k-1}}+\frac{2}{m_{K}4^{k}},$
where
$A_{k}:=\left\{t\in K\ \middle|\ \sum_{\ell=1}^{\infty}(\rho\circ\alpha^{\ell n_{k}})(t)\prod_{r=0}^{\ell n_{k}-1}\big(w\circ\alpha^{\,r}\big)(t)>\frac{1}{2^{k}}\right\}$
for all $k\in\mathbb{N}$. This proves $1.\Rightarrow 2.$
Next we will prove $2.\Rightarrow 1.$ Suppose that 2. holds. Given an open non-empty subset $O$ of $M_{\rho}(\Omega)$, by Remark 5.3 there exists a compact subset $K\subseteq\Omega$ and $\mu\in O$ such that $|\mu|(K^{c})=0$. Choose a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ and sequences of Borel subsets $\{A_{n}\}_{n},\{B_{k}\}_{k}$ of $K$ satisfying the assumption in (2) with respect to $\mu$ and $K$.
Let
$C_{k}=A_{k}\cup B_{k}\quad\text{for each }k\in\mathbb{N},$
and let $\mu_{k}\in M_{\rho}(\Omega)$ be given by
$\mu_{k}(E)=\mu(E\cap C_{k}^{c})$
for every Borel subset $E$ of $\Omega$. Then, $|\mu_{k}|(K^{c}\cup C_{k})=0$ for each $k\in\mathbb{N}$.
Let
$\mu_{k}=\mu_{1,+}^{(k)}-\mu_{1,-}^{(k)}+i\big(\mu_{2,+}^{(k)}-\mu_{2,-}^{(k)}\big)$
be the Jordan decomposition of $\mu_{k}$ for each $k\in\mathbb{N}$. As above in the proof of the implication $(1)\Rightarrow(2)$, if
$\big(T_{\alpha,\omega}^{*\,\ell n_{k}}(\mu_{k})\big)^{(\rho)}=\widetilde{\Theta}_{1,+}^{(k,\ell)}-\widetilde{\Theta}_{1,-}^{(k,\ell)}+i\big(\widetilde{\Theta}_{2,+}^{(k,\ell)}-\widetilde{\Theta}_{2,-}^{(k,\ell)}\big)$
and
$\big(S_{\alpha,\omega}^{*\,\ell n_{k}}(\mu_{k})\big)^{(\rho)}=\widetilde{\eta}_{1,+}^{(k,\ell)}-\widetilde{\eta}_{1,-}^{(k,\ell)}+i\big(\widetilde{\eta}_{2,+}^{(k,\ell)}-\widetilde{\eta}_{2,-}^{(k,\ell)}\big),$
are the Jordan decompositions of $\big(T_{\alpha,\omega}^{*\,\ell n_{k}}(\mu_{k})\big)^{(\rho)}$ and $\big(S_{\alpha,\omega}^{*\,\ell n_{k}}(\mu_{k})\big)^{(\rho)}$, respectively, for each $\ell,k\in\mathbb{N}$, then
$\widetilde{\Theta}_{j,\pm}^{(k,\ell)}(E)=\int_{E}\rho\Big(\prod_{r=0}^{\ell n_{k}-1}(w\circ\alpha^{\,r-n_{k}})\Big)\,d\big(\mu_{j,\pm}^{(k)}\circ\alpha^{-\ell n_{k}}\big)$
and
$\widetilde{\eta}_{j,\pm}^{(k,\ell)}(E)=\int_{E}\rho\Big(\prod_{r=1}^{\ell n_{k}}(w\circ\alpha^{\ell n_{k}-r})^{-1}\Big)\,d\big(\mu_{j,\pm}^{(k)}\circ\alpha^{\ell n_{k}}\big)$
for each $k,\ell\in\mathbb{N},j\in\{1,2\}$ and every Borel subset $E$ of $\Omega$. Therefore, for each $k\in\mathbb{N}$,
$\sum_{\ell=1}^{\infty}\left|\big(T_{\alpha,\omega}^{*\,\ell n_{k}}(\mu_{k})\big)^{(\rho)}\right|(\Omega)+\sum_{\ell=1}^{\infty}\left|\big(S_{\alpha,\omega}^{*\,\ell n_{k}}(\mu_{k})\big)^{(\rho)}\right|(\Omega)$
$\leq\sum_{j=1}^{2}\sum_{\ell=1}^{\infty}\big(\widetilde{\Theta}_{j,+}^{(k,\ell)}+\widetilde{\Theta}_{j,-}^{(k,\ell)}\big)(\Omega)+\sum_{j=1}^{2}\sum_{\ell=1}^{\infty}\big(\widetilde{\eta}_{j,+}^{(k,\ell)}+\widetilde{\eta}_{j,-}^{(k,\ell)}\big)(\Omega)$
$\sum_{j=1}^{2}\sum_{\ell=1}^{\infty}\int_{\Omega}(\rho\circ\alpha^{\ell n_{k}})\Big(\prod_{r=0}^{\ell n_{k}-1}(w\circ\alpha^{r})\Big)\,d\big(\mu_{j,+}^{(k)}+\mu_{j,-}^{(k)}\big)$
$\sum_{j=1}^{2}\sum_{\ell=1}^{\infty}\int_{\Omega}(\rho\circ\alpha^{-\ell n_{k}})\Big(\prod_{r=1}^{\ell n_{k}}(w\circ\alpha^{-r})^{-1}\Big)\,d\big(\mu_{j,+}^{(k)}+\mu_{j,-}^{(k)}\big)$
$\sum_{j=1}^{2}\sum_{\ell=1}^{\infty}\int_{C_{k}^{c}\cap K}(\rho\circ\alpha^{\ell n_{k}})\Big(\prod_{r=0}^{\ell n_{k}-1}(w\circ\alpha^{r})\Big)\,d\big(\mu_{j,+}^{(k)}+\mu_{j,-}^{(k)}\big)$
$\sum_{j=1}^{2}\sum_{\ell=1}^{\infty}\int_{C_{k}^{c}\cap K}(\rho\circ\alpha^{-\ell n_{k}})\Big(\prod_{r=1}^{\ell n_{k}}(w\circ\alpha^{-r})^{-1}\Big)\,d\big(\mu_{j,+}^{(k)}+\mu_{j,-}^{(k)}\big)$
$\sum_{j=1}^{2}\int_{C_{k}^{c}\cap K}\sum_{\ell=1}^{\infty}(\rho\circ\alpha^{\ell n_{k}})\Big(\prod_{r=0}^{\ell n_{k}-1}(w\circ\alpha^{r})\Big)\,d\big(\mu_{j,+}^{(k)}+\mu_{j,-}^{(k)}\big)$
$\sum_{j=1}^{2}\int_{C_{k}^{c}\cap K}\sum_{\ell=1}^{\infty}(\rho\circ\alpha^{-\ell n_{k}})\Big(\prod_{r=1}^{\ell n_{k}}(w\circ\alpha^{-r})^{-1}\Big)\,d\big(\mu_{j,+}^{(k)}+\mu_{j,-}^{(k)}\big)$
$\leq\sum_{j=1}^{2}\sup_{t\in C_{k}^{c}\cap K}\sum_{\ell=1}^{\infty}(\rho\circ\alpha^{\ell n_{k}})(t)\Big(\prod_{r=0}^{\ell n_{k}-1}(w\circ\alpha^{r})(t)\Big)\,(\mu_{j,+}^{(k)}+\mu_{j,-}^{(k)})(\Omega)$
$\sum_{j=1}^{2}\sup_{t\in C_{k}^{c}\cap K}\sum_{\ell=1}^{\infty}\Big((\rho\circ\alpha^{-\ell n_{k}})(t)\Big)\Big(\prod_{r=1}^{\ell n_{k}}(w\circ\alpha^{-r})^{-1}(t)\Big)\,(\mu_{j,+}^{(k)}+\mu_{j,-}^{(k)})(\Omega)$
$\leq 2\sup_{t\in A_{k}^{c}\cap K}\sum_{\ell=1}^{\infty}\Big((\rho\circ\alpha^{\ell n_{k}})(t)\Big)\Big(\prod_{r=0}^{\ell n_{k}-1}(w\circ\alpha^{r})(t)\Big)\,|\mu|(\Omega)$
$2\sup_{t\in B_{k}^{c}\cap K}\sum_{\ell=1}^{\infty}\Big((\rho\circ\alpha^{-\ell n_{k}})(t)\Big)\Big(\prod_{r=1}^{\ell n_{k}}(w\circ\alpha^{-r})^{-1}(t)\Big)\,|\mu|(\Omega)\longrightarrow 0,\quad k\to\infty,$
where in the first equality we have applied the change of variables, in the second equality we have used the fact that $|\mu_{k}|(K^{c}\cup C_{k})=0$ for each $k\in\mathbb{N}$, in the third equality we have applied the Monotone Convergence Theorem, and in the last inequality we have used the fact that
$(\mu_{j,+}^{(k)}+\mu_{j,-}^{(k)})(\Omega)\leq|\mu^{(k)}|(\Omega)\leq|\mu|(\Omega)$
for all $k\in\mathbb{N}$ and $j\in\{1,2\}$.
Further, if
$\mu=\mu_{1,+}-\mu_{1,-}+i(\mu_{2,+}-\mu_{2,-})$
is the Jordan decomposition of $\mu$, and $\widetilde{\mu}_{k,j,\pm}^{(p)}$ is the Borel measure on $\Omega$ given by
$\widetilde{\mu}_{k,j,\pm}^{(\rho)}(E)=\int_{E\cap C_{k}}\rho\,d\mu_{j,\pm},$
for every Borel subset $E$ of $\Omega,$ $k\in\mathbb{N}$ and $j\in\{1,2\}$, then, clearly,
$\widetilde{\mu}_{k,1,+}^{(\rho)}-\widetilde{\mu}_{k,1,-}^{(\rho)}+i\big(\widetilde{\mu}_{k,2,+}^{(\rho)}-\widetilde{\mu}_{k,2,-}^{(\rho)}\big)$
is the Jordan decomposition of $\mu^{(\rho)}-\mu_{k}^{(\rho)}$ for each $k\in\mathbb{N}$.
Since
$\int_{C_{k}}\rho\,d\mu_{j,\pm}\leq\sup_{t\in K}\rho(t)\,\mu_{j,\pm}(C_{k})\leq\sup_{t\in K}\rho(t)\,|\mu|(C_{k})\longrightarrow 0,\quad k\to\infty,$
we deduce that
$\big\|\mu^{(p)}-\mu_{k}^{(p)}\big\|(\Omega)\longrightarrow 0,\quad k\to\infty.$
As in the proof of [23, Proposition 10.3], for each $k\in\mathbb{N},$ set
$v_{k}=\mu_{k}+\sum_{l=1}^{\infty}T_{\alpha,w}^{*^{ln_{k}}}(\mu_{k})+\sum_{l=1}^{\infty}S_{\alpha,w}^{*^{ln_{k}}}(\mu_{k}).$
It is straightforward to check that $v_{k}=T_{\alpha,w}^{*ln_{k}}v_{k}$ for all $l,k\in\mathbb{N}.$ Moreover,
$v_{k}\longrightarrow\mu,\quad k\to\infty.$
This proves $2.\Rightarrow 1.$
Let $N\in\mathbb{N}$ and $\{\alpha_{z,1}\}_{z\in\mathbb{Z}},\dots,\{\alpha_{z,N}\}_{z\in\mathbb{Z}}$ be families of homeomorphisms of $\Omega$ which are $d\mathcal{F}$-aperiodic, that is for each compact subset $K\subseteq\Omega$ and every $E\in\mathcal{F}$, there exists some $F\subseteq E$ with $F\in\mathcal{F}$ such that
$(\alpha_{z,\ell}\circ\alpha_{z,s}^{-1})(K)\cap K=\varnothing$
for every $z\in F$ and each distinct $s,\ell\in\{1,\dots,N\}$ and in addition
$\alpha_{z,\ell}(K)\cap K=\varnothing$
for all $z\in F$ and $\ell\in\{1,\dots,N\}$.
We will assume that $\rho$ is a weight with respect to $\alpha_{z,\ell}$ for all $z\in\mathbb{Z}$ and $\ell\in\{1,\dots,N\}$.
Let $\{w_{z,1}\}_{z\in\mathbb{Z}},\dots,\{w_{z,N}\}_{z\in\mathbb{Z}}$ be families of continuous bounded positive functions on $\Omega$ such that $w_{z,\ell}^{-1}$ is also bounded for each $z\in\mathbb{Z}$ and $\ell\in\{1,\dots,N\}$.
We let
$T^{*}_{z,\ell}:=T^{*}_{\alpha_{z,\ell},\,w_{z,\ell}}\quad\text{for all }z\in\mathbb{Z}\text{ and }\ell\in\{1,\dots,N\}.$
Theorem 5.6.
Under the above notation and assumptions, the following are equivalent.
(1) The families $\{T^{*}_{z,1}\}_{z\in\mathbb{Z}},\dots,\{T^{*}_{z,N}\}_{z\in\mathbb{Z}}$ are $d\mathcal{F}$-semi-transitive on $M_{\rho}(\Omega)$.
(2) For each compact subset $K\subseteq\Omega,\varepsilon>0$ and any measures $\mu,\nu_{1},\dots,\nu_{N}\in M_{\rho}(\Omega)$ with
$|\mu|(K^{c})=|\nu_{1}|(K^{c})=\dots=|\nu_{N}|(K^{c})=0,$
there exists some $F\in\mathcal{F}$ and families of Borel subsets of $K$
$\{A_{z}\}_{z\in F},\quad\{B_{z,1}\}_{z\in F},\dots,\{B_{z,N}\}_{z\in F}$
such that
$|\mu|(A_{z})<2N\sqrt{\varepsilon}\left(1+\frac{\sqrt{\varepsilon}}{m_{K}}\right),$
$|\nu_{s}|(B_{z,s})<2(N+1)\sqrt{\varepsilon}\left(1+\frac{\sqrt{\varepsilon}}{m_{K}}\right),$
and
$\left(\sup_{t\in A_{z}^{c}\cap K}\big((\rho\circ\alpha_{z,\ell})(t)\,w_{z,\ell}(t)\big)\right)\left(\sup_{t\in B_{z,s}^{c}\cap K}\big((\rho\circ\alpha^{-1}_{z,s})(t)\,(w_{z,s}\circ\alpha^{-1}_{z,s})(t)\big)\right)<\varepsilon$
for all $z\in F$ and $\ell,s\in\{1,\dots,N\}$. Moreover, for all $z\in F$ and each distinct $s,\ell\in\{1,\dots,N\}$, it holds that
$\sup_{t\in B_{z,s}^{c}\cap K}\left((\rho\circ\alpha_{z,\ell}\circ\alpha^{-1}_{z,s})(t)\,\frac{\big(w_{z,\ell}\circ\alpha^{-1}_{z,s}\big)(t)}{\big(w_{z,s}\circ\alpha^{-1}_{z,s}\big)(t)}\right)<\sqrt{\varepsilon}.$
Proof.
First we will prove $(1)\Rightarrow(2)$. Given $\varepsilon>0,$ a compact subset $K\subseteq\Omega$ and measures $\mu,\nu_{1},\dots,\nu_{N}\in M(\Omega)$ such that
$0=|\mu|(K^{c})=|\nu_{1}|(K^{c})=\dots=|\nu_{N}|(K^{c}),$
we can find some $E\in\mathcal{F}$ and some $\lambda_{z}\in\mathbb{R}^{+}$ such that
$\frac{1}{\lambda_{z}}S^{*}_{z,1}\big(B(\nu_{1},\varepsilon)\big)\cap\dots\cap\frac{1}{\lambda_{z}}S^{*}_{z,N}\big(B(\nu_{N},\varepsilon)\big)\cap B(\mu,\varepsilon)\neq\varnothing$
for all $z\in E$, where $B(\nu_{\ell},\varepsilon)$ is the open ball in $M_{\rho}(\Omega)$ with centre in $\nu_{\ell}$ and radius $\varepsilon$ for all $\ell\in\{1,\dots,N\}$. Further, since $\{\alpha_{z,1}\}_{z\in\mathbb{R}},\dots,\{\alpha_{z,N}\}_{z\in\mathbb{R}}$ is a $d\mathcal{F}$-aperiodic system, we can find some $F\subseteq E$ with $F\in\mathcal{F}$ such that
$(\alpha_{z,\ell}\circ\alpha_{z,s}^{-1})(K)\cap K=\varnothing$
for every $z\in F$ and each distinct $s,\ell\in\{1,\dots,N\}$ and in addition
$\alpha_{z,\ell}(K)\cap K=\varnothing$
for all $z\in F$ and $\ell\in\{1,\dots,N\}$. Next, for each $z\in F$ there exists some $\eta_{z}\in M(\Omega)$ such that
$\big|\eta_{z}^{(\rho)}-\mu^{(\rho)}\big|(\Omega)<\varepsilon\quad\text{and}\quad\big|\lambda_{z}\gamma_{z,\ell}^{(\rho)}-\nu_{\ell}^{(\rho)}\big|(\Omega)<\varepsilon\quad\text{for all }\ell\in\{1,\dots,N\},$
where
$\gamma_{z,\ell}^{(\rho)}=\big(T^{*}_{z,\ell}(\eta_{z})\big)^{(\rho)},$
that is
$\gamma_{z,\ell}^{(\rho)}(E)=\int_{E}\rho\big(w_{z,\ell}\circ\alpha_{z,\ell}^{-1}\big)\,\,d\big(\eta_{z}\circ\alpha_{z,\ell}^{-1}\big)$
for every measurable subset $E$ of $\Omega$.
Observe that for each distinct $s,\ell\in\{1,\dots,N\}$ and all $z\in F$ we have
$T^{*}_{z,\ell}\Big(S^{*}_{z,s}\big(B(\nu_{s},\varepsilon)\big)\Big)\cap B(\nu_{\ell},\varepsilon)\neq\varnothing,$
since
$S^{*}_{z,s}\big(B(\nu_{s},\varepsilon)\big)\cap S^{*}_{z,\ell}\big(B(\nu_{\ell},\varepsilon)\big)\neq\varnothing.$
Hence, for each $z\in F$, we can find some $\beta_{z,s}\in B(\nu_{s},\varepsilon)$ with
$\Big|\big(T^{*}_{z,\ell}(S^{*}_{z,s}(\beta_{z,s}))\big)^{(\rho)}-\nu_{\ell}^{(\rho)}\Big|(\Omega)<\varepsilon.$
Let
$\Theta_{z,\ell,s}:=T^{*}_{z,\ell}\big(S^{*}_{z,s}(\beta_{z,s})\big).$
We notice that for each Borel subset $E$ of $\Omega$ we have
$\Theta_{z,\ell,s}^{(\rho)}(E)=\int_{E}\big(w_{z,\ell}\circ\alpha_{z,\ell}^{-1}\big)\,\rho\;d\Big(S^{*}_{z,s}(\beta_{z,s})\circ\alpha_{z,\ell}^{-1}\Big)=$
$\int_{\alpha_{z,\ell}^{-1}(E)}(\rho\circ\alpha_{z,\ell})\,w_{z,\ell}\;d\Big(S^{*}_{z,s}(\beta_{z,s})\Big)=\int_{\alpha_{z,\ell}^{-1}(E)}(\rho\circ\alpha_{z,\ell})\,w_{z,\ell}\,w_{z,s}^{-1}\;d\big(\beta_{z,s}\circ\alpha_{z,s}\big)=$
$\int_{E}\big(w_{z,\ell}\circ\alpha_{z,\ell}^{-1}\big)\,\big(w_{z,s}\circ\alpha_{z,\ell}^{-1}\big)^{-1}\,\rho\;d\big(\beta_{z,s}\circ\alpha_{z,s}\circ\alpha_{z,\ell}^{-1}\big).$
By using that $\alpha_{z,s}\circ\alpha_{z,\ell}^{-1}$ is a homeomorphism of $\Omega$ with
$(\alpha_{z,\ell}\circ\alpha_{z,s}^{-1})(K)\cap K=\varnothing,$
together with the fact that
$|\beta_{z,s}^{(\rho)}-\nu_{s}^{(\rho)}|(K^{c})=|\beta_{z,s}^{(\rho)}|(K^{c})<\varepsilon,$
by exactly the same arguments as in the proof of [23, Proposition 10.1] part $(i)\Rightarrow(ii)$, we get that
$\Theta_{z,\ell,s,j,\pm}^{(\rho)}(E)=\int_{E}\rho\big(w_{z,\ell}\circ\alpha_{z,\ell}^{-1}\big)\big(w_{z,s}\circ\alpha_{z,\ell}^{-1}\big)^{-1}\,d\big(\beta_{z,s,j,\pm}\circ\alpha_{z,s}^{-1}\circ\alpha_{z,\ell}^{-1}\big)$
for $j\in\{1,2\}$, where
$\Theta_{z,\ell,s}^{(\rho)}=\Theta_{z,\ell,s,1,+}^{(\rho)}-\Theta_{z,\ell,s,1,-}^{(\rho)}+i(\Theta_{z,\ell,s,2,+}^{(\rho)}-\Theta_{z,\ell,s,2,-}^{(\rho)})$
and
$\beta_{z,s}=\beta_{z,s,1,+}-\beta_{z,s,1,-}+i(\beta_{z,s,2,+}-\beta_{z,s,2,-}),$
are Jordan decompositions of $\Theta_{z,\ell,s}^{(\rho)}$ and $\beta_{z,s},$ respectively. Since
$|\Theta_{z,\ell,s}^{(\rho)}|(K^{c})=\big|\Theta_{z,\ell,s}^{(\rho)}-\nu_{\ell}^{(\rho)}\big|(K^{c})<\varepsilon$
and
$(\alpha_{z,\ell}\circ\alpha_{z,s}^{-1})(K)\cap K=\varnothing,$
we must have that
$\varepsilon>\int_{\alpha_{z,\ell}(\alpha_{z,s}^{-1}(K))}\rho\big(w_{z,\ell}\circ\alpha_{z,\ell}^{-1}\big)\big(w_{z,s}\circ\alpha_{z,\ell}^{-1}\big)^{-1}\,d\big((\beta_{z,s,j,+}+\beta_{z,s,j,-})\circ\alpha_{z,s}\circ\alpha_{z,\ell}^{-1}\big)=$
$\int_{K}(\rho\circ\alpha_{z},l\circ\alpha_{z,s}^{-1})\big(w_{z,\ell}\circ\alpha_{z,s}^{-1}\big)\big(w_{z,s}\circ\alpha_{z,s}^{-1}\big)^{-1}\,d\big(\beta_{z,s,j,+}+\beta_{z,s,j,-})$
for $j\in\{1,2\}.$ Let
$C_{z,\ell,s}:=\left\{t\in K\;\middle|\;(\rho\circ\alpha_{z,l}\circ\alpha_{z,s}^{-1})(t)\frac{\big(w_{z,\ell}\circ\alpha_{z,s}^{-1}\big)(t)}{\big(w_{z,s}\circ\alpha_{z,s}^{-1}\big)(t)}>\sqrt{\varepsilon}\right\}.$
Then
$\big(\beta_{z,s,j,+}+\beta_{z,s,j,-}\big)(C_{z,\ell,s})<\sqrt{\varepsilon}\quad\text{for }j\in\{1,2\}.$
This gives that
$|\beta_{z,s}|(C_{z,\ell,s})<2\sqrt{\varepsilon}.$
Now, since
$C_{z,\ell,s}\subseteq K,$
we get
$|\nu_{s}|(C_{z,\ell,s})\leq|\nu_{s}-\beta_{z,s}|(C_{z,\ell,s})+|\beta_{z,s}|(C_{z,\ell,s})\leq$
$\leq|v_{s}-\beta_{z,s}|(K)+2\sqrt{\varepsilon}\leq\dfrac{2\varepsilon}{m_{K}}+2\sqrt{\varepsilon},$
because
$|v_{s}^{(\rho)}-\beta_{z,s}^{(\rho)}|(K)\leq|v_{s}^{(\rho)}-\beta_{z,s}^{(\rho)}|(\Omega)<\varepsilon,$
and by Lemma 5.2 we then have that
$|v_{s}-\beta_{z,s}|(K)\leq\dfrac{2\varepsilon}{m_{K}}.$
By exactly the same arguments as in the proof of [23, Proposition 10.1] part $(i)\Rightarrow(ii)$ and [24, Proposition 4.1] part $i)\Rightarrow ii)$ , we can deduce that
$|\eta_{z}|\big(\widetilde{A}_{z,\ell}\big)<2\sqrt{\varepsilon}\quad\text{and}\quad|\lambda_{z}\gamma_{z,\ell}|\big(D_{z,\ell}\big)<2\sqrt{\varepsilon},$
where
$\widetilde{A}_{z,\ell}:=\left\{t\in K\;\middle|\;\big((\rho\circ\alpha_{z,\ell})(t)\big)\big(w_{z,\ell}(t)\big)>\frac{\sqrt{\varepsilon}}{\lambda_{z}}\right\}$
and
$D_{z,s}:=\left\{t\in K\;\middle|\;\big((\rho\circ\alpha_{z,s}^{-1})(t)\big)\big((w_{z,s}\circ\alpha_{z,s}^{-1})^{-1}(t)\big)>\lambda_{z}\sqrt{\varepsilon}\right\}$
for all $z\in F$ and $\ell,s\in\{1,\dots,N\}$. Again, since
$|\eta_{z}^{(\rho)}-\mu^{(\rho)}|(\Omega)<\varepsilon\quad\text{and}\quad|\lambda_{z}\gamma_{z,\ell}^{(\rho)}-\nu_{\ell}^{(\rho)}|(\Omega)<\varepsilon,$
we can conclude that
$|\mu|\big(\widetilde{A}_{z,\ell}\big)<2\sqrt{\varepsilon}+\frac{2\varepsilon}{m_{K}}$
and
$|\nu_{\ell}|\big(D_{z,\ell}\big)<2\sqrt{\varepsilon}+\frac{2\varepsilon}{m_{K}}$
for all $z\in F$ and $\ell\in\{1,\dots,N\}$.
Set
$A_{z}:=\bigcup_{\ell=1}^{N}\widetilde{A}_{z,\ell}$
and
$B_{z,s}:=D_{z,s}\;\cup\;\big(\bigcup_{\ell=1}^{N}C_{z,\ell,s}\big),$
for each $z\in F$ and $s\in\{1,\dots,N\}.$ Then
$|\mu|(A_{z})<2N\sqrt{\varepsilon}\left(1+\frac{\sqrt{\varepsilon}}{m_{K}}\right)$
and
$|\nu_{s}|(B_{z,s})<2(N+1)\sqrt{\varepsilon}\left(1+\frac{\sqrt{\varepsilon}}{m_{K}}\right)$
[…]
Now we will prove $(2)\Rightarrow(1)$. Given non-empty open subsets $O,O_{1},\dots,O_{N}$ of $M(\Omega)$, by Remark 5.3 we can find some $\mu\in O,$ $\nu_{\ell}\in O_{\ell}(\ell\in\{1,\dots,N\})$, and a compact $K\subseteq O$ such that
$|\mu|(K^{c})=0\quad\text{and}\quad|\nu_{\ell}|(K^{c})=0\quad\text{for each }\ell\in\{1,\dots,N\}.$
Also, there exists some $\delta>0$ such that
$B(\mu,\delta)\subseteq O\quad\text{and}\quad B(\nu_{\ell},\delta)\subseteq O_{\ell}$
for each $\ell\in\{1,\dots,N\}$. Let
$\varepsilon:=\frac{\delta^{2}}{\left(2(N+1)\left(2+\frac{1}{m_{K}}\right)+\delta\right)^{2}}.$
Choose $F\in\mathcal{F}$ and the families $\{A_{z}\}_{z\in F}$, $\{B_{z,1}\}_{z\in F},\dots,\{B_{z,N}\}_{z\in F}$ satisfying the assumptions of (2) with respect to $K$ and $\varepsilon$.
Let $\widetilde{\mu}_{z},\widetilde{\nu}_{z,1},\dots,\widetilde{\nu}_{z,N}$ be the measures in $M(\Omega)$ given by
$\widetilde{\mu}_{z}(E)=\mu(E\cap A_{z}^{c}\cap K)$
and
$\widetilde{\nu}_{z,\ell}(E)=\nu_{\ell}(E\cap B_{z,\ell}^{c}\cap K)$
for every Borel subset $E$ of $\Omega$, and each $z\in F,\ell\in\{1,\dots,N\}$.
Then, for each distinct $s,\ell\in\{1,\dots,N\}$,
$\Big(T^{*}_{z,\ell}\big(S^{*}_{z,s}(\widetilde{\nu}_{z,s})\big)\Big)^{(\rho)}(E)=\int_{E}\big(w_{z,\ell}\circ\alpha_{z,\ell}^{-1}\big)\big(w_{z,s}\circ\alpha_{z,\ell}^{-1}\big)^{-1}\rho\,d\big(\widetilde{\nu}_{z,s}\circ\alpha_{z,s}\circ\alpha_{z,\ell}^{-1}\big)$
for all $z\in F$ and every Borel subset $E$ of $\Omega$. Hence,
$\Big(T^{*}_{z,\ell}\big(S^{*}_{z,s}(\widetilde{\nu}_{z,s})\big)\Big)^{(\rho)}(E)=\int_{E\cap\alpha_{z,\ell}\big(\alpha_{z,s}^{-1}(B_{z,s}^{c}\cap K)\big)}\rho\frac{(w_{z,\ell}\circ\alpha_{z,\ell}^{-1})}{(w_{z,s}\circ\alpha_{z,\ell}^{-1})}\,d\big(\tilde{\nu}_{z,s}\circ\alpha_{z,s}\circ\alpha_{z,\ell}^{-1}\big)=$
$\int_{\alpha_{z,s}\!\big(\alpha_{z,\ell}^{-1}(E)\big)\cap\big(B_{z,s}^{c}\cap K\big)}(\rho\circ\alpha_{z,l}\circ\alpha_{z,s}^{-1})\frac{(w_{z,\ell}\circ\alpha_{z,s}^{-1})}{(w_{z,s}\circ\alpha_{z,s}^{-1})}\,d\tilde{v}_{z,s}$
for every $z\in F,$ each distinct $s,\ell\in\{1,\dots,N\}$ and every Borel subset $E$ of $\Omega$. By exactly the same arguments as in the proof of [23, Proposition 10.1] part $(ii)\Rightarrow(i)$ we can conclude that
$\Big\|\big(T^{*}_{z,\ell}(S^{*}_{z,s}(\widetilde{\nu}_{z,s}))\big)^{(\rho)}\Big\|\leq\sup_{t\in B_{z,s}^{c}\cap K}\left((\rho\circ\alpha_{z,\ell}\circ\alpha_{z,s}^{-1})(t)\right)\frac{(w_{z,\ell}\circ\alpha_{z,s}^{-1})(t)}{(w_{z,s}\circ\alpha_{z,s}^{-1})(t)}\|\nu_{s}\|.$
Similarly, as in the proof of [23, Proposition 10.1] part $(ii)\Rightarrow(i)$ and [24, Proposition 4.1] part $ii)\Rightarrow i)$, we can deduce that
$\Big\|\big(T^{*}_{z,\ell}(\widetilde{\mu}_{z})\big)^{(\rho)}\Big\|\leq\sup_{t\in A_{z}^{c}\cap K}\big(((\rho\circ\alpha_{z,\ell})(t))\,(w_{z,\ell}(t))\big)\|\mu\|$
and
$\Big\|\big(S^{*}_{z,s}(\widetilde{\nu}_{z,s})\big)^{(\rho)}\Big\|\leq\sup_{t\in B_{z,s}^{c}\cap K}\big((\rho\circ\alpha_{z,s}^{-1})(t)\big)\,\big((w_{z,s}\circ\alpha_{z,s}^{-1})^{-1}(t)\big)\|\nu_{s}\|$
for all $z\in F$ and $s,\ell\in\{1,\dots,N\}$. For each $z\in F$, set
$\psi_{z}=\widetilde{\mu}_{z}+\frac{\sqrt{\displaystyle\sum_{\ell=1}^{N}\sup_{t\in A_{z,\ell}^{c}\cap K}\Big((\rho\circ\alpha_{z,\ell})(t)\Big)\Big(\,w_{z,\ell}(t)\Big)}}{\sqrt{\displaystyle\sum_{l=1}^{N}\sup_{t\in B_{z,s}^{c}\cap K}\Big((\rho\circ\alpha_{z,s}^{-1})(t)\Big)\Big(\,(w_{z,s}\circ\alpha_{z,s}^{-1})^{-1}(t)\Big)}}\;\sum_{s=1}^{N}S^{*}_{z,s}(\widetilde{\nu}_{z,s}).$
By the choice of $\varepsilon$, the triangle inequality we can deduce that $\psi_{z}\in O$ and
$\quad\sqrt{\frac{\displaystyle\sum_{\ell=1}^{N}\sup_{t\in B_{z,s}^{c}\cap K}\Big((\rho\circ\alpha_{z,s}^{-1})(t)\Big)\Big(\,(w_{z,s}\circ\alpha_{z,s}^{-1})(t)\Big)}{\displaystyle\sum_{\ell=1}^{N}\sup_{t\in A_{z,\ell}^{c}\cap K}\Big((\rho\circ\alpha_{z,\ell})(t)\Big)\Big(\,w_{z,\ell}(t)\Big)}}\;T^{*}_{z,s}(\psi_{z})\in V_{s},$
for all $z\in F$ and $s\in\{1,\dots,N\}$. This proves the implication $(2)\Rightarrow(1)$. ∎
Acknowledgement: I am grateful to Professor S. M. Tabatabaie for suggesting linear dynamics on solid spaces as the topic for my research and for introducing to me the relevant literature.